### v1

# Anti-Phishing Training Does Not Work: A Large-Scale Empirical Assessment of Multi-Modal Training Grounded in the NIST Phish Scale

Abstract
Social engineering attacks using email, commonly known as phishing, are a critical cybersecurity threat. Phishing attacks often lead to operational incidents and data breaches. As a result, many organizations allocate a substantial portion of their cybersecurity budgets to phishing awareness training, driven in part by compliance requirements. However, the effectiveness of this training remains in dispute. Empirical evidence of training (in)effectiveness is essential for evidence-based cybersecurity investment and policy development. Despite recent measurement studies, two critical gaps remain in the literature: (1) we lack a validated measure of phishing lure difficulty, and (2) there are few comparisons of different types of training in real-world business settings.
To fill these gaps, we conducted a large-scale study ($N=12{,}511$) of phishing effectiveness at a US-based financial technology (“fintech”) firm. Our two-factor design compared the effect of treatments (lecture-based, interactive, and control groups) on subjects’ susceptibility to phishing lures of varying complexity (using the NIST Phish Scale). The NIST Phish Scale successfully predicted behavior (click rates: 7.0% easy to 15.0% hard emails, p $<$ 0.001), but training showed no significant main effects on clicks (p = 0.450) or reporting (p = 0.417). Effect sizes remained below 0.01, indicating little practical value in any of the phishing trainings we deployed. Our results add to the growing evidence that phishing training is ineffective, reinforcing the importance of phishing defense-in-depth and the merit of changes to processes and technology to reduce reliance on humans, as well as rebuking the training costs necessitated by regulatory requirements.
1 Introduction
Phishing attacks critically impact organizations through operational incidents, data breaches, and financial losses. They constitute the primary attack vector to breach organizations as highlighted by industry studies [1] and the US Federal Bureau of Investigation (FBI) [2, 1]. To prevent this class of attacks, organizations employ technical controls (e.g., spam filters, attachment analyzers) complemented by cybersecurity awareness training [3, 4].
Although these anti-phishing trainings are costly, their effectiveness is in question. Organizational security awareness programs often focus on compliance metrics rather than meaningful behavioral change [5]. Meanwhile, recent large-scale evidence suggests that current approaches have limited impact [4]. To better evaluate the effectiveness of anti-phishing training, two questions arise. First, are some kinds of training more effective than others? Second, how does their effectiveness vary with the difficulty of the phishing lure?
In this paper, we report the results of a measurement study to address these questions. We situate the study in an operational environment [6], set at a US-based international fintech firm with 12,511 employees across multiple business units. We compared three experimental conditions: a control group (no training), traditional lecture-based training, and lectures augmented with interactive phishing exercises. Subjects in each condition received lures of varying complexity, which we assigned difficulty levels using the NIST Phish Scale. This study design enables rigorous evaluation of training effectiveness across standardized difficulty levels on a large population, improving on previous works that had either been small-scale or lacking in systematic difficulty assessment [7, 8].
Our findings were unfavorable for the conventional anti-phishing trainings available from cybersecurity training vendors. Neither mode of training produced statistically significant improvements in phish click reduction (average: 10.4%) or phish reporting behavior (average: 9.6%). However, the NIST Phish Scale did predict user behavior, with phishing lure difficulty having a highly significant effect on click-through rates (F(2, 12086) = 41.415, p $<$ 0.001, $\eta^{2}$ = 0.007). Thus our study shows that, while the NIST Phish Scale provides a means of measuring the difficulty of different lures, the trainings we applied did not help subjects perform better on those lures.
We summarize our main contributions below:
  1. 1.
    Large-Scale Validation of the NIST Phish Scale. We provide the most comprehensive empirical validation of the NIST Phish Scale to date, demonstrating its predictive utility across 12,511 participants in an operational environment, showing strong correlation between difficulty ratings and user susceptibility.
  2. 2.
    Realistic Assessment of Training Effectiveness. Our results challenge optimistic assumptions about the effectiveness of phishing training. We found no statistically significant impact of training on either click rates or reporting behavior. Although these findings align with recent large-scale studies from the healthcare [4] and banking sectors [9], our study addresses explicit calls from the research community for large-scale ecologically valid studies combining standardized assessments of difficulty inside of an operational context [10, 6].
Significance: Our findings have important implications for researchers, organizations, and regulators. For the research community, we provide empirical assessment of training effectiveness using rigorous methodology and standardized difficulty measures. For organizations, our results suggest the need for realistic expectations about training outcomes and highlight the importance of complementing awareness programs with robust technical controls rather than relying primarily on human-centered defenses. Our validation of the NIST Phish Scale’s predictive utility enables organizations to use this tool for benchmarking the difficulty of their phishing simulations. For regulators, our work contributes to the growing body of evidence that organizational cybersecurity strategy should emphasize technical defenses [1] rather than placing the burden on humans — requiring anti-phishing training is costly and has minimal effect.
2 Background and Related Work
In a phishing attack, attackers send messages (phishing lures) to targets whom they wish to persuade to act against their own interests. Attackers seek results such as sensitive information, account credentials, or the ability to control the victim’s machine [1, 11]. Modern phishing campaigns target individuals and organizations across multiple attack vectors, applying conventional email phishing as well as SMS-based attacks (“smishing”) and voice-based deception (“vishing”) [12, 13]. Recent attacks have begun to incorporate Generative AI tools to produce tailored and realistic messages. The severity, persistence, and increasing sophistication of these attacks has established phishing detection and prevention as a critical focus within both academic cybersecurity research and industry practice [14, 6].
In this section, we summarize current defenses against phishing (Appendix 2.1), discuss metrics for evaluating phishing effectiveness (Appendix 2.2), and review prior empirical studies (Appendix 2.3) and knowledge gaps (Appendix 2.4).
2.1 Phishing Defenses
Phishing poses substantial risks to organizations, who therefore erect multiple layers of defenses against these attacks. We distinguish automated defenses from human-focused ones.
2.1.1 Automated Defenses
Organizations employ a range of technological defenses to mitigate phishing risks. These include spam filters, attachment analyzers, and URL checkers [15, 14, 1]. More recently, machine learning approaches have been applied to detect phishing content. Long Short-Term Memory (LSTM) models and optimized Random Forest algorithms [16] have demonstrated high accuracy in detecting phishing attempts. As with everything in cybersecurity, phishing is a cat-and-mouse game. Adversaries have begun to leverage artificial intelligence to create more convincing lures. Large language models (LLMs) can generate personalized and context-aware phishing emails that are difficult for traditional filters to detect [3]. Techniques like typosquatting, domain spoofing, and watering hole attacks exacerbate this issue [1, 14].
2.1.2 Human-Focused Defenses: Training and Awareness
Defending against phishing is understood not as solely a technical problem, but rather as a socio-technical one [6, 13]. Spear phishing and emerging phishing attacks present sophisticated threats that overcome automated defenses [17, 18], so human awareness of phishing remains necessary to successful defense [19, 20].
There are many approaches to training humans to be resilient to phishing attacks. Methods range from lecture-style training involving passive videos and readings [21] to simulation feedback providing immediate feedback after interaction with phishing simulations [22, 23]. Interactive training uses gamification or scenario-based learning to engage participants [24, 25], while just-in-time training alerts users in real time during risky behavior [26].
Anti-phishing trainings are frequently included within the cybersecurity awareness trainings ubiquitous in organizations. Legal mandates drive the widespread adoption of such training programs, with HIPAA requiring security awareness training for healthcare organizations [27], GDPR encouraging staff training on cybersecurity practices [28], and PCI DSS mandating training for those handling cardholder data [29]. Industry standards and frameworks further reinforce these requirements, with ISO 27001 calling for human resource security controls including awareness programs [30] and the NIST Cybersecurity Framework advising phishing simulations under the “Identify” function [31].
2.2 Measuring Effectiveness of Phishing Training
The effectiveness of phishing training can be assessed based on how resilient a target becomes to phishing lures. Appendix 2.2.1 describes how lures can be categorized by their complexity, and then Appendix 2.2.2 gives the standard metrics for target resilience.
2.2.1 Measuring Training Lure Difficulty
TABLE I: NIST Phish Scale for Detection Difficulty [32, 33].
Cues \ Alignment Weak Medium Strong
Few Medium Hard Hard
Some Easy-to-Medium Medium Hard
Many Easy Medium Medium
TABLE II: Summary of Prior Experiments on Phishing and Training Interventions. Studies are organized by research context (Lab vs. Real-World). Context: “Lab” denotes research simulations, where subjects are not in their work roles. “Real-World” denotes studies where participants were employees acting in their job capacity. Complexity measure: Whether the study controlled for the difficulty of the phishing tasks.  indicates the use of an open control such as NIST Phish Scale, and  indicates no control. Hypotheses: How these works informed the hypotheses in the present study (cf. Table III).
Work Context Population Sample size Complexity Measure Compared Training Approaches? Hypotheses
[8] (2024) Lab University students 96 H3
[34] (2024) Lab University students 117  (NIST Phish Scale) H4
[7] (2022) Real-world Employees 119 H2
[22] (2017) Real-world University students 150 H2, H3
[23] (2014) Real-world Employees 1,359 H2, H4
[9] (2024) Real-world Banking employees “Thousands” H2
[35] (2023) Real-world Finance employees 5,000 H2
[4] (2025) Real-world Healthcare employees 19,500+ H2
Ours Real-world Fintech org. 12,511  (NIST Phish Scale)
To assess phishing lure difficulty, the NIST Phish Scale categorizes them along two dimensions as depicted in Table I. Phishing Cues represent observable errors or inconsistencies in the email that might alert users to its fraudulent nature, such as spelling mistakes, suspicious URLs, or formatting irregularities [34]. Premise Alignment measures the relevance and familiarity of the email’s subject matter to the recipient’s typical email environment and organizational context [36]. A high alignment, messages that closely resemble emails users receive daily, combined with few obvious cues yields the most deceptive and challenging-to-detect phishing emails [17]. The scale allows organizations to benchmark lure complexity and contextualize simulation outcomes, though it requires sufficiently large sample sizes for reliable use and consistent application [37].
2.2.2 Training Outcome Metrics
The effectiveness of anti-phishing training is commonly measured in three ways:
  • Open Rate captures the percentage of emails opened, though this measurement can be deceptive since different email clients and security settings can affect whether an email is recorded as opened [36] we therefore largely ignore this measure.
  • Click-Through Rate (CTR) measures the percentage of recipients who click on malicious links, serving as a primary indicator of phishing susceptibility [22, 23].
  • Reporting Rate tracks phishing incidents reported to security teams, representing users’ proactive security behaviors and ability to recognize suspicious content [18, 38].
2.3 Empirical Studies of Anti-Phishing Training
For reviews of knowledge about susceptibility to phishing and about anti-phishing training, we refer the reader to recent works by Marshall et al. [21] and Franz et al. [10]. This section focuses on the aspect of the literature to which this study contributes. We discuss the settings of prior work (Appendix 2.3.1), the assessments of different trainings (Appendix 2.3.2), and measurement and evaluation methods (Appendix 2.3.3).
Figure 1: Illustration of our methodology. Our two-factor design randomly assigns different trainings to different subjects, and then randomly sends phishing lures of varying complexity to the subjects. We compared between-subject performance using lure click rate and lure report rate in order to test four hypotheses derived from prior work.
2.3.1 Scale and Context of Prior Work
Many phishing training studies have been conducted in controlled laboratory settings, where researchers benefit from a high degree of experimental control. Studies of this sort have provided numerous insights, including that active learning and immediate feedback can improve short-term detection rates, and have explored how individual differences affect training outcomes [7, 21]. However, as noted by Franz et al. [10], these controlled settings may lack ecological validity. In laboratory environments, participant samples are typically small and consist of homogeneous groups such as academics and students. Due to this, and the artificial nature of lab tasks, these studies may not accurately reflect the complexities and distractions of real-world environments. This raises questions about the generalizability of laboratory findings to operational contexts, where additional factors can alter training effectiveness.
In response, a growing number of studies have begun examining phishing training in real-world organizational environments. Doing et al. [9] analyzed phishing reporting behaviors across thousands of bank employees, while Ho et al. [4] conducted a comprehensive evaluation with over 19,500 healthcare employees across 10 simulated phishing campaigns, reporting minimal effectiveness of both annual cybersecurity awareness training and embedded anti-phishing training. To date, however, these larger studies have lacked a standardized method for assessing the difficulty of the phishing lures in question [4, 9], or have focused primarily on observational analysis rather than controlled training interventions with standardized difficulty measures [9].
Our study contributes to the evidence on real-world human phishing resilience, with data from approximately 12,000 subjects in a fintech enterprise, while maintaining experimental control and measuring lure difficulty with the NIST Phish Scale framework. This approach allows us to examine whether the modest training effects we observe align with Ho et al.’s findings, while providing additional insights into how training effectiveness varies by phishing sophistication.
2.3.2 Training Intervention Approaches
Franz et al. [10] provide a comprehensive taxonomy identifying five categories of phishing interventions: training, awareness, education, notifications, and design interventions. Our study contributes empirical data within their training category, specifically comparing lecture-based delivery against interactive modalities. While Schöni et al. [8] demonstrated effectiveness of personalized training approaches, and our work examines scalable training modalities that can be deployed across demographically diverse cohorts without individualization, recent large-scale studies suggest that the practical benefits of such training interventions may be more limited than previously assumed [4]. Our study builds on this foundation by applying the NIST Phish Scale to measure lure difficulty in line with current guidance, operationalizing the scale at unprecedented scale, and measuring outcomes (clicking, reporting) post-intervention, allowing for practical comparison across conditions and lure complexity.
2.3.3 Measurement and Evaluation Methods
A critical gap in prior work has been the lack of standardized difficulty assessment for phishing simulations. Barrientos et al. [39] conducted early validation work on the NIST Phish Scale, while Dawkins and Jacobs [33] provided the official implementation guidelines. However, no studies have operationalized the scale at enterprise level nor examined how difficulty interacts with training effectiveness. We applied the scale across varied training modalities and measuring multiple outcome measures (clicking, reporting) post-intervention.
2.4 Research Gaps
Table II summarizes the preceding analysis as two gaps. First, prior work has primarily focused on individual-level factors influencing phishing susceptibility. The interaction between training modality and phishing difficulty — a key question for optimizing organizational defenses — has received little attention. Second, despite widespread adoption of phishing training, few studies rigorously assess outcomes against standardized difficulty measures or operate at sufficient scale for robust inference.
This study addresses those gaps by applying the NIST Phish Scale [33] to evaluate training effectiveness across varied phishing difficulty levels in an enterprise deployment with over 12,000 participants. In this study, our objectives are to: (1) operationalize the NIST framework at scale, (2) examine how difficulty influences detection and reporting across training modalities, (3) provide comparative evidence on lecture-based versus interactive approaches, and (4) offer guidance for large-scale awareness efforts under conditions of modest effectiveness [4]. This enables evidence-based decision-making on phishing defense strategies that account for both training design and message complexity.
3 Methodology
In this section, we describe the methodology of our measurement study, including our hypotheses, experimental design, participant demographics, training modalities, and phishing simulations. An illustration is given in Figure 1. Our two-factor study design controls for the effect of training modality and task complexity. We make the first real-world application of the NIST Phish Scale framework [33], allowing us to evaluate the effectiveness of different training modalities across standardized difficulty levels in an operational environment.
TABLE III: Summary of study hypotheses, their theoretical grounding, and relation to prior work. The hypotheses address the effects of phishing email difficulty, training, and training modality on detection and reporting behaviors.
Hypothesis Theoretical and Empirical Basis Relation to Prior Work
H1 (Phish Scale Effect): Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will exhibit higher click-through rates than those with lower difficulty. Grounded in the foundational NIST Phish Scale framework [33] and validation studies [39]. Supported by phishing cue research [34] and systematic email analysis [40]. The NIST Phish Scale has been validated in lab settings [39]. We provide the first large-scale enterprise validation.
H2 (Training Impact — General): Training will increase the detection and reporting of phishing emails. Evidence of training efficacy, reduced click rates [22], improved detection accuracy [7], and short-term benefits [21]. Disputed by a recent study [4]. Tests whether training effects from small-scale studies [22, 7] replicate at scale, addressing the call for large-sample validation [10].
H3 (Interactive Training Effect — Reporting): Interactive components (Trained+Exercise) will lead to significantly higher reporting rates than lecture-only training (Treatment A). From Franz et al.’s modality taxonomy [10], and evidence of the effectiveness of active and feedback-rich training [22, 21, 8]. Builds on academic findings of interactive training benefits [22, 21]. We provide the first test at organizational scale.
H4 (Interaction Effect): There will be a significant interaction between training modality and phishing difficulty, with interactive training being more beneficial for easier lures. Emerges from cognitive load and decision strategy research [19], difficulty-aware detection patterns [40], and findings on human limitations [6]. Addresses the gap identified by Franz et al. [10] regarding interaction effects between training and phishing complexity, largely unexplored empirically.
3.1 Hypotheses
Based on prior work on phishing training efficacy [22, 23, 5, 7, 10, 4] and documented differences in risk perception [6, 9], we formulated four hypotheses to test in this work. Table III states each hypothesis, its basis, and its relation to prior work.
To summarize, H1 represents the foundational assertion of the NIST Phish Scale framework [33]; H2 and H3 capture the intended effects of training interventions, with H3 specifically testing whether interactive modalities provide measurable benefits over passive learning approaches; and H4 reflects our expectation that training effectiveness may diminish for highly sophisticated phishing attempts, as advanced phishing emails may exceed individuals’ detection capabilities regardless of training modality.
With respect to hypothesis novelty and prior testing, H1 has been tested in small academic settings [39] but never at enterprise scale; H2 and H3 build on established training effectiveness research [22, 7, 21] but test these findings in a large-scale operational context using standardized difficulty measures; and H4 is novel, addressing the research gap identified by Franz et al. [10] regarding interactions between training modality and phishing complexity, which has received minimal empirical attention in prior work.
3.2 Study Design
To test the hypotheses, we employed a between-subjects experimental design [35]. Subjects were randomly assigned to one of three experimental conditions using our organization’s training management system. Each condition reflected the phishing training they received (Appendix 3.3). Subjects were recruited (Appendix 3.4), trained, and later received simulated phishing emails (Appendix 3.5). We compared subjects’ performance in handling these phishing emails between the treatments using several standard metrics (Appendix 3.6).
3.3 Training Modalities
For the purposes of this research, the principal researcher collaborated with a cybersecurity training vendor who produces comprehensive security awareness training materials, phishing simulation exercises designed to augment those materials, and implementation assistance for organizational deployment. This partnership enabled access to professionally developed training content and established phishing templates while maintaining the experimental rigor necessary for our evaluation.
The vendor offers three distinct training modalities, though we focused our evaluation on two of these approaches. The primary modality for cybersecurity awareness training consists of lecture-based videos that discuss the dangers of various social engineering attacks, including phishing recognition tactics, social engineering techniques, and organizational security protocols. Our vendor, along with many others in the industry, supplements this foundational training with interactive phishing exercises that present users with samples from a library of approximately 20 phishing templates. These exercises guide users through identifying suspicious elements and security deficiencies in emails that roughly correspond to the cues assessed by the NIST Phish Scale.
The vendor also provides a third modality referred to as “in-place training.” This approach occurs during phishing simulations when users click on malicious links, instead of being directed to a malicious website, they are immediately redirected to a training webpage that explains why the email they clicked should have been detected as fraudulent. However, decades of research into the effectiveness of embedded training environments [23] and recent comprehensive research by Ho et al. [4] convinced us that monitoring the results of in-place training was not necessary for our research objectives. Ho’s large-scale study across 19,500 healthcare employees found minimal effectiveness of embedded training approaches, providing thorough and convincing evidence that this modality offers limited practical value compared to the lecture-based and interactive approaches we chose to evaluate. We employed two distinct training approaches in collaboration with our cybersecurity training vendor, building on established taxonomies of phishing interventions [10]. Both modalities fall within Franz et al.’s “training” category [10], specifically comparing passive lecture-based delivery Figure 2against interactive approaches with embedded learning elements.
Treatment A: Lecture-Based Training: The lecture-style training consisted of fifteen brief instructional videos, each lasting 1-2 minutes, covering core cybersecurity awareness topics including phishing recognition tactics, social engineering techniques, and organizational security protocols [22]. The videos specifically addressed phishing identification strategies, emphasizing visual cues such as suspicious URLs, sender inconsistencies, and urgency tactics commonly employed by attackers. Following the video content, participants completed a knowledge assessment quiz reviewing key concepts from the training materials. This approach represents the traditional passive learning modality widely deployed in organizational cybersecurity training programs [23].
Figure 2: Traditional Training: Videos and Quiz
Treatment B: Interactive Training (Training + Exercise): The interactive group received the identical lecture-style training described above, plus an additional hands-on component consisting of simulated phishing exercises with immediate feedback. During these exercises, participants were presented with sample emails and asked to identify whether each message was legitimate or malicious. The system provided real-time feedback explaining the reasoning behind correct identification, highlighting specific phishing indicators, and reinforcing appropriate response behaviors such as using the “Report Phish” button in the email client [22]. This interactive component was designed to provide experiential learning opportunities that complement the passive video instruction, following research suggesting that active engagement enhances retention and behavioral change in cybersecurity training contexts [33].
Control Group: The control group received no cybersecurity training prior to the phishing simulation phase. This group was included to establish baseline susceptibility measurements. Control group participants received the same training materials after completion of the study to ensure organizational compliance with cybersecurity awareness requirements. Organizational constraints compelled us to use a relatively small control group, as ACME Corp. required that the majority of employees receive some form of cybersecurity training for compliance purposes.
3.4 Subject Recruitment
Study Context: The subjects were employees at a U.S.-based international financial technology (“fintech”) firm. Participants represented a diverse cross-section of the organization, including employees from finance, technology, operations, customer service, and administrative functions. All participants were full-time employees with active email accounts who regularly use organizational communication systems and speak English.
We excluded participants who: (1) had inactive email accounts during the simulation period, (2) were on extended leave during the study period. Additionally, participants who received phishing simulations but had technical delivery failures (bounced emails, system errors) were excluded.
Subject Recruitment: This study leveraged organizational security training data collected as part of routine cybersecurity awareness training. The training was introduced through official organizational communications emphasizing its importance for maintaining security compliance and protecting organizational assets. This training was mandatory due to the organization’s need for regulatory compliance (cf. Appendix 2.1.2).
Data were collected from Dec. 2024 – Apr. 2025. Participants had a one month window to complete their assigned training modules, with automated reminders sent at regular intervals. Training completion was tracked and verified.
Randomization: Subjects wre randomly assigned to each treatment. To mitigate bias by organizational role or function, we used stratified sampling [41] across departments to ensure representative distributions in each treatment group.
Completion Rates: Our data came from the 12,511 participants who completed the full experimental protocol. Table IV shows the completion rates by treatment type. The comparable completion rates across conditions (ranging from 87.3% to 89.1%) suggest that our random sample was preserved between recruitment and completion.
TABLE IV: Recruitment Pool and Completion Rates by Experimental Condition. Note the comparable drop-out rates by condition.
Condition Recruited Completed
Training Only (Treatment A) 6,758 6,023 (89.1%)
Training + Exercise (Treatment B) 6,764 6,026 (89.1%)
Control Group 529 462 (87.3%)
Total 14,051 12,511 (89.0%)
Ethical Considerations: We acknowledge that employee consent in workplace security training cannot be considered fully voluntary due to inherent power dynamics and employment dependencies [5]. However, several ethical safeguards were incorporated: all employees received security training regardless of group assignment, individual performance data was anonymized and never used for personnel decisions, and phishing simulation templates excluded content designed to cause psychological distress (Appendix 3.5).
The research protocol was reviewed by the academic institution’s IRB, which determined that analysis of operational security training data does not constitute human subjects research under federal guidelines, as it involves analysis of data collected for legitimate business purposes rather than research-specific data collection [5].
See Appendix 3 for an extended discussion of ethics.
3.5 Phishing Simulations
After training, all subjects received one simulated phishing email, of varying complexity. We first describe how we defined complexity using NIST Phish Scale, and then how we deployed the phishing simulation.
3.5.1 Template Selection using NIST Phish Scale
To establish controlled and consistent evaluation of participant responses, we systematically assessed and categorized the difficulty of each simulated phishing email using a structured expert review process informed by the NIST Phish Scale framework [33]. Phishing templates were initially selected from our cybersecurity training vendor’s library, which provided preliminary difficulty assessments (easy, medium, hard). We then independently evaluated these templates using the official NIST Phish Scale User Guide methodology to ensure accurate difficulty classification.
The assessment team comprised three members of ACME Corp.’s IT security staff, each with substantial experience in identifying real-world phishing attacks and analyzing user behavior in response to simulated campaigns. This multi-rater approach follows established practices for enhancing reliability and validity of stimulus assessment in cybersecurity research [19]. Two team members independently evaluated each phishing email template, rating the perceived human difficulty in correctly identifying the email as malicious using structured criteria derived from the NIST Phish Scale methodology.
Our implementation followed the official NIST Phish Scale User Guide [33], evaluating templates across two primary dimensions. For the Phishing Cues dimension, raters assessed observable indicators that might alert users to fraudulent content, including spelling errors, suspicious URLs, formatting inconsistencies, and sender authenticity markers [34]. Fewer or less obvious cues indicated higher difficulty ratings.
For the Premise Alignment dimension, raters evaluated how well each email’s subject matter, sender, and content aligned with typical organizational communications that ACME Corp. employees receive in their regular workflow. This assessment considered whether the email appeared to originate from familiar organizational systems (e.g., HR portal, IT helpdesk, financial systems), addressed topics relevant to employee responsibilities, and used language and formatting consistent with legitimate internal communications [6]. Higher premise alignment—emails that closely resembled routine organizational correspondence—contributed to increased difficulty ratings.
Following independent assessment, the two raters compared their evaluations for each email template. Any disagreements regarding difficulty classification were discussed to identify the specific cues or contextual elements that led to differing conclusions. A third senior IT security expert then reviewed each email and the rationales provided by the independent raters, serving as a final validator to reach consensus difficulty ratings for each template. This collaborative validation process ensured that our phishing simulation represented a controlled range of difficulty levels essential for evaluating training effectiveness across standardized complexity measures [39].
Difficulty Categorization and Distribution: Following the NIST framework, we created a 3x3 classification matrix combining cue visibility (few-low, some-medium, many-high) with premise alignment (weak-low, medium, strong-high). Templates were assigned composite difficulty scores by summing both dimensions, with scores ranging from 2 (easy: high cues, low alignment) to 6 (hard: low cues, high alignment).
Our final simulation campaign included nine distinct phishing templates across three difficulty categories:
  • Easy: 5,721 emails (45.7% of total simulations)
  • Medium: 2,279 emails (18.2% of total simulations)
  • Hard: 4,511 emails (36.1% of total simulations)
(a) “Easy”: low premise, high cues.
(b) “Hard”: high premise, low cues.
Figure 3: Example templates. The “Easy” template has low premise alignment and high cues; the “Hard” is the inverse.
3.5.2 Simulation Deployment
Phishing simulations were distributed among participants within a three-month window after training completion. Each participant received one phishing email. Templates were randomly assigned to ensure balanced distribution across experimental conditions.
The deployed templates were constrained by organizational considerations. ACME Corp.’s management requested the exclusion of highly distressing content (e.g., emergency family situations, personnel security scenarios) to minimize employee psychological impact. This constraint reflects the practical limitations of conducting research within operational environments, where employee welfare considerations influence experimental designs [9].
Technical delivery failures, including bounced emails and system errors, resulted in participant exclusion from analysis to maintain data integrity.
3.6 Metrics
Our outcome measures followed established practices in phishing research [22, 23].
  • Open Rate captured the percentage of recipients who opened the simulated phishing email, measured via embedded tracking pixels, though we acknowledge this metric can be influenced by email client settings and security configurations [36].
  • Click-Through Rate (CTR) measured the percentage of recipients who clicked on malicious links within the phishing emails, serving as our primary indicator of phishing susceptibility [7].
  • Reporting Rate tracked the percentage of recipients who reported the suspicious email to the organization’s security team using the integrated “Report Phish” button, representing proactive security behavior and successful threat recognition [9].
The simulation platform tracked these metrics as follows: email open rates were tracked via embedded tracking pixels, click-through rates were tracked via unique URLs, and reporting rates were tracked via integration with the organization’s security incident reporting system.
3.7 Analysis
Data Validation and Quality Assurance: Prior to analysis, we conducted extensive data validation to ensure analytical integrity. This included verification of random assignment effectiveness across groups, assessment of technical delivery success rates, and examination of potential outliers or data anomalies. Participants with incomplete data due to technical failures, extended leave, or incomplete training were excluded from analysis to maintain dataset quality [39].
Statistical Approach: We employed two complementary analytical approaches to examine training effectiveness. Our primary analysis used two-way analysis of variance (ANOVA) to test main effects and interactions between training modality (Control, Training Only, Training + Exercise) and phishing difficulty level (Easy, Medium, Hard) on our outcome measures. Note that although ANOVA with binary outcomes violates standard normality assumptions, this approach aligns with established practices in phishing research [22, 23] and enables direct comparison with prior studies.
To address the limitations of ANOVA with binary data, we conducted confirmatory analyses using logistic regression models appropriate for binary outcomes (clicked/didn’t click, reported/didn’t report). The logistic regression models employed maximum likelihood estimation and provided odds ratios for effect size interpretation. Both analytical approaches yielded consistent conclusions regarding training effectiveness, with ANOVA providing F-statistics for comparison with prior literature and logistic regression offering more statistically appropriate modeling of our binary outcomes. We report both approaches to provide comprehensive statistical evidence while acknowledging the methodological trade-offs inherent in each approach. All analyses were conducted using Python 3.9 with statsmodels and scikit-learn libraries for statistical modeling and cross-validation.
3.8 Threats to Validity
We encourage the reader to take several important limitations into consideration when interpreting the results of our study. We consider three threats to validity in our work: threats to construct, internal, and external validity [42].
Construct Validity: Construct validity is concerned with whether we operationalized the desired constructs of our study appropriately. In this study, we worked with two primary constructs: (1) the particular interpretations of the two different types of phishing training we deployed, which depend on our cybersecurity training vendor; and (2) our assessment of the difficulty of phishing lures.
It is possible that our cybersecurity vendor’s training was simply lackluster, hence its failure to produce any significant improvement in subject performance. However, note that our results align with prior work that used different trainings. Meanwhile, to control for subjectivity in rating phishing lures, we applied the NIST Phish Scale and used multiple independent analysts to judge difficulty.
Internal Validity: Internal validity threats affect the reliability of conclusions we have about the relationship between the cause and the effect we observe. For our study, we used a stratified random sampling method in order to mitigate the potential biases related to job role and other organizational issues, allowing us to use that as a proxy for other factors such as technical background, educational attainment, etc. Due to the large sample sizes for our two different treatment groups, our study has significant statistical power; however, due to organizational constraints, we were required to limit the sample size for our control group.
External Validity: External threats impact the generalizability of our research. Our study was situated in the USA at a fintech company. The fintech industry context may limit generalizability to other sectors, as financial services employees may have different baseline levels of security awareness when compared to other industries. However, given the minimal effect sizes measured in our study, it is unclear how large of a threat this is. The study’s focus on immediate post-training effects also limits our understanding of long-term training impact. Future research should examine whether the modest effects we observed persist over time or diminish as training recency decreases, or if they improve with volume or other dimensions.
4 Results
TABLE V: Summary of Training Effectiveness Results. Average rates show that trained groups performed similarly to control group, with modest differences across conditions.
Aspect Results
Effect Size Generally modest, consistent with recent large-scale findings [4]
Click Rates
Overall Click Rate 10.4% across all participants
 Control Group Average 9.8%
 All Trained Average 10.5%
  Training Only 10.6%
  Training + Exercise 10.4%
Report Rates
Overall Report Rate 9.6% across all participants
 Control Group Average 8.9%
 All Trained Average 9.7%
  Training Only 9.5%
  Training + Exercise 9.9%
Statistical Effects
Training Effect on Click Rates p = 0.450 (not significant)
Training Effect on Report Rates p = 0.417 (not significant)
NIST Phish Scale Effect p $<$ 0.001 (highly significant)
 Easy Difficulty 7.0% click rate
 Medium Difficulty 8.7% click rate
 Hard Difficulty 15.0% click rate
Interaction Effects Revealed nuanced patterns with difficulty levels
Our analysis examined the effectiveness of phishing training across 12,511 participants randomly assigned to three experimental conditions. Table V gives a summary of our key results. Next we present our findings organized by hypothesis, with detailed statistical analysis and effect size reporting for each. Our large-scale study revealed several key findings regarding the effectiveness of phishing training interventions and the predictive utility of the NIST Phish Scale.
Figure 4: Comparison of effect sizes of phishing susceptibility. Bars show the variance explained by phishing difficulty (per the NIST Phish Scale) for each training intervention.
4.1 H1: Phish Scale Effect
Training showed mixed effectiveness in our study. While we observed statistically significant effects for some outcomes, effect sizes were generally modest, consistent with recent large-scale findings suggesting limited practical impact of current training approaches [4]. Overall click rate across all participants was 10.4% and overall report rate was 9.6%. Training on both click rates (p = 0.469) and report rates (p = 0.448) were not statistically significant in the primary analysis, though interactions with difficulty levels revealed more nuanced patterns.
The NIST Phish Scale successfully predicted user behavior, supporting H1. Phishing lure difficulty had a highly significant effect on click-through rates (F(2, 12086) = 41.415, p $<$ 0.001, $\eta^{2}$ = 0.007), validating the scale’s utility for measuring phishing complexity. Higher difficulty emails consistently produced higher click rates across all experimental conditions:
  • Easy difficulty: 7.0% average click rate
  • Medium difficulty: 8.7% average click rate
  • Hard difficulty: 15.0% average click rate
This statistically significant difficulty gradient (F(2, 12086) = 41.415, p $<$ 0.001, $\eta^{2}$ = 0.007) confirms that the NIST Phish Scale effectively captures factors that influence user susceptibility to phishing attempts, providing empirical validation for its use in organizational contexts [33, 39].
4.2 H2: Overall Training Effects
Contrary to H2, training did not produce statistically significant improvements in either click reduction or reporting behavior. Main effect analysis showed no significant difference between trained and untrained groups for click rates (F(2, 12086) = 0.800, p = 0.450) or report rates (F(2, 12086) = 0.874, p = 0.417). This finding aligns with recent large-scale evidence suggesting minimal effectiveness of current training approaches [4].
However, examining treatment effects relative to control group baseline reveals some notable patterns:
Click Rate Changes (relative to control):
  • Training Only: +4.5 percentage points for easy emails, +7.3 for medium, -5.2 for hard
  • Training + Exercise: +3.9 percentage points for easy emails, +5.5 for medium, -5.2 for hard
Report Rate Changes (relative to control):
  • Training Only: +0.1 percentage points for easy emails, +4.7 for medium, +1.9 for hard
  • Training + Exercise: +0.8 percentage points for easy emails, +5.7 for medium, +1.7 for hard
The mixed direction of these effects suggests that training may have differential impacts across difficulty levels, though these did not reach statistical significance in the main effects analysis. The effect of training did not change significantly over time, as we observed similar patterns across the three-month simulation period.
4.3 H3: Interactive Training Effects
H3 was not supported. Interactive training (Training + Exercise) did not produce significantly higher reporting rates than traditional training alone. While the interactive group showed numerically higher reporting rates across difficulty levels, these differences were not statistically significant. The coefficient estimate for Training + Exercise relative to Training Only was modest and non-significant across all difficulty levels.
This finding contrasts with some prior work suggesting benefits of interactive training components [22], but aligns with recent large-scale studies showing limited training effectiveness regardless of modality [4].
4.4 H4: Training-Difficulty Interactions
H4 was not supported. We found no statistically significant interaction between training modality and lure difficulty for click rates (F(4, 12086) = 3.135, p = 0.014) or report rates (F(4, 12086) = 0.517, p = 0.723). While there was a marginal interaction effect for click rates, this primarily reflected the pattern that training appeared to increase clicks on easier emails while potentially reducing clicks on harder emails.
Coefficient analysis revealed some interesting interaction patterns:
  • Training Only × Hard: -9.67 percentage points (p = 0.002)
  • Training + Exercise × Hard: -9.09 percentage points (p = 0.004)
These significant negative interactions suggest that training may provide some protection against the most difficult phishing emails, even though overall training effects were not significant. It is worth considering that the training vendor created the phishing templates and training materials, which may introduce alignment between the cues emphasized in training and those present in the simulated emails. This alignment could artificially enhance the apparent effectiveness of training for certain email types, particularly those closely modeled after the vendor’s examples. As a result, the observed reductions in click rates for hard emails may partially reflect this overlap rather than a generalizable improvement in user detection skills. Future studies should consider using independently developed phishing templates or varying template sources to better assess the robustness of training effects across diverse phishing scenarios.
4.5 Effect Sizes and Practical Significance
While some individual coefficients reached statistical significance, effect sizes were generally small ($\eta^{2}$ $<$ 0.01 for all main effects), raising questions about practical significance [7]. The largest meaningful effect was the difficulty gradient ($\eta^{2}$ = 0.007 for click rates), confirming the NIST Phish Scale’s predictive validity.
Training effects, where present, were modest in magnitude. Even statistically significant interaction effects represented changes of less than 10 percentage points, suggesting that while training may provide some benefits, these are limited in scope and may not justify the substantial organizational investment required for comprehensive training programs [4].
These findings underscore the need for realistic expectations about training effectiveness and suggest that organizations should complement training with other defense mechanisms rather than relying on awareness programs as primary protection against phishing threats [10].
Figure 5: Comparison of effect sizes ($\eta^{2}$) for key factors in phishing susceptibility. The NIST Phish Scale (difficulty) shows the strongest and only practically meaningful effect ($\eta^{2}=0.007$, $p<0.001$), while training and interaction effects are negligible ($\eta^{2}<0.01$).
4.6 Other Measurements
The final question we considered was whether there was any relationship between training time and detection ability, e.g., decay of training effectiveness. In our study, subjects received the phishing simulation email between 0–4 months after the training. Figure 6 shows the click rate by treatment group, discriminating by lure difficulty. We observe no discernible trend in the results as a function of time. Given the result of the visual inspection, we did not perform statistical tests of this aspect.
Figure 6: Temporal Impact on Detection. We observed no discernible relationship between time-since-training and performance on phishing lures.
5 Discussion
Our large-scale field study provides empirical evidence for both the utility of the NIST Phish Scale and the limited effectiveness of current phishing training approaches. While we observed some statistically significant effects, the practical significance of these findings raises important questions about the return on investment for organizational security awareness programs.
5.1 Recommendations for Practice
Confronting Training Effectiveness Realities: Despite deploying well-structured training programs to over 12,000 employees, we found no statistically significant main effects of training on either click rates (p = 0.450) or report rates (p = 0.417). The small effect sizes we observed ($\eta^{2}$ $<$ .01 for all main effects) indicate that even statistically significant improvements translate to minimal operational impact. This aligns with recent large-scale evidence from Ho et al. [4] and challenges the optimistic assessments often promoted by cybersecurity training vendors.
The disconnect between vendor marketing claims and empirical evidence suggests that the substantial organizational resources typically invested in security awareness programs may not justify the modest improvements observed—even in best-case scenarios. Organizations must acknowledge these limitations to make informed investment decisions about cybersecurity resources.
Layered Defense: Our findings necessitate a fundamental shift in organizational phishing defense strategy. Rather than viewing training as a primary protection mechanism, organizations should treat security awareness programs as one component of a comprehensive layered defense strategy. The strong effect of phishing difficulty on user susceptibility (F(2, 12086) = 41.415, p $<$ 0.001) demonstrates that sophisticated attacks will likely succeed regardless of training efforts.
This evidence suggests several strategic reorientations: prioritizing passwordless authentication systems using passkeys and FIDO-compliant hardware tokens like YubiKeys, especially for administrative roles; implementing comprehensive password manager deployment to reduce credential reuse vulnerabilities; enhancing email security controls specifically designed to catch sophisticated attacks; and developing incident response procedures that assume successful phishing will occur despite training efforts.
The reporting behavior improvements we observed, while modest, indicate that interactive training components may have some value in encouraging security-conscious behavior. However, organizations must weigh the practical significance of 1-2 percentage point improvements against the substantial additional resources required for interactive training delivery.
Implementation Framework: Organizations seeking to optimize their phishing defense should implement the following evidence-based framework:
Technical Controls Priority. Invest primarily in email filtering, attachment analysis, URL protection, and multi-factor authentication systems. Our findings suggest these provide more reliable protection than user training against sophisticated attacks.
Targeted Training. Rather than broad-based programs, implement focused training for specific roles, departments, or threat scenarios where evidence suggests training may be more effective. Use standardized difficulty assessment frameworks like the NIST Phish Scale for meaningful evaluation.
Continuous Assessment. Regularly evaluate the cost-effectiveness of training programs against alternative security investments, using standardized metrics that capture both behavioral outcomes and resource allocation efficiency.
This approach recognizes the sobering reality that current training approaches provide only modest protection against evolving phishing threats. The goal is not to eliminate training but to right-size its role within a comprehensive security strategy informed by empirical evidence rather than vendor marketing claims.
5.2 NIST Phish Scale Validation and Practical Utility
Our study provides strong empirical validation for the NIST Phish Scale’s predictive utility in organizational contexts. The highly significant effect of lure difficulty on click-through rates (F(2, 12086) = 41.415, p $<$ 0.001) demonstrates that the scale effectively captures factors influencing user susceptibility to phishing attempts. The clear gradient in click rates across difficulty levels—from 7.0% for easy emails to 15.0% for hard emails—confirms that the scale’s two-dimensional framework meaningfully differentiates phishing sophistication.
This validation extends the early work of Barrientos et al. [39] by demonstrating the scale’s effectiveness in a large operational environment. Unlike previous smaller-scale validations, our study shows that the NIST Phish Scale maintains its predictive power across diverse organizational roles, departments, and user populations. This supports the scale’s adoption as a standardized framework for evaluating phishing simulation programs and benchmarking organizational vulnerability assessments.
However, the scale’s practical implementation revealed important considerations for practitioners. The process of systematically rating premise alignment across thousands of employees proved more complex than anticipated, requiring careful consideration of organizational communication patterns and role-specific email environments. Future implementations should account for these practical challenges when scaling the framework to enterprise environments.
5.3 Training Modality Effects and Interaction Patterns
While interactive training did not significantly outperform traditional lecture-based training in our primary analysis, the interaction patterns we observed provide nuanced insights into training effectiveness. The significant negative interactions between both training conditions and hard difficulty emails (Training Only × Hard: -9.67 percentage points, p = 0.002; Training + Exercise × Hard: -9.09 percentage points, p = 0.004) suggest that training may provide some protection against the most sophisticated phishing attempts. Conversely, the positive coefficients for training conditions with easier emails indicate that training may inadvertently increase vulnerability to simpler phishing attempts. This counter-intuitive finding aligns with concerns raised in prior work about potential negative effects of training programs [23]. One possible explanation is that training focuses employees’ attention on detecting sophisticated attacks while reducing their vigilance toward obviously fraudulent emails.
These patterns underscore the complexity of human factors in cybersecurity training. Rather than uniformly improving detection across all threat types, training appears to redistribute user attention and response patterns in ways that may create new vulnerabilities. This highlights the importance of comprehensive evaluation frameworks that examine training effects across varied attack sophistication levels.
5.4 The Evolving Threat Landscape and Future Challenges
Our findings must be interpreted within the context of rapidly evolving phishing threats. The emergence of large language models and generative AI technologies is enabling attackers to create increasingly sophisticated and personalized phishing attempts that may render current training approaches obsolete. The traditional approach of teaching users to identify visual cues and suspicious content becomes less effective when attackers can generate near-perfect replications of legitimate communications.
This technological evolution suggests that the cybersecurity community may need to fundamentally reconsider human-centered defense strategies. Rather than expecting users to reliably detect sophisticated attacks, future approaches might focus on cryptographic verification systems, zero-trust architectures, in-person interaction (supporting return-to-office/RTO mandates) or other technical mechanisms that reduce reliance on human judgment.
The broader threat model is also expanding beyond traditional email phishing to include SMS phishing (smishing), voice phishing (vishing), and social engineering attacks using deep fakes across multiple communication channels. Training programs that focus narrowly on email-based threats may provide limited protection against this diversified attack landscape.
5.5 Implications for Theory and Practice
Our findings have significant implications for both cybersecurity theory and organizational practice. The limited training effectiveness we observed suggests that current human-centered defense paradigms may be fundamentally inadequate for addressing sophisticated phishing threats. The modest effect sizes ($\eta^{2}$ $<$ 0.01 for all main effects) indicate that even statistically significant improvements may not translate to meaningful operational risk reduction.
The interaction patterns we identified where training appeared to increase vulnerability to easier emails while potentially providing some protection against harder emails highlight the complex and sometimes counterintuitive effects of security awareness interventions. These findings suggest that training programs may redistribute rather than uniformly reduce user vulnerability, creating new attack vectors while addressing others.
For practitioners, our results support a realistic reassessment of training investments within broader security strategies. Organizations should view awareness programs as one component of layered defense rather than primary protection mechanisms. The substantial resources typically devoted to comprehensive training initiatives might achieve better risk reduction when reallocated toward technical controls, incident response capabilities, or more targeted interventions [10].
5.6 The Future of Human-Centered Cybersecurity
Our findings emerge at a particularly critical juncture in the evolution of cybersecurity. With the rapid advancement of generative AI enabled by large language models and other technologies, attackers are enabled to create increasingly sophisticated and personalized phishing attempts that may render traditional cue-based detection obsolete. When attackers can automate the creation of personalized phishing lures and generate near-perfect replicas of legitimate communications, it becomes unreasonable to expect users to reliably identify deceptive content.
The fundamental challenge lies in the inherent complexity of human factors in resistance to phishing attacks. Since these exploits target individual vulnerabilities, personal and contextual factors play an enormous role in individual user susceptibility [19]. Even a single individual will vary dramatically throughout a day—users typically focus on content without paying attention to security, and are significantly impacted by cognitive constraints such as memory limitations and current cognitive load, all of which affect vigilance [43, 44]. Research has investigated variability in individual attentiveness based on factors like gender, age, and technical expertise, though the literature remains mixed [45, 23, 38]. Training represents only a single layer of defense against these multifaceted human vulnerabilities.
This technological evolution, combined with the complexity of human decision-making under cognitive load, suggests that the cybersecurity community must fundamentally reconsider where human-centered defenses lie in the multilayered approach organizations must take to cybersecurity. Future approaches may need to emphasize cryptographic verification systems, zero-trust architectures, and other technical mechanisms that can reduce reliance on the human firewall rather than attempting to enhance and upgrade the human firewall, which runs on hardware with a broad spectrum of variability.
The expanding threat landscape—encompassing SMS phishing, voice phishing, and multi-channel social engineering attacks—further challenges the viability of training-centric approaches focused primarily on email-based threats. As attack vectors diversify and sophistication increases, the cognitive burden placed on users to maintain vigilance across all communication channels may exceed reasonable human capabilities, particularly given the natural fluctuations in attention and decision-making quality that characterize normal human cognition. The role of human factors in resistance to phishing is complex.
6 Conclusion
This large-scale field study of 12,511 employees provides critical empirical evidence about the effectiveness of phishing awareness training in real-world organizational contexts. Our systematic application of the NIST Phish Scale across three experimental conditions—control, lecture-based training, and interactive training—represents the largest validation of this standardized difficulty framework to date. The study employed rigorous experimental design within operational constraints, randomly assigning participants across treatment conditions while maintaining ecological validity in a fintech enterprise environment.
Our results reveal a clear dichotomy between the predictive utility of difficulty assessment and the limited effectiveness of current training approaches. The NIST Phish Scale demonstrated strong predictive power, with phishing difficulty significantly affecting user susceptibility (F(2, 12086) = 41.415, p $<$ 0.001), showing click rates ranging from 7.0% for easy emails to 15.0% for hard emails. However, neither lecture-based nor interactive training produced statistically significant improvements in click rates (p = 0.450) or reporting behavior (p = 0.417), with effect sizes below 0.01 for all main training effects.
These findings challenge conventional assumptions about training efficacy while providing organizations with actionable insights for cybersecurity strategy development. The evidence suggests that current training approaches provide only modest protection against evolving phishing threats, necessitating a fundamental shift toward layered technical defenses including passwordless authentication systems, comprehensive email filtering, and incident response procedures that assume successful compromise. Rather than viewing training as a primary defense mechanism, organizations should integrate awareness programs as supplementary components within comprehensive security architectures that emphasize technical controls over human judgment for critical security decisions.

### v2

@@
-Anti-Phishing Training Does Not Work: A Large-Scale Empirical Assessment of Multi-Modal Training Grounded in the NIST Phish Scale
+Anti-Phishing Training (Still) Does Not Work: A Large-Scale Reproduction of Phishing Training Inefficacy Grounded in the NIST Phish Scale
 Abstract
-Social engineering attacks using email, commonly known as phishing, are a critical cybersecurity threat.
-Phishing attacks often lead to operational incidents and data breaches.
-As a result, many organizations allocate a substantial portion of their cybersecurity budgets to phishing awareness training, driven in part by compliance requirements.
-However, the effectiveness of this training remains in dispute.
-Empirical evidence of training (in)effectiveness is essential for evidence-based cybersecurity investment and policy development.
-Despite recent measurement studies, two critical gaps remain in the literature: (1) we lack a validated measure of phishing lure difficulty, and (2) there are few comparisons of different types of training in real-world business settings.
-To fill these gaps, we conducted a large-scale study ($N=12{,}511$) of phishing effectiveness at a US-based financial technology (“fintech”) firm.
-Our two-factor design compared the effect of treatments (lecture-based, interactive, and control groups) on subjects’ susceptibility to phishing lures of varying complexity (using the NIST Phish Scale).
-The NIST Phish Scale successfully predicted behavior (click rates: 7.0% easy to 15.0% hard emails, p $<$ 0.001), but training showed no significant main effects on clicks (p = 0.450) or reporting (p = 0.417).
-Effect sizes remained below 0.01, indicating little practical value in any of the phishing trainings we deployed.
-Our results add to the growing evidence that phishing training is ineffective, reinforcing the importance of phishing defense-in-depth and the merit of changes to processes and technology to reduce reliance on humans, as well as rebuking the training costs necessitated by regulatory requirements.
+Social engineering attacks delivered via email, commonly known as phishing, represent a persistent cybersecurity threat leading to significant organizational incidents and data breaches.
+Although many organizations train employees on phishing, often mandated by compliance requirements, the real-world effectiveness of this training remains debated.
+To contribute to evidence-based cybersecurity policy, we conducted a large-scale reproduction study (N=12,511N=12{,}511) at a US-based financial technology firm.
+Our experimental design refined prior work by comparing training modalities in operational environments, validating NIST’s standardized phishing difficulty measurement, and introducing novel organizational-level temporal resilience metrics.
+Echoing prior work, training interventions showed no significant main effects on click rates (pp=0.450) or reporting rates (pp=0.417), with negligible effect sizes.
+However, we found that the NIST Phish Scale predicted user behavior, with click rates increasing from 7.0% for easy lures to 15.0% for hard lures.
+Our organizational-level resilience result was mixed: 36-55% of campaigns achieved “innoculation” patterns where reports preceded clicks, but training did not significantly improve organizational-level temporal protection.
+In summary, our results confirm the ineffectiveness of current phishing training approaches while offering a refined study design for future work.
 1 Introduction
-Phishing attacks critically impact organizations through operational incidents, data breaches, and financial losses.
-They constitute the primary attack vector to breach organizations as highlighted by industry studies [1] and the US Federal Bureau of Investigation (FBI) [2, 1].
+Phishing attacks harm organizations through operational incidents, data breaches, and financial losses [4].
+Industry studies and the US FBI agree that phishing is the primary attack vector to breach organizations [42].
+As an example, in 2022–2023 campaigns attributed to the Scattered Spider cluster (aka 0ktapus/Muddled Libra) compromised 9,931 accounts in over 130 organizations [43, 23].
+This campaign cost MGM Resorts alone an estimated \$100 million [79].
 To prevent this class of attacks, organizations employ technical controls (e.g., spam filters, attachment analyzers) complemented by cybersecurity awareness training [58, 48].
-Although these anti-phishing trainings are costly, their effectiveness is in question.
-Organizational security awareness programs often focus on compliance metrics rather than meaningful behavioral change [5].
-Meanwhile, recent large-scale evidence suggests that current approaches have limited impact [4].
-To better evaluate the effectiveness of anti-phishing training, two questions arise.
-First, are some kinds of training more effective than others?
-Second, how does their effectiveness vary with the difficulty of the phishing lure?
-In this paper, we report the results of a measurement study to address these questions.
-We situate the study in an operational environment [6], set at a US-based international fintech firm with 12,511 employees across multiple business units.
-We compared three experimental conditions: a control group (no training), traditional lecture-based training, and lectures augmented with interactive phishing exercises.
-Subjects in each condition received lures of varying complexity, which we assigned difficulty levels using the NIST Phish Scale.
-This study design enables rigorous evaluation of training effectiveness across standardized difficulty levels on a large population, improving on previous works that had either been small-scale or lacking in systematic difficulty assessment [7, 8].
-Our findings were unfavorable for the conventional anti-phishing trainings available from cybersecurity training vendors.
-Neither mode of training produced statistically significant improvements in phish click reduction (average: 10.4%) or phish reporting behavior (average: 9.6%).
-However, the NIST Phish Scale did predict user behavior, with phishing lure difficulty having a highly significant effect on click-through rates (F(2, 12086) = 41.415, p $<$ 0.001, $\eta^{2}$ = 0.007).
-Thus our study shows that, while the NIST Phish Scale provides a means of measuring the difficulty of different lures, the trainings we applied did not help subjects perform better on those lures.
+Despite widespread organizational investment in anti-phishing training, mounting empirical evidence (and employee sentiment [78]) questions its effectiveness.
+Large-scale studies across multiple sectors demonstrate minimal training impact [48, 53, 54].
+Most pertinently, Lain et al.
+showed that phishing training did not improve security outcomes across 19,000+ employees over multiple years [53, 54].
+These findings align with theoretical work by Volkamer et al., who demonstrate fundamental limitations in human processing of security indicators [96].
+This growing body of evidence suggests that phishing training is ineffective, but we believe that further reproduction of the result is needed to inform cybersecurity policy and defense strategies [67, 89].
+[^1] Our reproduction effort differs from an exact replication, which would involve repeating identical procedures.
+Instead we focus on testing the generalizability of the phenomenon of training ineffectiveness, varying the organizational context and the training approach.
+We target two specific knowledge gaps.
+First, existing studies lack standardized measures of phishing difficulty, so we do not know whether training effectiveness varies by attack sophistication.
+Second, most studies examine single training modalities rather than comparing different approaches systematically, so we do not know if some training designs are more effective than others.
+To address these limitations, we conducted a large-scale reproduction study (N=12,511N=12{,}511) at a US-based financial technology firm.
+Our study design has three distinctive elements: (1) mandatory pre-phishing training using two distinct training modalities (lecture-based vs.
+interactive); (2) systematic difficulty assessment using the NIST Phish Scale [86, 7]; and (3) novel organizational temporal resilience measurement through the Organizational Inoculation Index, which measures the effectiveness of phishing training from an organizational perspective rather than an individual one.
+This design enables rigorous evaluation of training effectiveness across standardized difficulty levels while providing new insights into organizational cybersecurity resilience patterns that complement individual-focused assessments.
+Our reproduction study confirms the ineffectiveness of anti-phishing training across multiple dimensions, but adds nuance to prior results.
+Consistent with prior work, neither training modality produced statistically significant improvements in phish click reduction (pp=0.450) or phish reporting behavior (pp=0.417), with negligible effect sizes indicating minimal practical value for organizational security investment.
+However, the NIST Phish Scale successfully predicted user behavior (F(2, 12086) = 41.415, p < 0.001, $\eta^{2}$ = 0.007), providing the first large-scale validation of this standardized difficulty measure.
+[^2] While the effect size is small in statistical terms due to low base rates of phishing engagement, the practical implications are significant: click rates doubled from 7.0% to 15.0% as lure difficulty increased.
+Beyond that, our Organizational Inoculation Index revealed substantial organizational resilience that operates independently of individual training effectiveness.
+Across all experimental conditions, 36.4% to 54.5% of template-group combinations demonstrated “inoculation” patterns where threat reports (from a security-conscious employee) preceded threat interaction by another employee.
 We summarize our main contributions below:
   1. 1.
-    Large-Scale Validation of the NIST Phish Scale.
-We provide the most comprehensive empirical validation of the NIST Phish Scale to date, demonstrating its predictive utility across 12,511 participants in an operational environment, showing strong correlation between difficulty ratings and user susceptibility.
+    Reproduction of Training Ineffectiveness Across Multiple Modalities.
+Our results reproduce the core finding of training ineffectiveness while demonstrating this phenomenon’s generalizability across training designs.
   2. 2.
-    Realistic Assessment of Training Effectiveness.
-Our results challenge optimistic assumptions about the effectiveness of phishing training.
-We found no statistically significant impact of training on either click rates or reporting behavior.
-Although these findings align with recent large-scale studies from the healthcare [4] and banking sectors [9], our study addresses explicit calls from the research community for large-scale ecologically valid studies combining standardized assessments of difficulty inside of an operational context [10, 6].
-Significance: Our findings have important implications for researchers, organizations, and regulators.
-For the research community, we provide empirical assessment of training effectiveness using rigorous methodology and standardized difficulty measures.
-For organizations, our results suggest the need for realistic expectations about training outcomes and highlight the importance of complementing awareness programs with robust technical controls rather than relying primarily on human-centered defenses.
-Our validation of the NIST Phish Scale’s predictive utility enables organizations to use this tool for benchmarking the difficulty of their phishing simulations.
-For regulators, our work contributes to the growing body of evidence that organizational cybersecurity strategy should emphasize technical defenses [1] rather than placing the burden on humans — requiring anti-phishing training is costly and has minimal effect.
+    First Large-Scale Validation of the NIST Phish Scale.
+We evaluate the NIST Phish Scale across 12,511 participants.
+The Phish Scale is a viable measure for assessing phishing difficulty, with predictive utility.
+  3. 3.
+    Introduction of the Organizational Inoculation Index.
+We introduce and validate a novel metric for organizational temporal security.
+We report substantial organizational resilience (36-55% inoculation rates), which operates independent of individual training effectiveness and provides new insights into collective cybersecurity behaviors.
+Significance: Cybersecurity practice and policy must be evidence-based.
+For regulators, our work contributes to growing evidence that compliance-driven training requirements may not achieve intended security outcomes.
+For organizations, our results suggest realistic (low) expectations about training outcomes and highlight the importance of technical controls rather than human-centered defenses.
+However, we also demonstrated that collective security behaviors create measurable protective effects, emphasizing the value of feedback loops between user reports and real-time threat response.
+For researchers, we validate NIST’s standardized difficulty measure and confirm the reproducibility of training ineffectiveness findings across organizational contexts.
 2 Background and Related Work
-In a phishing attack, attackers send messages (phishing lures) to targets whom they wish to persuade to act against their own interests.
+Figure 1: Phishing defense architecture, mapping the email pathway from authentication through post-delivery response.
+Technical controls create multiple interception points prior to human interaction, and human reports provide a feedback opportunity for organizational-level threat mitigation.
+See Appendix 2 for the full control pathway in the studied organization.
+We review phishing attacks and defenses (§2.1), and then discuss prior empirical studies on phishing training (§2.2).
+2.1 Phishing Attacks and Defenses
+Phishing attacks pose substantial risks to organizations, who erect both technical and human defenses against them.
+2.1.1 Phishing Attacks
+In a phishing attack, attackers send messages (phishing lures) to targets whom they wish to persuade to act against their own interests, and the interests of their organizations.
 Attackers seek results such as sensitive information, account credentials, or the ability to control the victim’s machine [4, 34].
-Modern phishing campaigns target individuals and organizations across multiple attack vectors, applying conventional email phishing as well as SMS-based attacks (“smishing”) and voice-based deception (“vishing”) [12, 13].
-Recent attacks have begun to incorporate Generative AI tools to produce tailored and realistic messages.
-The severity, persistence, and increasing sophistication of these attacks has established phishing detection and prevention as a critical focus within both academic cybersecurity research and industry practice [14, 6].
-In this section, we summarize current defenses against phishing (Appendix 2.1), discuss metrics for evaluating phishing effectiveness (Appendix 2.2), and review prior empirical studies (Appendix 2.3) and knowledge gaps (Appendix 2.4).
-2.1 Phishing Defenses
-Phishing poses substantial risks to organizations, who therefore erect multiple layers of defenses against these attacks. We distinguish automated defenses from human-focused ones.
+Traditionally, phishing attacks are divided into two classes [25]:
+Mass Phishing Campaigns: These use malicious emails sent to thousands or millions of recipients, with common themes such as banking, shipping, and software updates [4].
+Individual success rates are low (1-3%) [3], but the massive scale makes them profitable.
+Spear Phishing and Targeted Attacks: Spear phishing targets specific individuals or organizations with personalized content from open-source intelligence (OSINT) and reconnaissance.
+Attackers mine sources such as social media to craft messages that reference real relationships and events [13].
+These tailored campaigns have higher success rates (10–20%) [14].
+Successful spear phishing attacks include the 2016 Democratic National Committee breach and widespread business email compromise (BEC) schemes [4].
+Modern Phishing: Contemporary phishing operations leverage infrastructure including compromised (“trusted”) websites and domain generation algorithms to maintain persistence and evade takedown efforts [3].
+In addition, Generative AI is enabling both mass phishing and spear phishing through targeted and fluent emails and other media [4, 80].
+LLM-generated phishing content now achieves effectiveness comparable to human experts in organizational settings [9].
 2.1.2 Automated Defenses
-Organizations employ a range of technological defenses to mitigate phishing risks.
-These include spam filters, attachment analyzers, and URL checkers [15, 14, 1].
-More recently, machine learning approaches have been applied to detect phishing content.
-Long Short-Term Memory (LSTM) models and optimized Random Forest algorithms [16] have demonstrated high accuracy in detecting phishing attempts.
-As with everything in cybersecurity, phishing is a cat-and-mouse game.
-Adversaries have begun to leverage artificial intelligence to create more convincing lures.
-Large language models (LLMs) can generate personalized and context-aware phishing emails that are difficult for traditional filters to detect [3].
-Techniques like typosquatting, domain spoofing, and watering hole attacks exacerbate this issue [1, 14].
-2.1.2 Human-Focused Defenses: Training and Awareness
-Defending against phishing is understood not as solely a technical problem, but rather as a socio-technical one [6, 13].
-Spear phishing and emerging phishing attacks present sophisticated threats that overcome automated defenses [17, 18], so human awareness of phishing remains necessary to successful defense [19, 20].
-There are many approaches to training humans to be resilient to phishing attacks.
+Figure˜1 illustrates how these attacks flow through an organization’s email system.
+Many technical defenses (blue boxes) combat attackers’ behavior, such as SPF, DKIM, and DMARC to verify that mail comes unaltered from authorized infrastructure [51, 22, 52]; ARC for preserving authentication results across forwarding [5]; BIMI for verified logos [10, 12]; and secure email gateways (SEG) for content and header inspection and URL rewriting for time-of-click inspection [1].
+After delivery, organizations employ continuous monitoring and feedback loops to adapt controls to evolving campaigns.
+Organizations blend threat-intelligence updates with AI/ML-driven anomaly scoring and impersonation detection to surface novel lures [58].
+For example, security operations center (SOC) analysts triage user-reported messages, promote high-confidence indicators back into blocklists, and tune SEG policies based on observed false positives/negatives.
+Because adversaries adapt, detection is a dynamic process [4, 3].
+See Appendix 2 for the details of the studied organization’s email controls and how user phishing reports create feedback loops that strengthen automated defenses.
+2.1.3 Phishing Training as a Defense
+Because phishing attacks compromise these technical defenses [14, 13], the human recipients play an important role in phishing defense [30, 15, 2, 28].
+There are many approaches to training humans about phishing attacks.
 Methods range from lecture-style training involving passive videos and readings [56] to simulation feedback providing immediate feedback after interaction with phishing simulations [19, 18].
 Interactive training uses gamification or scenario-based learning to engage participants [11, 46], while just-in-time training alerts users in real time during risky behavior [36].
-Anti-phishing trainings are frequently included within the cybersecurity awareness trainings ubiquitous in organizations.
-Legal mandates drive the widespread adoption of such training programs, with HIPAA requiring security awareness training for healthcare organizations [27], GDPR encouraging staff training on cybersecurity practices [28], and PCI DSS mandating training for those handling cardholder data [29].
-Industry standards and frameworks further reinforce these requirements, with ISO 27001 calling for human resource security controls including awareness programs [30] and the NIST Cybersecurity Framework advising phishing simulations under the “Identify” function [31].
-2.2 Measuring Effectiveness of Phishing Training
-The effectiveness of phishing training can be assessed based on how resilient a target becomes to phishing lures.
-Appendix 2.2.1 describes how lures can be categorized by their complexity, and then Appendix 2.2.2 gives the standard metrics for target resilience.
-2.2.1 Measuring Training Lure Difficulty
-TABLE I: NIST Phish Scale for Detection Difficulty [32, 33].
-Cues \ Alignment Weak Medium Strong
-Few Medium Hard Hard
-Some Easy-to-Medium Medium Hard
-Many Easy Medium Medium
-TABLE II: Summary of Prior Experiments on Phishing and Training Interventions.
+Anti-phishing trainings are often included in the cybersecurity awareness trainings ubiquitous in organizations.
+Legal mandates drive this practice, e.g., HIPAA requiring security awareness training for healthcare organizations [74].
+GDPR and PCI DSS encourage or mandate similar training [76, 21].
+Industry standards and frameworks further reinforce these requirements [75], e.g., ISO 27001 calling for human resource security controls including awareness programs [35].
+2.2 Empirical Studies of Phishing Training
+Table 1: Summary of Prior Experiments on Phishing and Training Interventions.
 Studies are organized by research context (Lab vs.
@@
 Hypotheses: How these works informed the hypotheses in the present study (cf.
-Table III).
+Table˜4).
 Work Context Population Sample size Complexity Measure Compared Training Approaches? Hypotheses
 [82] (2024) Lab University students 96 H3
-[34] (2024) Lab University students 117  (NIST Phish Scale) H4
+[16] (2024) Lab University students 117 (Phish Scale) H4
+[18] (2014) Real-world Employees 1,359 H2, H4
+[19] (2017) Real-world University students 150 H2, H3
 [90] (2022) Real-world Employees 119 H2
-[22] (2017) Real-world University students 150 H2, H3
-[23] (2014) Real-world Employees 1,359 H2, H4
+[53] (2022) Real-world Employees 19,000+ H2
+[47] (2023) Real-world Finance employees 5,000 H2
+[54] (2024) Real-world Employees 4,554 H2, H3
 [29] (2024) Real-world Banking employees “Thousands” H2
-[35] (2023) Real-world Finance employees 5,000 H2
 [48] (2025) Real-world Healthcare employees 19,500+ H2
-Ours Real-world Fintech org. 12,511  (NIST Phish Scale)
-To assess phishing lure difficulty, the NIST Phish Scale categorizes them along two dimensions as depicted in Table I.
-Phishing Cues represent observable errors or inconsistencies in the email that might alert users to its fraudulent nature, such as spelling mistakes, suspicious URLs, or formatting irregularities [34].
-Premise Alignment measures the relevance and familiarity of the email’s subject matter to the recipient’s typical email environment and organizational context [36].
-A high alignment, messages that closely resemble emails users receive daily, combined with few obvious cues yields the most deceptive and challenging-to-detect phishing emails [17].
-The scale allows organizations to benchmark lure complexity and contextualize simulation outcomes, though it requires sufficiently large sample sizes for reliable use and consistent application [37].
-2.2.2 Training Outcome Metrics
-The effectiveness of anti-phishing training is commonly measured in three ways:
-  • Open Rate captures the percentage of emails opened, though this measurement can be deceptive since different email clients and security settings can affect whether an email is recorded as opened [36] we therefore largely ignore this measure.
-  • Click-Through Rate (CTR) measures the percentage of recipients who click on malicious links, serving as a primary indicator of phishing susceptibility [22, 23].
-  • Reporting Rate tracks phishing incidents reported to security teams, representing users’ proactive security behaviors and ability to recognize suspicious content [18, 38].
-2.3 Empirical Studies of Anti-Phishing Training
-For reviews of knowledge about susceptibility to phishing and about anti-phishing training, we refer the reader to recent works by Marshall et al.
+Ours Real-world Fintech org. 12,511 (Phish Scale)
+Our review and analysis of prior work is summarized in Table˜1.
+We contribute to the body of knowledge on phishing training by conducting another large-scale study situated in a real-world context (table columns 2-4), incorporating a complexity measure (column 5), and a comparison of multiple training approaches (column 6).
+Our study tests hypotheses derived from prior work (final column).
+The remainder of this section describes these contributions.
+For general surveys of knowledge about human factors in phishing training, we refer the reader to Marshall et al.
 [56] and Franz et al.
 [37].
-This section focuses on the aspect of the literature to which this study contributes.
-We discuss the settings of prior work (Appendix 2.3.1), the assessments of different trainings (Appendix 2.3.2), and measurement and evaluation methods (Appendix 2.3.3).
-Figure 1: Illustration of our methodology.
-Our two-factor design randomly assigns different trainings to different subjects, and then randomly sends phishing lures of varying complexity to the subjects.
-We compared between-subject performance using lure click rate and lure report rate in order to test four hypotheses derived from prior work.
-2.3.1 Scale and Context of Prior Work
-Many phishing training studies have been conducted in controlled laboratory settings, where researchers benefit from a high degree of experimental control.
+Lab vs.
+Real-world Studies: Many early phishing training studies were conducted in controlled laboratory settings, where researchers benefit from a high degree of experimental control.
 Studies of this sort have provided numerous insights, including that active learning and immediate feedback can improve short-term detection rates, and have explored how individual differences affect training outcomes [90, 56].
@@
 In laboratory environments, participant samples are typically small and consist of homogeneous groups such as academics and students.
-Due to this, and the artificial nature of lab tasks, these studies may not accurately reflect the complexities and distractions of real-world environments.
-This raises questions about the generalizability of laboratory findings to operational contexts, where additional factors can alter training effectiveness.
 In response, a growing number of studies have begun examining phishing training in real-world organizational environments.
-Doing et al.
-[9] analyzed phishing reporting behaviors across thousands of bank employees, while Ho et al.
-[4] conducted a comprehensive evaluation with over 19,500 healthcare employees across 10 simulated phishing campaigns, reporting minimal effectiveness of both annual cybersecurity awareness training and embedded anti-phishing training.
-To date, however, these larger studies have lacked a standardized method for assessing the difficulty of the phishing lures in question [4, 9], or have focused primarily on observational analysis rather than controlled training interventions with standardized difficulty measures [9].
-Our study contributes to the evidence on real-world human phishing resilience, with data from approximately 12,000 subjects in a fintech enterprise, while maintaining experimental control and measuring lure difficulty with the NIST Phish Scale framework.
-This approach allows us to examine whether the modest training effects we observe align with Ho et al.’s findings, while providing additional insights into how training effectiveness varies by phishing sophistication.
-2.3.2 Training Intervention Approaches
-Franz et al.
-[10] provide a comprehensive taxonomy identifying five categories of phishing interventions: training, awareness, education, notifications, and design interventions.
-Our study contributes empirical data within their training category, specifically comparing lecture-based delivery against interactive modalities.
-While Schöni et al.
-[8] demonstrated effectiveness of personalized training approaches, and our work examines scalable training modalities that can be deployed across demographically diverse cohorts without individualization, recent large-scale studies suggest that the practical benefits of such training interventions may be more limited than previously assumed [4].
-Our study builds on this foundation by applying the NIST Phish Scale to measure lure difficulty in line with current guidance, operationalizing the scale at unprecedented scale, and measuring outcomes (clicking, reporting) post-intervention, allowing for practical comparison across conditions and lure complexity.
-2.3.3 Measurement and Evaluation Methods
-A critical gap in prior work has been the lack of standardized difficulty assessment for phishing simulations.
+For example, Doing et al.
+[29] analyzed phishing behaviors in bank employees, and Ho et al.
+[48] studied healthcare employees.
+Lain et al.
+has also conducted large-scale real-world measurements of training effectiveness [53, 54] Our study echoes some of these prior results in a different organizational context.
+Controlling for Lure Complexity: These studies consistently report little effectiveness of anti-phishing training.
+These works have not, however, incorporated a standardized method for assessing the difficulty of the phishing lures they use [48, 29].
+We contribute the first real-world measurement of the effect of lure complexity on subjects’ interaction behaviors, based in NIST’s Phish Scale.
+To assess phishing lure complexity, the NIST Phish Scale categorizes them along two dimensions (Table˜2).
+Phishing Cues represent observable errors or inconsistencies in the email that might alert users to its fraudulent nature, such as spelling mistakes, suspicious URLs, or formatting irregularities [16].
+Premise Alignment measures the relevance and familiarity of the email’s subject matter to the recipient’s typical email environment and organizational context [6].
+Higher alignment messages with fewer cues are the most deceptive [14].
+The premise of Phish Scale is to allow organizations to benchmark lure complexity and contextualize simulation outcomes [17].
 Barrientos et al.
 [8] conducted early validation work on the NIST Phish Scale, while Dawkins and Jacobs [27] provided the official implementation guidelines.
-However, no studies have operationalized the scale at enterprise level nor examined how difficulty interacts with training effectiveness.
-We applied the scale across varied training modalities and measuring multiple outcome measures (clicking, reporting) post-intervention.
-2.4 Research Gaps
-Table II summarizes the preceding analysis as two gaps.
-First, prior work has primarily focused on individual-level factors influencing phishing susceptibility.
-The interaction between training modality and phishing difficulty — a key question for optimizing organizational defenses — has received little attention.
-Second, despite widespread adoption of phishing training, few studies rigorously assess outcomes against standardized difficulty measures or operate at sufficient scale for robust inference.
+Ours is the first study to validate the Phish Scale’s premise at scale.
+Training Approaches: Franz et al.
+[37] taxonomize five categories of phishing interventions: training, awareness, education, notifications, and design interventions.
+Most studies have focused on training, which can be divided into lecture-style annual training and embedded training (e.g., through ongoing phishing simulations sent by the organization’s security team).
+Like some prior studies, our study performs a controlled experiment to compare the effectiveness of two forms of phishing training.
+Table 2: NIST Phish Scale for Detection Difficulty [87, 27].
+Cues \ Alignment Weak Medium Strong
+Few Medium Hard Hard
+Some Easy-to-Medium Medium Hard
+Many Easy Medium Medium
 This study addresses those gaps by applying the NIST Phish Scale [27] to evaluate training effectiveness across varied phishing difficulty levels in an enterprise deployment with over 12,000 participants.
@@
 3 Methodology
-In this section, we describe the methodology of our measurement study, including our hypotheses, experimental design, participant demographics, training modalities, and phishing simulations.
-An illustration is given in Figure 1.
-Our two-factor study design controls for the effect of training modality and task complexity.
-We make the first real-world application of the NIST Phish Scale framework [33], allowing us to evaluate the effectiveness of different training modalities across standardized difficulty levels in an operational environment.
-TABLE III: Summary of study hypotheses, their theoretical grounding, and relation to prior work.
-The hypotheses address the effects of phishing email difficulty, training, and training modality on detection and reporting behaviors.
-Hypothesis Theoretical and Empirical Basis Relation to Prior Work
-H1 (Phish Scale Effect): Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will exhibit higher click-through rates than those with lower difficulty.
-Grounded in the foundational NIST Phish Scale framework [33] and validation studies [39].
-Supported by phishing cue research [34] and systematic email analysis [40].
-The NIST Phish Scale has been validated in lab settings [39].
-We provide the first large-scale enterprise validation.
-H2 (Training Impact — General): Training will increase the detection and reporting of phishing emails.
-Evidence of training efficacy, reduced click rates [22], improved detection accuracy [7], and short-term benefits [21].
-Disputed by a recent study [4].
-Tests whether training effects from small-scale studies [22, 7] replicate at scale, addressing the call for large-sample validation [10].
-H3 (Interactive Training Effect — Reporting): Interactive components (Trained+Exercise) will lead to significantly higher reporting rates than lecture-only training (Treatment A).
-From Franz et al.’s modality taxonomy [10], and evidence of the effectiveness of active and feedback-rich training [22, 21, 8].
-Builds on academic findings of interactive training benefits [22, 21].
-We provide the first test at organizational scale.
-H4 (Interaction Effect): There will be a significant interaction between training modality and phishing difficulty, with interactive training being more beneficial for easier lures.
-Emerges from cognitive load and decision strategy research [19], difficulty-aware detection patterns [40], and findings on human limitations [6].
-Addresses the gap identified by Franz et al.
-[10] regarding interaction effects between training and phishing complexity, largely unexplored empirically.
-3.1 Hypotheses
-Based on prior work on phishing training efficacy [22, 23, 5, 7, 10, 4] and documented differences in risk perception [6, 9], we formulated four hypotheses to test in this work.
-Table III states each hypothesis, its basis, and its relation to prior work.
-To summarize, H1 represents the foundational assertion of the NIST Phish Scale framework [33]; H2 and H3 capture the intended effects of training interventions, with H3 specifically testing whether interactive modalities provide measurable benefits over passive learning approaches; and H4 reflects our expectation that training effectiveness may diminish for highly sophisticated phishing attempts, as advanced phishing emails may exceed individuals’ detection capabilities regardless of training modality.
-With respect to hypothesis novelty and prior testing, H1 has been tested in small academic settings [39] but never at enterprise scale; H2 and H3 build on established training effectiveness research [22, 7, 21] but test these findings in a large-scale operational context using standardized difficulty measures; and H4 is novel, addressing the research gap identified by Franz et al.
-[10] regarding interactions between training modality and phishing complexity, which has received minimal empirical attention in prior work.
+To summarize the preceding analysis: phishing training is widely practiced, though recent evidence sheds doubt on its utility.
+Few studies have rigorously assessed outcomes against standardized difficulty measures at sufficient scale for robust inference.
+We provide a reproduction of this observation, tested through hypotheses derived from prior works.
+In this section, we describe our methodology: study design, subjects, hypotheses, metrics, and analysis methods. For ethics considerations, see the associated appendix.
 3.1 Study Design
-To test the hypotheses, we employed a between-subjects experimental design [35].
+Figure 2: Illustration of our methodology.
+Our two-factor design randomly assigns different trainings to different subjects, and then randomly sends phishing lures of varying complexity to the subjects.
+We compared between-subject performance using a mix of standard and novel metrics to test our five hypotheses.
+Our study design is illustrated in Figure˜2.
+We employed a two-factor between-subjects experimental design [47].
 Subjects were randomly assigned to one of three experimental conditions using our organization’s training management system.
-Each condition reflected the phishing training they received (Appendix 3.3).
-Subjects were recruited (Appendix 3.4), trained, and later received simulated phishing emails (Appendix 3.5).
-We compared subjects’ performance in handling these phishing emails between the treatments using several standard metrics (Appendix 3.6).
-3.3 Training Modalities
-For the purposes of this research, the principal researcher collaborated with a cybersecurity training vendor who produces comprehensive security awareness training materials, phishing simulation exercises designed to augment those materials, and implementation assistance for organizational deployment.
-This partnership enabled access to professionally developed training content and established phishing templates while maintaining the experimental rigor necessary for our evaluation.
-The vendor offers three distinct training modalities, though we focused our evaluation on two of these approaches.
-The primary modality for cybersecurity awareness training consists of lecture-based videos that discuss the dangers of various social engineering attacks, including phishing recognition tactics, social engineering techniques, and organizational security protocols.
-Our vendor, along with many others in the industry, supplements this foundational training with interactive phishing exercises that present users with samples from a library of approximately 20 phishing templates.
-These exercises guide users through identifying suspicious elements and security deficiencies in emails that roughly correspond to the cues assessed by the NIST Phish Scale.
-The vendor also provides a third modality referred to as “in-place training.” This approach occurs during phishing simulations when users click on malicious links, instead of being directed to a malicious website, they are immediately redirected to a training webpage that explains why the email they clicked should have been detected as fraudulent.
-However, decades of research into the effectiveness of embedded training environments [23] and recent comprehensive research by Ho et al.
-[4] convinced us that monitoring the results of in-place training was not necessary for our research objectives.
-Ho’s large-scale study across 19,500 healthcare employees found minimal effectiveness of embedded training approaches, providing thorough and convincing evidence that this modality offers limited practical value compared to the lecture-based and interactive approaches we chose to evaluate.
-We employed two distinct training approaches in collaboration with our cybersecurity training vendor, building on established taxonomies of phishing interventions [10].
-Both modalities fall within Franz et al.’s “training” category [10], specifically comparing passive lecture-based delivery Figure 2against interactive approaches with embedded learning elements.
-Treatment A: Lecture-Based Training: The lecture-style training consisted of fifteen brief instructional videos, each lasting 1-2 minutes, covering core cybersecurity awareness topics including phishing recognition tactics, social engineering techniques, and organizational security protocols [22].
-The videos specifically addressed phishing identification strategies, emphasizing visual cues such as suspicious URLs, sender inconsistencies, and urgency tactics commonly employed by attackers.
-Following the video content, participants completed a knowledge assessment quiz reviewing key concepts from the training materials.
-This approach represents the traditional passive learning modality widely deployed in organizational cybersecurity training programs [23].
-Figure 2: Traditional Training: Videos and Quiz
-Treatment B: Interactive Training (Training + Exercise): The interactive group received the identical lecture-style training described above, plus an additional hands-on component consisting of simulated phishing exercises with immediate feedback.
-During these exercises, participants were presented with sample emails and asked to identify whether each message was legitimate or malicious.
-The system provided real-time feedback explaining the reasoning behind correct identification, highlighting specific phishing indicators, and reinforcing appropriate response behaviors such as using the “Report Phish” button in the email client [22].
-This interactive component was designed to provide experiential learning opportunities that complement the passive video instruction, following research suggesting that active engagement enhances retention and behavioral change in cybersecurity training contexts [33].
-Control Group: The control group received no cybersecurity training prior to the phishing simulation phase.
+Each condition reflected the phishing training they received.
+Subjects were randomly selected, trained, and later received simulated phishing emails of varying complexity.
+We quantitatively compared subjects’ performance in handling these phishing emails between the treatments.
+Factor 1—Training: The research team collaborated with a cybersecurity training vendor who produces comprehensive security awareness training materials, phishing simulation exercises designed to augment those materials, and implementation assistance for organizational deployment.
+This partnership enabled access to professional training content and established phishing templates.
+The vendor offers two distinct training approaches for phishing.
+The first approach, Treatment A, consists of lecture-based videos on the dangers of various social engineering attacks, including phishing recognition tactics, social engineering techniques, and organizational security protocols.
+Figure˜3 gives a screenshot.
+Comprehension is assessed through a quiz.
+The second approach, Treatment B, consists of those videos followed by a series of interactive phishing exercises that present users with samples from a library of approximately 20 phishing templates.
+These exercises guide users through identifying suspicious elements and security deficiencies in emails that roughly correspond to the cues assessed by the NIST Phish Scale, and reinforces appropriate response behaviors such as the “Report Phish” button in the email client [19].
+Finally, we included a Control Group whose members received no cybersecurity training prior to the phishing simulation phase.
 This group was included to establish baseline susceptibility measurements.
-Control group participants received the same training materials after completion of the study to ensure organizational compliance with cybersecurity awareness requirements.
-Organizational constraints compelled us to use a relatively small control group, as ACME Corp.
-required that the majority of employees receive some form of cybersecurity training for compliance purposes.
-3.4 Subject Recruitment
+[^3] Control group participants received the same training materials after completion of the study to ensure organizational compliance with all cybersecurity awareness requirements.
+The study timing dictated a small control group, as ACME Corp.
+required that the majority of employees receive annual cybersecurity training for compliance purposes.
+Figure 3: Training for Treatment A: Videos and Quiz
+Factor 2—Phishing Lure Complexity: After training, all subjects received one simulated phishing email, of varying complexity.
+To establish controlled and consistent evaluation of participant responses, we systematically assessed and categorized the difficulty of each simulated phishing email using a structured expert review process informed by the NIST Phish Scale framework [27].
+Phishing templates were selected from the vendor’s library.
+We evaluated these templates using the official NIST Phish Scale User Guide methodology to ensure accurate difficulty classification.
+For the Phishing Cues dimension, raters assessed observable indicators that might alert users to fraudulent content, including spelling errors, suspicious URLs, formatting inconsistencies, and sender authenticity markers [16].
+For the Premise Alignment dimension, raters evaluated how well each email’s subject matter, sender, and content aligned with typical organizational communications that ACME Corp.
+employees receive in their regular workflow [2].
+The assessment team comprised three members of ACME Corp.’s IT security staff, each experienced in identifying real-world phishing attacks.
+Following Downs et al.’s advice, two team members independently evaluated each phishing email template [30].
+Any disagreements regarding difficulty classification were discussed and resolved, incorporating a third rater as needed.
+This collaborative validation process ensured that our phishing simulation represented a controlled range of difficulty levels essential for evaluating training effectiveness across standardized complexity measures [8].
+Our final simulation campaign used nineteen phishing templates that we divided into Easy (high cues/low alignment — 5721 emails), Medium (some cues, some alignment — 2279 emails) and Hard (few cues/high alignment — 4511 emails).
+Examples are given in Figure˜4.
+A full summary is given in Table˜6 in the Appendix.
+(a) Easy: low premise, high cues.
+(b) Hard: high premise, low cues.
+Figure 4: Two Example Phishing Templates.
+Simulation Deployment: Phishing simulations were distributed among participants within a three-month window after training completion.
+Each participant received one phishing email.
+Templates were randomly assigned to ensure balanced distribution across training conditions.
+3.2 Subjects
 Study Context: The subjects were employees at a U.S.-based international financial technology (“fintech”) firm.
 Participants represented a diverse cross-section of the organization, including employees from finance, technology, operations, customer service, and administrative functions.
-All participants were full-time employees with active email accounts who regularly use organizational communication systems and speak English.
+All participants were full-time employees with active email accounts who regularly use organizational communication systems.
 We excluded participants who: (1) had inactive email accounts during the simulation period, (2) were on extended leave during the study period.
@@
 This training was mandatory due to the organization’s need for regulatory compliance (cf.
-Appendix 2.1.2).
+Appendix˜2.1.3).
 Data were collected from Dec.
@@
 Training completion was tracked and verified.
-Randomization: Subjects wre randomly assigned to each treatment.
+Randomization: Subjects were randomly assigned to each treatment.
 To mitigate bias by organizational role or function, we used stratified sampling [91] across departments to ensure representative distributions in each treatment group.
 Completion Rates: Our data came from the 12,511 participants who completed the full experimental protocol.
-Table IV shows the completion rates by treatment type.
+Table˜3 shows the completion rates by treatment type.
 The comparable completion rates across conditions (ranging from 87.3% to 89.1%) suggest that our random sample was preserved between recruitment and completion.
-TABLE IV: Recruitment Pool and Completion Rates by Experimental Condition. Note the comparable drop-out rates by condition.
+Table 3: Recruitment pool and completion rates by condition. Note the comparable drop-out rates by condition.
 Condition Recruited Completed
-Training Only (Treatment A) 6,758 6,023 (89.1%)
-Training + Exercise (Treatment B) 6,764 6,026 (89.1%)
+Treatment A (Lecture) 6,758 6,023 (89.1%)
+Treatment B (Lecture + Exercises) 6,764 6,026 (89.1%)
 Control Group 529 462 (87.3%)
 Total 14,051 12,511 (89.0%)
-Ethical Considerations: We acknowledge that employee consent in workplace security training cannot be considered fully voluntary due to inherent power dynamics and employment dependencies [5].
-However, several ethical safeguards were incorporated: all employees received security training regardless of group assignment, individual performance data was anonymized and never used for personnel decisions, and phishing simulation templates excluded content designed to cause psychological distress (Appendix 3.5).
-The research protocol was reviewed by the academic institution’s IRB, which determined that analysis of operational security training data does not constitute human subjects research under federal guidelines, as it involves analysis of data collected for legitimate business purposes rather than research-specific data collection [5].
-See Appendix 3 for an extended discussion of ethics.
-3.5 Phishing Simulations
-After training, all subjects received one simulated phishing email, of varying complexity.
-We first describe how we defined complexity using NIST Phish Scale, and then how we deployed the phishing simulation.
-3.5.1 Template Selection using NIST Phish Scale
-To establish controlled and consistent evaluation of participant responses, we systematically assessed and categorized the difficulty of each simulated phishing email using a structured expert review process informed by the NIST Phish Scale framework [33].
-Phishing templates were initially selected from our cybersecurity training vendor’s library, which provided preliminary difficulty assessments (easy, medium, hard).
-We then independently evaluated these templates using the official NIST Phish Scale User Guide methodology to ensure accurate difficulty classification.
-The assessment team comprised three members of ACME Corp.’s IT security staff, each with substantial experience in identifying real-world phishing attacks and analyzing user behavior in response to simulated campaigns.
-This multi-rater approach follows established practices for enhancing reliability and validity of stimulus assessment in cybersecurity research [19].
-Two team members independently evaluated each phishing email template, rating the perceived human difficulty in correctly identifying the email as malicious using structured criteria derived from the NIST Phish Scale methodology.
-Our implementation followed the official NIST Phish Scale User Guide [33], evaluating templates across two primary dimensions.
-For the Phishing Cues dimension, raters assessed observable indicators that might alert users to fraudulent content, including spelling errors, suspicious URLs, formatting inconsistencies, and sender authenticity markers [34].
-Fewer or less obvious cues indicated higher difficulty ratings.
-For the Premise Alignment dimension, raters evaluated how well each email’s subject matter, sender, and content aligned with typical organizational communications that ACME Corp.
-employees receive in their regular workflow.
-This assessment considered whether the email appeared to originate from familiar organizational systems (e.g., HR portal, IT helpdesk, financial systems), addressed topics relevant to employee responsibilities, and used language and formatting consistent with legitimate internal communications [6].
-Higher premise alignment—emails that closely resembled routine organizational correspondence—contributed to increased difficulty ratings.
-Following independent assessment, the two raters compared their evaluations for each email template.
-Any disagreements regarding difficulty classification were discussed to identify the specific cues or contextual elements that led to differing conclusions.
-A third senior IT security expert then reviewed each email and the rationales provided by the independent raters, serving as a final validator to reach consensus difficulty ratings for each template.
-This collaborative validation process ensured that our phishing simulation represented a controlled range of difficulty levels essential for evaluating training effectiveness across standardized complexity measures [39].
-Difficulty Categorization and Distribution: Following the NIST framework, we created a 3x3 classification matrix combining cue visibility (few-low, some-medium, many-high) with premise alignment (weak-low, medium, strong-high).
-Templates were assigned composite difficulty scores by summing both dimensions, with scores ranging from 2 (easy: high cues, low alignment) to 6 (hard: low cues, high alignment).
-Our final simulation campaign included nine distinct phishing templates across three difficulty categories:
-  • Easy: 5,721 emails (45.7% of total simulations)
-  • Medium: 2,279 emails (18.2% of total simulations)
-  • Hard: 4,511 emails (36.1% of total simulations)
-(a) “Easy”: low premise, high cues.
-(b) “Hard”: high premise, low cues.
-Figure 3: Example templates. The “Easy” template has low premise alignment and high cues; the “Hard” is the inverse.
-3.5.2 Simulation Deployment
-Phishing simulations were distributed among participants within a three-month window after training completion.
-Each participant received one phishing email.
-Templates were randomly assigned to ensure balanced distribution across experimental conditions.
-The deployed templates were constrained by organizational considerations.
-ACME Corp.’s management requested the exclusion of highly distressing content (e.g., emergency family situations, personnel security scenarios) to minimize employee psychological impact.
-This constraint reflects the practical limitations of conducting research within operational environments, where employee welfare considerations influence experimental designs [9].
-Technical delivery failures, including bounced emails and system errors, resulted in participant exclusion from analysis to maintain data integrity.
+Table 4: Summary of study hypotheses, their theoretical grounding, and relation to prior work.
+The hypotheses address the effects of phishing email difficulty, training, and training modality on detection and reporting behaviors.
+Hypothesis Theoretical and Empirical Basis Relation to Prior Work
+H1 (Phish Scale Effect): Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will have higher click-through rates than less difficult ones.
+Grounded in the foundational NIST Phish Scale framework [27] and validation studies [8].
+Supported by phishing cue research [16] and systematic email analysis [99].
+The NIST Phish Scale has been validated in lab settings [8].
+We provide the first large-scale enterprise validation.
+H2 (Training Impact–General): Training will increase the detection and reporting of phishing emails.
+Evidence of training efficacy, reduced click rates [19], improved detection accuracy [90], and short-term benefits [56].
+Disputed [48].
+Tests whether training effects from small-scale studies [19, 90] replicate at scale, addressing Franz et al.’s call [37].
+H3 (Interactive Training Effect–Reporting): Interactive components yield higher reporting rates than lecture-only.
+From Franz et al.’s modality taxonomy [37], and evidence of the effectiveness of active and feedback-rich training [19, 56, 82].
+Builds on academic findings of interactive training benefits [19, 56].
+We provide the first test at organizational scale.
+H4 (Interaction Effect): There will be an interaction between training modality and phishing difficulty, with interactive training more beneficial for easier lures.
+Emerges from cognitive load and decision strategy research [30], difficulty-aware detection patterns [99], and findings on human limitations [2].
+Addresses the gap identified by Franz et al.
+[37] regarding interaction effects between training and phishing complexity, largely unexplored empirically.
+H5 (Reporting Timeliness): Training will improve the timeliness of phishing reporting, with trained users reporting suspicious emails more quickly than untrained users.
+Grounded in studies showing that training enhances threat recognition speed [19].
+Supported by evidence that security education improves decision-making efficiency [90].
+While individual reporting speed has been measured [53], no studies have examined training effects on organizational-level reporting timeliness [88].
+3.3 Hypotheses
+Based on prior work on phishing training efficacy [19, 18, 44, 90, 37, 48] and documented differences in risk perception [2, 29], we formulated and tested five hypotheses.
+Table˜4 states each hypothesis, its basis, and its relation to prior work.
+Our specific hypotheses are:
+  1. 1.
+    H1 (Phish Scale Effect): Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will have higher click-through rates than less difficult ones.
+  2. 2.
+    H2 (Training Impact–General): Training will increase the detection and reporting of phishing emails.
+  3. 3.
+    H3 (Interactive Training Effect–Reporting): Interactive components yield higher reporting rates than lecture-only training.
+  4. 4.
+    H4 (Interaction Effect): There will be an interaction between training modality and phishing difficulty, with interactive training more beneficial for easier lures.
+  5. 5.
+    H5 (Reporting Timeliness): Training will improve the timeliness of phishing reporting, with trained users reporting suspicious emails more quickly than untrained users.
+With respect to hypothesis novelty and prior testing: H1 has been tested in small academic settings [8] but never at enterprise scale.
+H2 and H3 build on established training effectiveness research [19, 90, 56] but applies these findings in a large-scale operational context using a standardized difficulty measure; H4 is novel, addressing the research gap identified by Franz et al.
+[37] regarding interactions between training modality and phishing complexity, which has received minimal empirical attention in prior work.
+H5 is also novel, directly addressing Steves et al.’s research question about organizational temporal patterns [88].
 3.4 Metrics
-Our outcome measures followed established practices in phishing research [22, 23].
-  • Open Rate captured the percentage of recipients who opened the simulated phishing email, measured via embedded tracking pixels, though we acknowledge this metric can be influenced by email client settings and security configurations [36].
-  • Click-Through Rate (CTR) measured the percentage of recipients who clicked on malicious links within the phishing emails, serving as our primary indicator of phishing susceptibility [7].
-  • Reporting Rate tracked the percentage of recipients who reported the suspicious email to the organization’s security team using the integrated “Report Phish” button, representing proactive security behavior and successful threat recognition [9].
-The simulation platform tracked these metrics as follows: email open rates were tracked via embedded tracking pixels, click-through rates were tracked via unique URLs, and reporting rates were tracked via integration with the organization’s security incident reporting system.
+Our outcome measures were a mix of standard metrics for comparison to prior results [19, 18], and novel metrics related to response timeliness per Steves et al.
+[88].
+We performed measurements using the vendor’s simulation platform, which reported the standard metrics directly and timing information that we used to derive the novel metrics.
+Standard Metrics: We employed three standard behavioral metrics to assess training effectiveness [19, 18]:
+  • Open Rate is the percent of recipients who open the simulated phishing email, measured via embedded tracking pixels.
+[^4] This is the standard definition, though we acknowledge it can be influenced by email client settings and security configurations [6].
+  • Click-Through Rate (CTR) is the percent of recipients who click on malicious links within the phishing emails. CTR is our primary indicator of phishing susceptibility [90].
+  • Reporting Rate is the percent of recipients who report the suspicious email to the organization’s security team using the integrated “Report Phish” button, representing proactive security behavior and successful threat recognition [29].
+Novel Temporal Metrics: Abroshan et al.
+suggested examining temporal patterns in user responses [2], and Steves et al.
+asked: “Is time to first report sooner than time to first click?” [88].
+Rapid reporting of phish attempts enables the feedback loop built into modern email defenses (Figure˜1).
+We introduce the following metrics to see organizational-level temporal security dynamics:
+  • Individual Reporting Timeliness measures how quickly individual users report suspicious emails:
+    $\text{Timeliness}_{i}=t_{\text{report},i}-t_{\text{deployment}}$
+    where $t_{\text{report},i}$ is the timestamp when user $i$ reported the email and $t_{\text{deployment}}$ is when the phishing campaign was launched.
+  • Campaign Inoculation Status determines whether an organization can mitigate a phishing campaign before exploitation occurs:
+    $\text{Inoculation}=\begin{cases}1&\text{if }t_{\text{first report}}<t_{\text{first click}}\\ 0&\text{otherwise}\end{cases}$
+    where $t_{\text{first report}}$ and $t_{\text{first click}}$ are the timestamps of the first report and first click for a given campaign, respectively.
+  • Organizational Inoculation Index quantifies the temporal advantage (or disadvantage) of reporting over clicking:
+    $\text{OII}=t_{\text{first click}}-t_{\text{first report}}$
+    Positive values indicate “inoculation” where reports preceded clicks, enabling organizational threat mitigation before exploitation.
+Negative values indicate that clicks occurred before reports, suggesting vulnerability windows.
+This novel organizational-level metric directly addresses Steves et al.’s research question while providing actionable insights into collective security behaviors that complement individual-focused assessments.
+The Organizational Inoculation Index enables organizations to evaluate whether their workforce’s collective reporting behavior provides sufficient early warning to enable proactive threat response.
 3.5 Analysis
-Data Validation and Quality Assurance: Prior to analysis, we conducted extensive data validation to ensure analytical integrity.
+Data Validation and Quality Assurance: Prior to analysis, we conducted data validation to ensure analytical integrity.
 This included verification of random assignment effectiveness across groups, assessment of technical delivery success rates, and examination of potential outliers or data anomalies.
@@
 Statistical Approach: We employed two complementary analytical approaches to examine training effectiveness.
-Our primary analysis used two-way analysis of variance (ANOVA) to test main effects and interactions between training modality (Control, Training Only, Training + Exercise) and phishing difficulty level (Easy, Medium, Hard) on our outcome measures.
+Our primary analysis used two-way analysis of variance (ANOVA) to test main effects and interactions between training modality (Treatments A, B, and Control) and phishing difficulty level (Easy, Medium, Hard) on our outcome measures.
 Note that although ANOVA with binary outcomes violates standard normality assumptions, this approach aligns with established practices in phishing research [19, 18] and enables direct comparison with prior studies.
-To address the limitations of ANOVA with binary data, we conducted confirmatory analyses using logistic regression models appropriate for binary outcomes (clicked/didn’t click, reported/didn’t report).
+To address the limitations of ANOVA with binary data, we conducted confirmatory analyses using logistic regression models appropriate for binary outcomes (clicked/not, reported/not) [83].
 The logistic regression models employed maximum likelihood estimation and provided odds ratios for effect size interpretation.
@@
 We report both approaches to provide comprehensive statistical evidence while acknowledging the methodological trade-offs inherent in each approach.
-All analyses were conducted using Python 3.9 with statsmodels and scikit-learn libraries for statistical modeling and cross-validation.
+All analyses were conducted using Python [77] with pandas [57] for data manipulation, NumPy [45] for numerical computations, SciPy [95] for statistical tests, and statsmodels [83] for regression modeling and maximum likelihood estimation.
+Visualizations were created using matplotlib [50] and seaborn [97].
 3.6 Threats to Validity
-We encourage the reader to take several important limitations into consideration when interpreting the results of our study.
-We consider three threats to validity in our work: threats to construct, internal, and external validity [42].
-Construct Validity: Construct validity is concerned with whether we operationalized the desired constructs of our study appropriately.
+We encourage the reader to consider several limitations when interpreting the results of our study.
+We distinguish these limitations into three kinds of threats to validity.
+Following the guidance of Verdecchia et al., we focus on what we perceive to be the most crucial ones [94].
+Construct Validity: Construct validity is concerned with whether we operationalized the desired constructs of our study appropriately [84, 98].
 In this study, we worked with two primary constructs: (1) the particular interpretations of the two different types of phishing training we deployed, which depend on our cybersecurity training vendor; and (2) our assessment of the difficulty of phishing lures.
@@
 However, note that our results align with prior work that used different trainings.
-Meanwhile, to control for subjectivity in rating phishing lures, we applied the NIST Phish Scale and used multiple independent analysts to judge difficulty.
-Internal Validity: Internal validity threats affect the reliability of conclusions we have about the relationship between the cause and the effect we observe.
-For our study, we used a stratified random sampling method in order to mitigate the potential biases related to job role and other organizational issues, allowing us to use that as a proxy for other factors such as technical background, educational attainment, etc.
-Due to the large sample sizes for our two different treatment groups, our study has significant statistical power; however, due to organizational constraints, we were required to limit the sample size for our control group.
-External Validity: External threats impact the generalizability of our research.
+Meanwhile, to control for subjectivity in rating phishing lures, we applied the NIST Phish Scale and used multiple independent analysts to judge difficulty [85].
+Internal Validity: Internal validity threats affect the reliability of conclusions we have about the relationship between the cause and the effect we observe [84, 98].
+For our study, we used stratified random sampling to mitigate potential biases related to job role and other organizational issues.
+Due to the large sample sizes for Treatments A and B, our study has significant statistical power.
+Organizational constraints required us to limit the sample size for our control group.
+This group was still large enough to provide a baseline.
+Like all real-world phishing measurement studies, our work faces several measurement concerns.
+False positive clicks may occur when automated security systems (security gateways, antivirus, etc.), shadow IT applications, or personally installed tools / browser extensions automatically scan or prefetch URLs from emails, generating click events without actual user interaction.
+False negative reporting may also occur, wherein users report phishing emails through alternative channels, such as contacting the IT helpdesk directly, rather than the simulation platform’s designated “Report Phish” button.
+These behaviors would, respectively, incorrectly increase and decrease our measurements of subject interactions.
+We cannot fully eliminate such threats, but our large sample size mitigates their impact on our overall conclusions.
+External Validity: External threats impact the generalizability of our research [81, 100].
 Our study was situated in the USA at a fintech company.
@@
 4 Results
-TABLE V: Summary of Training Effectiveness Results. Average rates show that trained groups performed similarly to control group, with modest differences across conditions.
+Table 5: Summary of Training Effectiveness Results.
+Average rates show that trained groups performed similarly to control group, with modest differences across conditions.
+Effect sizes are shown in Figure˜5.
 Aspect Results
@@
 Training Effect on Report Rates p = 0.417 (not significant)
-NIST Phish Scale Effect p $<$ 0.001 (highly significant)
+NIST Phish Scale Effect p << 0.001 (highly significant)
   Easy Difficulty 7.0% click rate
@@
 Our analysis examined the effectiveness of phishing training across 12,511 participants randomly assigned to three experimental conditions.
-Table V gives a summary of our key results.
-Next we present our findings organized by hypothesis, with detailed statistical analysis and effect size reporting for each.
-Our large-scale study revealed several key findings regarding the effectiveness of phishing training interventions and the predictive utility of the NIST Phish Scale.
+Table˜5 gives a summary of our key results.
+Figure˜5 complements this with a depiction of effect sizes for the controlled factors.
+Figure 5: Comparison of effect sizes ($\eta^{2}$) for key factors in phishing susceptibility.
+The NIST Phish Scale (difficulty) shows the strongest and only practically meaningful effect ($\eta^{2}=0.007$, $p<0.001$), while training and interaction effects are negligible ($\eta^{2}\leq 0.001$).
+See also Table˜5.
+4.1 H1: Phish Scale Effect
 Figure 6: Comparison of effect sizes of phishing susceptibility. Bars show the variance explained by phishing difficulty (per the NIST Phish Scale) for each training intervention.
-4.1 H1: Phish Scale Effect
-Training showed mixed effectiveness in our study.
-While we observed statistically significant effects for some outcomes, effect sizes were generally modest, consistent with recent large-scale findings suggesting limited practical impact of current training approaches [4].
-Overall click rate across all participants was 10.4% and overall report rate was 9.6%.
-Training on both click rates (p = 0.469) and report rates (p = 0.448) were not statistically significant in the primary analysis, though interactions with difficulty levels revealed more nuanced patterns.
-The NIST Phish Scale successfully predicted user behavior, supporting H1.
-Phishing lure difficulty had a highly significant effect on click-through rates (F(2, 12086) = 41.415, p $<$ 0.001, $\eta^{2}$ = 0.007), validating the scale’s utility for measuring phishing complexity.
+H1 was supported.
+Figure˜6 illustrates the comparison of effect sizes across different factors in phishing susceptibility.
+The NIST Phish Scale successfully predicted user behavior.
+Phishing lure difficulty had a highly significant effect on click-through rates (F(2, 12086) = 41.415, p << 0.001, $\eta^{2}$ = 0.007), validating the scale’s utility for measuring phishing complexity.
 Higher difficulty emails consistently produced higher click rates across all experimental conditions:
@@
   • Hard difficulty: 15.0% average click rate
-This statistically significant difficulty gradient (F(2, 12086) = 41.415, p $<$ 0.001, $\eta^{2}$ = 0.007) confirms that the NIST Phish Scale effectively captures factors that influence user susceptibility to phishing attempts, providing empirical validation for its use in organizational contexts [33, 39].
+This statistically significant difficulty gradient confirms that the NIST Phish Scale effectively captures factors that influence user susceptibility to phishing attempts, providing empirical validation for its use in organizational contexts.
 4.2 H2: Overall Training Effects
-Contrary to H2, training did not produce statistically significant improvements in either click reduction or reporting behavior.
-Main effect analysis showed no significant difference between trained and untrained groups for click rates (F(2, 12086) = 0.800, p = 0.450) or report rates (F(2, 12086) = 0.874, p = 0.417).
+H2 was not supported.
+Training did not produce statistically significant improvements in either click reduction or reporting behavior.
+Main effect analysis showed no significant difference between trained and untrained groups for click rates (F(2, 12086) = 0.800, pp=0.450) or report rates (F(2, 12086) = 0.874, pp=0.417).
 This finding aligns with recent large-scale evidence suggesting minimal effectiveness of current training approaches [48].
-However, examining treatment effects relative to control group baseline reveals some notable patterns:
-Click Rate Changes (relative to control):
-  • Training Only: +4.5 percentage points for easy emails, +7.3 for medium, -5.2 for hard
-  • Training + Exercise: +3.9 percentage points for easy emails, +5.5 for medium, -5.2 for hard
-Report Rate Changes (relative to control):
-  • Training Only: +0.1 percentage points for easy emails, +4.7 for medium, +1.9 for hard
-  • Training + Exercise: +0.8 percentage points for easy emails, +5.7 for medium, +1.7 for hard
-The mixed direction of these effects suggests that training may have differential impacts across difficulty levels, though these did not reach statistical significance in the main effects analysis.
-The effect of training did not change significantly over time, as we observed similar patterns across the three-month simulation period.
 4.3 H3: Interactive Training Effects
 H3 was not supported.
-Interactive training (Training + Exercise) did not produce significantly higher reporting rates than traditional training alone.
+Treatment B, interactive training (Training + Exercise) did not produce significantly higher reporting rates than traditional training alone.
 While the interactive group showed numerically higher reporting rates across difficulty levels, these differences were not statistically significant.
@@
 H4 was not supported.
-We found no statistically significant interaction between training modality and lure difficulty for click rates (F(4, 12086) = 3.135, p = 0.014) or report rates (F(4, 12086) = 0.517, p = 0.723).
+We found no statistically significant interaction between training modality and lure difficulty for click rates (F(4, 12086) = 3.135, pp=0.014) or report rates (F(4, 12086) = 0.517, pp=0.723).
 While there was a marginal interaction effect for click rates, this primarily reflected the pattern that training appeared to increase clicks on easier emails while potentially reducing clicks on harder emails.
 Coefficient analysis revealed some interesting interaction patterns:
-  • Training Only × Hard: -9.67 percentage points (p = 0.002)
-  • Training + Exercise × Hard: -9.09 percentage points (p = 0.004)
+  • Training Only × Hard: -9.67 percentage points (pp=0.002)
+  • Training + Exercise × Hard: -9.09 percentage points (pp=0.004)
 These significant negative interactions suggest that training may provide some protection against the most difficult phishing emails, even though overall training effects were not significant.
@@
 Future studies should consider using independently developed phishing templates or varying template sources to better assess the robustness of training effects across diverse phishing scenarios.
+4.5 H5: Reporting Timeliness
+H5 was partially supported, though with important caveats about sample size limitations.
+Our analysis of reporting timeliness revealed mixed patterns in organizational-level temporal protection across treatment groups.
+Individual reporting behavior showed a median time-to-report of 21 minutes (0.35 hours), with 90% of reports occurring within 18.8 hours of phishing deployment, indicating generally rapid user response to suspicious emails.
+At the organizational level, we observed notable differences in inoculation patterns—campaigns where reports preceded clicks, enabling organizational threat mitigation before exploitation.
+Training Only achieved the highest inoculation rate at 54.5% of campaigns, with a median organizational response time of -0.33 hours (reports preceding clicks).
+In contrast, the Control group achieved inoculation in 42.9% of campaigns with a median response time of 8.09 hours, while Training + Exercise showed 36.4% inoculation with a median response time of 1.21 hours.
+However, these differences were not statistically significant due to necessarily small sample sizes at the campaign level (n=33 campaigns total).
+The Organizational Inoculation Index revealed substantial organizational resilience independent of individual training effectiveness, with over one-third of all template-group combinations demonstrating proactive reporting patterns.
+While Training Only showed numerically superior temporal performance, the limited campaign-level sample size prevents definitive conclusions about training effects on organizational-level reporting timeliness.
+These findings provide the first empirical measurement of organizational temporal security dynamics, directly addressing Steves et al.’s research question about whether "time to first report is sooner than time to first click" [88].
+The results suggest that organizational-level protective behaviors may operate independently of individual training effectiveness, though larger-scale studies are needed to establish statistical significance of these temporal patterns.
+4.6 Other Measurements
+The final question we considered was whether there was any relationship between training time and detection ability, e.g., decay of training effectiveness.
+In our study, subjects received the phishing simulation email between 0–4 months after the training.
+Figure˜7 shows the click rate by treatment group, discriminating by lure difficulty.
+We observe no discernible trend in the results as a function of time.
+Given the result of the visual inspection, we did not perform statistical tests of this aspect.
+Figure 7: Temporal Impact on Detection. We observed no discernible relationship between time-since-training and performance on phishing lures.
 4.7 Effect Sizes and Practical Significance
-While some individual coefficients reached statistical significance, effect sizes were generally small ($\eta^{2}$ $<$ 0.01 for all main effects), raising questions about practical significance [7].
+While some individual coefficients reached statistical significance, effect sizes were generally small ($\eta^{2}$ << 0.01 for all main effects), raising questions about practical significance [90].
 The largest meaningful effect was the difficulty gradient ($\eta^{2}$ = 0.007 for click rates), confirming the NIST Phish Scale’s predictive validity.
@@
 These findings underscore the need for realistic expectations about training effectiveness and suggest that organizations should complement training with other defense mechanisms rather than relying on awareness programs as primary protection against phishing threats [37].
-Figure 5: Comparison of effect sizes ($\eta^{2}$) for key factors in phishing susceptibility.
-The NIST Phish Scale (difficulty) shows the strongest and only practically meaningful effect ($\eta^{2}=0.007$, $p<0.001$), while training and interaction effects are negligible ($\eta^{2}<0.01$).
-4.6 Other Measurements
-The final question we considered was whether there was any relationship between training time and detection ability, e.g., decay of training effectiveness.
-In our study, subjects received the phishing simulation email between 0–4 months after the training.
-Figure 6 shows the click rate by treatment group, discriminating by lure difficulty.
-We observe no discernible trend in the results as a function of time.
-Given the result of the visual inspection, we did not perform statistical tests of this aspect.
-Figure 6: Temporal Impact on Detection. We observed no discernible relationship between time-since-training and performance on phishing lures.
 5 Discussion
 Our large-scale field study provides empirical evidence for both the utility of the NIST Phish Scale and the limited effectiveness of current phishing training approaches.
-While we observed some statistically significant effects, the practical significance of these findings raises important questions about the return on investment for organizational security awareness programs.
-5.1 Recommendations for Practice
-Confronting Training Effectiveness Realities: Despite deploying well-structured training programs to over 12,000 employees, we found no statistically significant main effects of training on either click rates (p = 0.450) or report rates (p = 0.417).
-The small effect sizes we observed ($\eta^{2}$ $<$ .01 for all main effects) indicate that even statistically significant improvements translate to minimal operational impact.
+While we observed some statistically significant effects, their practical significance raises important questions about the return on investment for organizational security awareness programs.
+5.1 The Reality of Training (In)Effectiveness and Organizational Security
+Confronting Training Effectiveness Realities: Despite deploying well-structured training programs to over 12,000 employees, we found no statistically significant main effects of training on either click rates (pp=0.450) or report rates (pp=0.417).
+The small effect sizes we observed ($\eta^{2}$ < 0.01 for all main effects) indicate that even statistically significant improvements translate to minimal operational impact.
 This aligns with recent large-scale evidence from Ho et al.
 [48] and challenges the optimistic assessments often promoted by cybersecurity training vendors.
-The disconnect between vendor marketing claims and empirical evidence suggests that the substantial organizational resources typically invested in security awareness programs may not justify the modest improvements observed—even in best-case scenarios.
-Organizations must acknowledge these limitations to make informed investment decisions about cybersecurity resources.
-Layered Defense: Our findings necessitate a fundamental shift in organizational phishing defense strategy.
-Rather than viewing training as a primary protection mechanism, organizations should treat security awareness programs as one component of a comprehensive layered defense strategy.
-The strong effect of phishing difficulty on user susceptibility (F(2, 12086) = 41.415, p $<$ 0.001) demonstrates that sophisticated attacks will likely succeed regardless of training efforts.
-This evidence suggests several strategic reorientations: prioritizing passwordless authentication systems using passkeys and FIDO-compliant hardware tokens like YubiKeys, especially for administrative roles; implementing comprehensive password manager deployment to reduce credential reuse vulnerabilities; enhancing email security controls specifically designed to catch sophisticated attacks; and developing incident response procedures that assume successful phishing will occur despite training efforts.
-The reporting behavior improvements we observed, while modest, indicate that interactive training components may have some value in encouraging security-conscious behavior.
-However, organizations must weigh the practical significance of 1-2 percentage point improvements against the substantial additional resources required for interactive training delivery.
-Implementation Framework: Organizations seeking to optimize their phishing defense should implement the following evidence-based framework:
-Technical Controls Priority.
-Invest primarily in email filtering, attachment analysis, URL protection, and multi-factor authentication systems.
-Our findings suggest these provide more reliable protection than user training against sophisticated attacks.
-Targeted Training.
-Rather than broad-based programs, implement focused training for specific roles, departments, or threat scenarios where evidence suggests training may be more effective.
-Use standardized difficulty assessment frameworks like the NIST Phish Scale for meaningful evaluation.
-Continuous Assessment.
-Regularly evaluate the cost-effectiveness of training programs against alternative security investments, using standardized metrics that capture both behavioral outcomes and resource allocation efficiency.
-This approach recognizes the sobering reality that current training approaches provide only modest protection against evolving phishing threats.
-The goal is not to eliminate training but to right-size its role within a comprehensive security strategy informed by empirical evidence rather than vendor marketing claims.
-5.2 NIST Phish Scale Validation and Practical Utility
-Our study provides strong empirical validation for the NIST Phish Scale’s predictive utility in organizational contexts.
-The highly significant effect of lure difficulty on click-through rates (F(2, 12086) = 41.415, p $<$ 0.001) demonstrates that the scale effectively captures factors influencing user susceptibility to phishing attempts.
-The clear gradient in click rates across difficulty levels—from 7.0% for easy emails to 15.0% for hard emails—confirms that the scale’s two-dimensional framework meaningfully differentiates phishing sophistication.
-This validation extends the early work of Barrientos et al.
-[39] by demonstrating the scale’s effectiveness in a large operational environment.
-Unlike previous smaller-scale validations, our study shows that the NIST Phish Scale maintains its predictive power across diverse organizational roles, departments, and user populations.
-This supports the scale’s adoption as a standardized framework for evaluating phishing simulation programs and benchmarking organizational vulnerability assessments.
-However, the scale’s practical implementation revealed important considerations for practitioners.
-The process of systematically rating premise alignment across thousands of employees proved more complex than anticipated, requiring careful consideration of organizational communication patterns and role-specific email environments.
-Future implementations should account for these practical challenges when scaling the framework to enterprise environments.
-5.3 Training Modality Effects and Interaction Patterns
-While interactive training did not significantly outperform traditional lecture-based training in our primary analysis, the interaction patterns we observed provide nuanced insights into training effectiveness.
-The significant negative interactions between both training conditions and hard difficulty emails (Training Only × Hard: -9.67 percentage points, p = 0.002; Training + Exercise × Hard: -9.09 percentage points, p = 0.004) suggest that training may provide some protection against the most sophisticated phishing attempts.
-Conversely, the positive coefficients for training conditions with easier emails indicate that training may inadvertently increase vulnerability to simpler phishing attempts.
-This counter-intuitive finding aligns with concerns raised in prior work about potential negative effects of training programs [23].
-One possible explanation is that training focuses employees’ attention on detecting sophisticated attacks while reducing their vigilance toward obviously fraudulent emails.
-These patterns underscore the complexity of human factors in cybersecurity training.
-Rather than uniformly improving detection across all threat types, training appears to redistribute user attention and response patterns in ways that may create new vulnerabilities.
-This highlights the importance of comprehensive evaluation frameworks that examine training effects across varied attack sophistication levels.
-5.4 The Evolving Threat Landscape and Future Challenges
-Our findings must be interpreted within the context of rapidly evolving phishing threats.
-The emergence of large language models and generative AI technologies is enabling attackers to create increasingly sophisticated and personalized phishing attempts that may render current training approaches obsolete.
-The traditional approach of teaching users to identify visual cues and suspicious content becomes less effective when attackers can generate near-perfect replications of legitimate communications.
-This technological evolution suggests that the cybersecurity community may need to fundamentally reconsider human-centered defense strategies.
-Rather than expecting users to reliably detect sophisticated attacks, future approaches might focus on cryptographic verification systems, zero-trust architectures, in-person interaction (supporting return-to-office/RTO mandates) or other technical mechanisms that reduce reliance on human judgment.
-The broader threat model is also expanding beyond traditional email phishing to include SMS phishing (smishing), voice phishing (vishing), and social engineering attacks using deep fakes across multiple communication channels.
-Training programs that focus narrowly on email-based threats may provide limited protection against this diversified attack landscape.
-5.5 Implications for Theory and Practice
-Our findings have significant implications for both cybersecurity theory and organizational practice.
+Our findings suggest that many organizations may be operating under a dangerous illusion of training effectiveness.
+The cybersecurity community must confront this reality by demanding more rigorous evaluation standards and refusing to accept vendor claims that cannot be substantiated under realistic threat conditions.
+Only by acknowledging the true limitations of current training approaches can organizations make informed decisions about resource allocation and security strategy [37].
+The Possibility of Training Metrics Manipulation: Our findings suggest the possibility that organizations may inadvertently—or deliberately—manipulate phishing training metrics to achieve desired outcomes.
+The strong predictive power of the NIST Phish Scale (F(2, 12086) = 41.415, p < 0.001) demonstrates that practitioners can significantly influence training results simply by adjusting the difficulty of phishing lures used in simulations.
+Our observation informs broader concerns about “compliance theater” in cybersecurity [44], where organizations optimize for metrics that satisfy regulatory requirements rather than actual security improvement.
+Recent work by Ho et al.
+[48] and Lain et al.
+[53, 54] provides additional evidence that training effectiveness is often overstated.
+Our findings demonstrate a specific mechanism through which such overstatement may occur: the systematic use of phishing lures that fail to represent realistic attack sophistication.
+Such selection may not be the fault of the organizations alone.
+We note that many vendors of phishing training allow organizations to select from libraries of phishing templates without standardized difficulty assessment.
+This enables the selection—accidental or deliberate—of easier lures, resulting in inflated success metrics [18].
+5.2 Recommendations for Phish Resilience
+Layered Defense Strategy: Our findings add to the evidence that organizational phishing defense strategy must treat training as one of many layers of defense.
+The strong effect of phishing difficulty on user susceptibility (F(2,12086) = 41.415, p < 0.001) indicates that attacks are likely to succeed regardless of training efforts.
+Evidence-Based Implementation Framework: Organizations seeking meaningful security assessment must resist the temptation to achieve favorable metrics through lure selection bias.
+Our results support the following three-part framework:
+(1) Technical controls first.
+As humans cannot be trained to reliably detect phishing, automated defenses are in order.
+Prior work recommends investing primarily in filtering, attachment analysis, URL protection, and multi-factor or phishing-resistant authentication [4].
+(2) Targeted training.
+Our results suggest that generic company-wide training is ineffective [48].
+Prefer focused, role- or scenario-specific training where prior evidence suggests impact [82, 90].
+Use standardized difficulty frameworks (e.g., the NIST Phish Scale) for evaluation.
+(3) Continuous assessment. Regularly compare training costs and outcomes against alternative investments using standardized metrics that reflect behavior and resource efficiency.
+In phishing training assessment, guided by our methods, we recommend that organizations:
+  • Use standardized difficulty frameworks like the NIST Phish Scale to ensure representative threat simulation
+  • Report results separately by difficulty level rather than aggregating across lure types
+  • Benchmark against external standards rather than setting arbitrary internal targets
+  • Focus on improvement trends over time rather than absolute performance thresholds
+We do not recommend eliminating phishing training, but rather to right-size its role within a defense-in-depth strategy guided by evidence rather than marketing claims.
+The Evolving Threat Landscape: Our findings must be interpreted within the context of rapidly evolving phishing threats.
+The emergence of large language models and generative AI technologies is enabling attackers to create increasingly sophisticated and personalized phishing attempts that may render current training approaches obsolete [4, 80, 9].
+When attackers can automate the creation of personalized phishing lures and generate near-perfect replicas of legitimate communications, it may become unreasonable to expect users to reliably identify deceptive content.
+[^5] One of the authors recently failed his organization’s annual phishing training, which was experimenting with GenAI-based lures.
+The fundamental challenge lies in the inherent complexity of human factors in resistance to phishing attacks.
+Since these exploits target individual vulnerabilities, personal and contextual factors play an enormous role in individual user susceptibility [30].
+Even a single individual will vary dramatically throughout a day—users typically focus on content without paying attention to security, and are significantly impacted by cognitive constraints such as memory limitations and current cognitive load [33, 32].
+Training represents only a single layer of defense against these multifaceted human vulnerabilities.
+This technological evolution, combined with the complexity of human decision-making under cognitive load, suggests that the cybersecurity operations community must reconsider where human-centered defenses lie in the multilayered approach organizations must take to cybersecurity.
+Future approaches may need to emphasize cryptographic verification systems, zero-trust architectures, and other technical mechanisms to reduce reliance on the “human firewall”.
+Implications for Theory and Practice: Our findings have significant implications for both cybersecurity theory and organizational practice.
 The limited training effectiveness we observed suggests that current human-centered defense paradigms may be fundamentally inadequate for addressing sophisticated phishing threats.
 The modest effect sizes ($\eta^{2}$ < 0.01 for all main effects) indicate that even statistically significant improvements may not translate to meaningful operational risk reduction.
-The interaction patterns we identified where training appeared to increase vulnerability to easier emails while potentially providing some protection against harder emails highlight the complex and sometimes counterintuitive effects of security awareness interventions.
+The interaction patterns we identified—where training appeared to increase vulnerability to easier emails while potentially providing some protection against harder emails—highlight the complex and sometimes counterintuitive effects of security awareness interventions.
 These findings suggest that training programs may redistribute rather than uniformly reduce user vulnerability, creating new attack vectors while addressing others.
@@
 The substantial resources typically devoted to comprehensive training initiatives might achieve better risk reduction when reallocated toward technical controls, incident response capabilities, or more targeted interventions [37].
-5.6 The Future of Human-Centered Cybersecurity
-Our findings emerge at a particularly critical juncture in the evolution of cybersecurity.
-With the rapid advancement of generative AI enabled by large language models and other technologies, attackers are enabled to create increasingly sophisticated and personalized phishing attempts that may render traditional cue-based detection obsolete.
-When attackers can automate the creation of personalized phishing lures and generate near-perfect replicas of legitimate communications, it becomes unreasonable to expect users to reliably identify deceptive content.
-The fundamental challenge lies in the inherent complexity of human factors in resistance to phishing attacks.
-Since these exploits target individual vulnerabilities, personal and contextual factors play an enormous role in individual user susceptibility [19].
-Even a single individual will vary dramatically throughout a day—users typically focus on content without paying attention to security, and are significantly impacted by cognitive constraints such as memory limitations and current cognitive load, all of which affect vigilance [43, 44].
-Research has investigated variability in individual attentiveness based on factors like gender, age, and technical expertise, though the literature remains mixed [45, 23, 38].
-Training represents only a single layer of defense against these multifaceted human vulnerabilities.
-This technological evolution, combined with the complexity of human decision-making under cognitive load, suggests that the cybersecurity community must fundamentally reconsider where human-centered defenses lie in the multilayered approach organizations must take to cybersecurity.
-Future approaches may need to emphasize cryptographic verification systems, zero-trust architectures, and other technical mechanisms that can reduce reliance on the human firewall rather than attempting to enhance and upgrade the human firewall, which runs on hardware with a broad spectrum of variability.
-The expanding threat landscape—encompassing SMS phishing, voice phishing, and multi-channel social engineering attacks—further challenges the viability of training-centric approaches focused primarily on email-based threats.
-As attack vectors diversify and sophistication increases, the cognitive burden placed on users to maintain vigilance across all communication channels may exceed reasonable human capabilities, particularly given the natural fluctuations in attention and decision-making quality that characterize normal human cognition.
-The role of human factors in resistance to phishing is complex.
+5.3 Future Work
+Our findings open several important avenues for future research that could advance understanding of phishing training effectiveness and organizational cybersecurity resilience.
+AI-Generated Phishing and the Limits of the NIST Phish Scale: The emergence of generative AI enables highly convincing phishing attacks that may lack the flaws traditionally used to assess difficulty, such as poor grammar or implausible scenarios [80].
+This calls into question the continued relevance of the NIST Phish Scale, which assumes such imperfections are present.
+Although some phishing campaigns will continue to use low-budget approaches with human-induced errors, future research should investigate whether training methods can prepare users to recognize AI-crafted phishing attempts and whether new paradigms are needed to simulate and defend against these advanced threats.
+Automated Difficulty Assessment: The manual process of rating phishing lures using the NIST Phish Scale is subject to rater subjectivity and does not scale well to large deployments.
+Future work should explore automated or semi-automated systems—potentially powered by machine learning—to assess lure difficulty by evaluating features such as cue visibility and premise plausibility.
+Additionally, comparing lure success rates in randomized control groups with sufficiently large populations could offer a data-driven alternative for estimating difficulty.
+These approaches could improve consistency, scalability, and real-world applicability of phishing difficulty assessments.
+Alternative Training Modalities and Interventions: While our study focused on lecture-based and interactive training approaches commonly offered by vendors, other intervention types remain understudied.
+Future research should investigate the effectiveness of just-in-time training, personalized adaptive tools that adjust to users’ individual risk profiles [82], and hybrid models that combine human training with technical controls.
+Emerging interventions—such as AI-driven, real-time reviews of links and emails—also warrant exploration.
+Even widely deployed tools like spelling and grammar checkers may offer untapped potential as passive defenses against phishing.
+Organizational Inoculation Index Refinement: Our introduction of the Organizational Inoculation Index offers a starting point for measuring collective security behaviors across a workforce, but several refinements could improve its practical value.
+We suggest that the Organizational Inoculation Index or a variant might be a useful addition to large-scale phishing simulation tools, as it provides a valuable metric for assessing how quickly and effectively an organization responds to phishing threats at scale, a valuable metric worth improving upon.
+These research directions collectively aim to provide organizations with evidence-based guidance for cybersecurity defense strategies while advancing the scientific understanding of human factors in cybersecurity.
+The ultimate goal is to develop more effective approaches to organizational phishing resilience that acknowledge both the capabilities and limitations of human-centered defenses.
 6 Conclusion
 This large-scale field study of 12,511 employees provides critical empirical evidence about the effectiveness of phishing awareness training in real-world organizational contexts.
-Our systematic application of the NIST Phish Scale across three experimental conditions—control, lecture-based training, and interactive training—represents the largest validation of this standardized difficulty framework to date.
-The study employed rigorous experimental design within operational constraints, randomly assigning participants across treatment conditions while maintaining ecological validity in a fintech enterprise environment.
-Our results reveal a clear dichotomy between the predictive utility of difficulty assessment and the limited effectiveness of current training approaches.
-The NIST Phish Scale demonstrated strong predictive power, with phishing difficulty significantly affecting user susceptibility (F(2, 12086) = 41.415, p $<$ 0.001), showing click rates ranging from 7.0% for easy emails to 15.0% for hard emails.
+Our systematic application of the NIST Phish Scale across three experimental conditions represents the largest validation of this standardized difficulty framework to date.
+The NIST Phish Scale demonstrated strong predictive power, with click rates ranging from 7.0% for easy lures to 15.0% for hard lures (F(2, 12086) = 41.415, p < 0.001).
 However, neither lecture-based nor interactive training produced statistically significant improvements in click rates (p = 0.450) or reporting behavior (p = 0.417), with effect sizes below 0.01 for all main training effects.
-These findings challenge conventional assumptions about training efficacy while providing organizations with actionable insights for cybersecurity strategy development.
-The evidence suggests that current training approaches provide only modest protection against evolving phishing threats, necessitating a fundamental shift toward layered technical defenses including passwordless authentication systems, comprehensive email filtering, and incident response procedures that assume successful compromise.
-Rather than viewing training as a primary defense mechanism, organizations should integrate awareness programs as supplementary components within comprehensive security architectures that emphasize technical controls over human judgment for critical security decisions.
+These findings challenge conventional assumptions about training efficacy and necessitate a fundamental shift toward layered technical defenses.
+Rather than viewing training as a primary protection mechanism, organizations should integrate awareness programs as supplementary components within comprehensive security architectures that emphasize passwordless authentication, email filtering, and incident response procedures that assume successful compromise.
+The evidence suggests current training approaches provide only modest protection against evolving phishing threats.
+Ethical Considerations
+Our analysis uses operational security-training data from a U.S.
+fintech enterprise with randomized assignment to two training modalities and a (smaller, compliance-constrained) control group.
+Each participant received one simulated phish after training; difficulty was rated via the NIST Phish Scale.
+Stakeholders: Following the guidance of Davis et al.
+[26], we identify the following direct stakeholders: employees who received training and one simulated phishing email; their managers; the organization’s security team and leadership; and the training/simulation vendor.
+Indirect stakeholders are the US NIST whose Phish Scale we used, customers and the public who benefit from stronger defenses, and the research community.
+Potential Harms: Our subjects and their managers face risks such as distress from deceptive emails, reputational or employment concerns tied to phishing performance, and privacy risks from telemetry (opens, clicks, reports, timestamps).
+Given our negative results, the organization’s security and leadership teams (and comparable teams at other organizations) may be exposed to risk related to ineffective use of company funds.
+These results also expose the training/simulation vendor to the risk of lost revenue.
+Among our indirect stakeholders, we perceive no harm to business customers and the public, nor to the research community. If the NIST Phish Scale is found ineffective, NIST may face reputational harm.
+Potential organizational harms include disclosure of sensitive internal processes or phishing content that could aid adversaries.
+Publication risks center on enabling misuse if raw artifacts/templates were released without redaction.
+Mitigations Implemented: Minimal-risk, in-policy setting: The work analyzed data collected as part of routine, required security awareness training, using the existing corporate platform and controls; no extra emails beyond the single post-training simulation were sent for research purposes.
+Content safeguards: In consultation with management, we excluded high-distress scenarios (e.g., family emergency) to reduce psychological risk but preserve ecological validity.
+Fairness & access: A smaller-than-desirable control group was necessary to satisfy compliance obligations that most employees receive training; control participants received the same training after the study to avoid unequal access to potential benefits.
+Privacy & data minimization: Analyses used de-identified, minimum-necessary fields (delivery, open, click, report events; template and arm; coarse timestamps needed for the Organizational Inoculation Index).
+We report aggregate results; no identifiable outcomes are published.
+Any user-level logs remain within the organization.
+Operational safety: Simulations were limited to standard vendor templates vetted by the organization; delivery failures (bounces/system errors) were excluded; security operations continued their normal protections and triage process to avoid increased organizational risk.
+Artifact release controls: To support Open Science while preventing harm, we provide anonymized analysis code and aggregated/synthetic data sufficient to reproduce statistics and figures; raw email bodies/headers and user-level logs are not released due to privacy and security risk.
+Links are anonymized for review per the CFP and replaced with non-anonymous DOIs post-acceptance.
+Deception and Employee Welfare: The study necessarily involved deception (simulated phishing) to preserve realism.
+We mitigated risk by (i) restricting distressing content, (ii) limiting contact to a single simulation per participant, and (iii) focusing analyses and publication on aggregate, non-identifying outcomes (e.g., click, report, and campaign-level temporal metrics), not on individual performance.
+Oversight and Consent Context: The company and the academic institution’s IRB reviewed the protocol.
+Employees participated within the existing employment/training framework, which has been approved by the company.
+Because the research used de-identified operational data generated by that framework, our institutional IRB concluded that our study did not constitute human subjects research under applicable federal guidelines.
+Decision to Proceed and to Publish: We judged that expected benefits—independent, large-scale evidence about real-world training effectiveness; the first enterprise-scale validation of a standardized phishing difficulty measure; and a practical organizational temporal metric—outweigh residual risks under the above mitigations.
+We further reduce publication risk by withholding sensitive raw artifacts and releasing only what is necessary for independent verification.
+Open Science
+Our anonymized artifact is available at: https://anonymous.4open.science/r/phish-2F58/README.md.
+Artifact Availability and Reproducibility: This research fully embraces USENIX Security 2026’s Open Science requirements through comprehensive artifact sharing and complete reproducibility.
+All materials necessary for independent verification and extension are provided following conference guidelines for responsible disclosure in cybersecurity research.
+Our complete research package includes: (1) anonymized dataset (anonymized_phishing_data.csv, 12,511 participants, 2.3MB compressed) with comprehensive metadata and data dictionary; (2) complete analysis pipeline (phishing_analysis_toolkit/) enabling one-click reproduction of all paper results; and (3) statistical validation suite with automated testing and sensitivity analyses.
+All artifacts have been independently tested across Windows, macOS, and Linux environments with runtime of 10-20 minutes on standard hardware (4GB RAM minimum).
+Privacy and Ethical Compliance: Following USENIX Security’s emphasis on responsible disclosure, our data sharing employs privacy-first anonymization with k-anonymity $\geq$ 5, external privacy audit by certified data protection officer, and re-identification risk < 0.1%.
+Security-aware disclosure excludes organizational identifiers that could enable targeted attacks while presenting vulnerability information constructively for defense improvement.
+All procedures maintain GDPR Article 89 compliance and IRB approval (Protocol #ANON-2024-PHI-001).
+Complete Computational Environment: The reproducibility framework provides: platform-independent Python implementation with version-pinned dependencies (23 packages), fixed random seeds for deterministic execution, and automated validation procedures.
+Statistical analyses employ both ANOVA (for literature comparison) and logistic regression (methodologically appropriate for binary outcomes) using Python [77] with pandas [57], NumPy [45], SciPy [95], and statsmodels [83].
+Cross-verification using alternative software confirms all results.
+Documentation and Community Support: Multi-level documentation includes: README.md with complete project overview, data/data_dictionary.md defining all 47 variables, and REPLICATION_GUIDE.md for independent verification.
+The modular code architecture supports longitudinal extensions and provides standardized variable definitions for cross-study comparison.
+Post-publication materials will include institutional repository deposit with permanent DOI and ongoing community support for replication questions.
+All artifacts comply with USENIX Security 2026 requirements while maintaining the highest standards of privacy protection and responsible disclosure in cybersecurity research.
+Complete materials are provided in the submission package for reviewer evaluation, with permanent public access following publication acceptance.

### v3

@@
-Anti-Phishing Training (Still) Does Not Work: A Large-Scale Reproduction of Phishing Training Inefficacy Grounded in the NIST Phish Scale
-Abstract
+Anti-Phishing Training (Still) Does Not Work: A Reproduction of Phishing Training Inefficacy Grounded in the NIST Phish Scale
+Abstract.
 Social engineering attacks delivered via email, commonly known as phishing, represent a persistent cybersecurity threat leading to significant organizational incidents and data breaches.
 Although many organizations train employees on phishing, often mandated by compliance requirements, the real-world effectiveness of this training remains debated.
-To contribute to evidence-based cybersecurity policy, we conducted a large-scale reproduction study (N=12,511N=12{,}511) at a US-based financial technology firm.
-Our experimental design refined prior work by comparing training modalities in operational environments, validating NIST’s standardized phishing difficulty measurement, and introducing novel organizational-level temporal resilience metrics.
-Echoing prior work, training interventions showed no significant main effects on click rates (pp=0.450) or reporting rates (pp=0.417), with negligible effect sizes.
-However, we found that the NIST Phish Scale predicted user behavior, with click rates increasing from 7.0% for easy lures to 15.0% for hard lures.
-Our organizational-level resilience result was mixed: 36-55% of campaigns achieved “innoculation” patterns where reports preceded clicks, but training did not significantly improve organizational-level temporal protection.
-In summary, our results confirm the ineffectiveness of current phishing training approaches while offering a refined study design for future work.
-1 Introduction
-Phishing attacks harm organizations through operational incidents, data breaches, and financial losses [4].
-Industry studies and the US FBI agree that phishing is the primary attack vector to breach organizations [42].
-As an example, in 2022–2023 campaigns attributed to the Scattered Spider cluster (aka 0ktapus/Muddled Libra) compromised 9,931 accounts in over 130 organizations [43, 23].
-This campaign cost MGM Resorts alone an estimated \$100 million [79].
-To prevent this class of attacks, organizations employ technical controls (e.g., spam filters, attachment analyzers) complemented by cybersecurity awareness training [58, 48].
-Despite widespread organizational investment in anti-phishing training, mounting empirical evidence (and employee sentiment [78]) questions its effectiveness.
-Large-scale studies across multiple sectors demonstrate minimal training impact [48, 53, 54].
+Past work has demonstrated the ineffectiveness of training, but reproduction across different organizations, training approaches, and with a standardized threat assessment will help the generalizability of this phenomenon.
+To contribute to evidence-based cybersecurity policy, we conducted a large-scale reproduction study (NN=12,511) at a US-based financial technology firm.
+Our design refined prior work by comparing training modalities in operational environments, applying NIST’s standardized phishing difficulty measurement, and introducing novel organizational-level temporal resilience metrics.
+Echoing prior work, training interventions showed no significant main effects on click rates (pp=0.450) nor reporting rates (pp=0.417), with negligible effect sizes.
+However, we found that the NIST Phish Scale predicted user behavior, with click rates increasing from 7.0% (easy lures) to 15.0% (hard lures).
+Our organizational-level resilience result was mixed: 36-55% of campaigns achieved “inoculation” patterns where reports preceded clicks, but training did not significantly improve organizational-level temporal protection.
+Our results confirm the ineffectiveness of current phishing training approaches and offer a refined study design for future work.
+copyright: none
+1. Introduction
+Phishing is a primary attack vector (greitzerExperimentalInvestigationTechnical2021, ) and regularly harms organizations through operational incidents, data breaches, supply chain attacks(valsorda-compromise-survey, ),and financial losses (alkhalilPhishingAttacksRecent2021, ).
+For, in 2022–2023 campaigns attributed to the Scattered Spider cluster compromised 9,931 accounts in over 130 organizations (groupib2022oktapus, ; cisa2023scatteredspider, ).
+This campaign cost MGM Resorts alone an estimated \$100 million (reuters2024mgm100m, ).
+To prevent this class of attacks, organizations employ technical controls (e.g., spam filters, attachment analyzers) complemented by cybersecurity awareness training (mengesContrastingSynergizingCISOs2024, ; hoUnderstandingEfficacyPhishing2025, ).
+Despite organizational and regulatory emphasis on anti-phishing training, its effectiveness is questioned by mounting empirical evidence (hoUnderstandingEfficacyPhishing2025, ; lain2022phishing, ; lain2024content, ) and employee sentiment (reevesGetRedhotPoker2021, ).
 Most pertinently, Lain et al.
-showed that phishing training did not improve security outcomes across 19,000+ employees over multiple years [53, 54].
-These findings align with theoretical work by Volkamer et al., who demonstrate fundamental limitations in human processing of security indicators [96].
-This growing body of evidence suggests that phishing training is ineffective, but we believe that further reproduction of the result is needed to inform cybersecurity policy and defense strategies [67, 89].
-[^1] Our reproduction effort differs from an exact replication, which would involve repeating identical procedures.
-Instead we focus on testing the generalizability of the phenomenon of training ineffectiveness, varying the organizational context and the training approach.
-We target two specific knowledge gaps.
+found that phishing training did not improve security outcomes across 19,000+ employees over years (lain2022phishing, ; lain2024content, ).
+These findings align with theoretical work by Volkamer et al., who show fundamental limitations in human processing of security indicators (volkamer2020making, ).
+These prior studies support the general contention that phishing training is ineffective.
+In this work we refine our understanding by addressing two specific knowledge gaps.
 First, existing studies lack standardized measures of phishing difficulty, so we do not know whether training effectiveness varies by attack sophistication.
 Second, most studies examine single training modalities rather than comparing different approaches systematically, so we do not know if some training designs are more effective than others.
-To address these limitations, we conducted a large-scale reproduction study (N=12,511N=12{,}511) at a US-based financial technology firm.
+To address these limitations, we conducted a large-scale reproduction study (NN=12,511) at a US-based financial technology firm.
+[^1] Our reproduction effort differs from replication, which would repeat identical procedures.
+Instead, we focus on testing the generalizability of the phenomenon of training ineffectiveness, varying organizational context and training approach.
 Our study design has three distinctive elements: (1) mandatory pre-phishing training using two distinct training modalities (lecture-based vs.
-interactive); (2) systematic difficulty assessment using the NIST Phish Scale [86, 7]; and (3) novel organizational temporal resilience measurement through the Organizational Inoculation Index, which measures the effectiveness of phishing training from an organizational perspective rather than an individual one.
-This design enables rigorous evaluation of training effectiveness across standardized difficulty levels while providing new insights into organizational cybersecurity resilience patterns that complement individual-focused assessments.
-Our reproduction study confirms the ineffectiveness of anti-phishing training across multiple dimensions, but adds nuance to prior results.
-Consistent with prior work, neither training modality produced statistically significant improvements in phish click reduction (pp=0.450) or phish reporting behavior (pp=0.417), with negligible effect sizes indicating minimal practical value for organizational security investment.
-However, the NIST Phish Scale successfully predicted user behavior (F(2, 12086) = 41.415, p < 0.001, $\eta^{2}$ = 0.007), providing the first large-scale validation of this standardized difficulty measure.
-[^2] While the effect size is small in statistical terms due to low base rates of phishing engagement, the practical implications are significant: click rates doubled from 7.0% to 15.0% as lure difficulty increased.
+interactive); (2) systematic difficulty assessment using the NIST Phish Scale (stevesCategorizingHumanPhishing2020, ; barrientosScalingPhishAdvancing2021, ); and (3) a novel organizational temporal resilience measurement through our Organizational Inoculation Index.
+Our design evaluates training effectiveness across standardized difficulty levels, and brings new insights into organizational cybersecurity resilience patterns to complement individual ones.
+We confirm the ineffectiveness of anti-phishing training across multiple dimensions, but adds nuance to prior results.
+Consistent with prior work, neither training modality produced statistically significant improvements in phish click reduction (pp=0.450) nor phish reporting behavior (pp=0.417), with negligible effect sizes indicating minimal practical value for individuals.
+However, the NIST Phish Scale successfully predicted user behavior (F(2, 12086)=41.415, p ¡ 0.001, $\eta^{2}$=0.007), providing the first large-scale validation of this standardized difficulty measure.
+[^2] The effect size is (statistically) small due to low base rates of phishing engagement.
+Practically, click rates doubled from 7.0% to 15.0% as lure difficulty increased.
 Beyond that, our Organizational Inoculation Index revealed substantial organizational resilience that operates independently of individual training effectiveness.
-Across all experimental conditions, 36.4% to 54.5% of template-group combinations demonstrated “inoculation” patterns where threat reports (from a security-conscious employee) preceded threat interaction by another employee.
-We summarize our main contributions below:
-  1. 1.
-    Reproduction of Training Ineffectiveness Across Multiple Modalities.
-Our results reproduce the core finding of training ineffectiveness while demonstrating this phenomenon’s generalizability across training designs.
-  2. 2.
-    First Large-Scale Validation of the NIST Phish Scale.
-We evaluate the NIST Phish Scale across 12,511 participants.
-The Phish Scale is a viable measure for assessing phishing difficulty, with predictive utility.
-  3. 3.
+Across all experimental conditions, 36%–55% of template-group combinations demonstrated “inoculation” patterns where threat reports from one employee preceded threat interaction by others.
+In summary, we contribute:
+  1. (1)
+    Reproduction of Training Ineffectiveness. We reproduce the core finding of phishing training ineffectiveness.
+  2. (2)
+    First Large-Scale Validation of NIST Phish Scale. We show the NIST Phish Scale predicts phishing lure difficulty.
+  3. (3)
     Introduction of the Organizational Inoculation Index.
-We introduce and validate a novel metric for organizational temporal security.
-We report substantial organizational resilience (36-55% inoculation rates), which operates independent of individual training effectiveness and provides new insights into collective cybersecurity behaviors.
+We introduce a novel metric for organizational temporal security.
+We find substantial organizational resilience (36-55% inoculation rates), independent of individual training effectiveness.
 Significance: Cybersecurity practice and policy must be evidence-based.
-For regulators, our work contributes to growing evidence that compliance-driven training requirements may not achieve intended security outcomes.
-For organizations, our results suggest realistic (low) expectations about training outcomes and highlight the importance of technical controls rather than human-centered defenses.
-However, we also demonstrated that collective security behaviors create measurable protective effects, emphasizing the value of feedback loops between user reports and real-time threat response.
+For regulators, our work adds more evidence that phishing training is ineffective.
+For organizations, our results suggest realistic (low) expectations about training outcomes and highlight the importance of technical controls to assist humans.
+We also found that collective security behaviors may create measurable protective effects, emphasizing the value of user/IT feedback loops for real-time threat response.
 For researchers, we validate NIST’s standardized difficulty measure and confirm the reproducibility of training ineffectiveness findings across organizational contexts.
-2 Background and Related Work
-Figure 1: Phishing defense architecture, mapping the email pathway from authentication through post-delivery response.
-Technical controls create multiple interception points prior to human interaction, and human reports provide a feedback opportunity for organizational-level threat mitigation.
-See Appendix 2 for the full control pathway in the studied organization.
-We review phishing attacks and defenses (§2.1), and then discuss prior empirical studies on phishing training (§2.2).
-2.1 Phishing Attacks and Defenses
-Phishing attacks pose substantial risks to organizations, who erect both technical and human defenses against them.
-2.1.1 Phishing Attacks
-In a phishing attack, attackers send messages (phishing lures) to targets whom they wish to persuade to act against their own interests, and the interests of their organizations.
-Attackers seek results such as sensitive information, account credentials, or the ability to control the victim’s machine [4, 34].
-Traditionally, phishing attacks are divided into two classes [25]:
-Mass Phishing Campaigns: These use malicious emails sent to thousands or millions of recipients, with common themes such as banking, shipping, and software updates [4].
-Individual success rates are low (1-3%) [3], but the massive scale makes them profitable.
-Spear Phishing and Targeted Attacks: Spear phishing targets specific individuals or organizations with personalized content from open-source intelligence (OSINT) and reconnaissance.
-Attackers mine sources such as social media to craft messages that reference real relationships and events [13].
-These tailored campaigns have higher success rates (10–20%) [14].
-Successful spear phishing attacks include the 2016 Democratic National Committee breach and widespread business email compromise (BEC) schemes [4].
-Modern Phishing: Contemporary phishing operations leverage infrastructure including compromised (“trusted”) websites and domain generation algorithms to maintain persistence and evade takedown efforts [3].
-In addition, Generative AI is enabling both mass phishing and spear phishing through targeted and fluent emails and other media [4, 80].
-LLM-generated phishing content now achieves effectiveness comparable to human experts in organizational settings [9].
-2.1.2 Automated Defenses
-Figure˜1 illustrates how these attacks flow through an organization’s email system.
-Many technical defenses (blue boxes) combat attackers’ behavior, such as SPF, DKIM, and DMARC to verify that mail comes unaltered from authorized infrastructure [51, 22, 52]; ARC for preserving authentication results across forwarding [5]; BIMI for verified logos [10, 12]; and secure email gateways (SEG) for content and header inspection and URL rewriting for time-of-click inspection [1].
-After delivery, organizations employ continuous monitoring and feedback loops to adapt controls to evolving campaigns.
-Organizations blend threat-intelligence updates with AI/ML-driven anomaly scoring and impersonation detection to surface novel lures [58].
-For example, security operations center (SOC) analysts triage user-reported messages, promote high-confidence indicators back into blocklists, and tune SEG policies based on observed false positives/negatives.
-Because adversaries adapt, detection is a dynamic process [4, 3].
-See Appendix 2 for the details of the studied organization’s email controls and how user phishing reports create feedback loops that strengthen automated defenses.
-2.1.3 Phishing Training as a Defense
-Because phishing attacks compromise these technical defenses [14, 13], the human recipients play an important role in phishing defense [30, 15, 2, 28].
-There are many approaches to training humans about phishing attacks.
-Methods range from lecture-style training involving passive videos and readings [56] to simulation feedback providing immediate feedback after interaction with phishing simulations [19, 18].
-Interactive training uses gamification or scenario-based learning to engage participants [11, 46], while just-in-time training alerts users in real time during risky behavior [36].
-Anti-phishing trainings are often included in the cybersecurity awareness trainings ubiquitous in organizations.
-Legal mandates drive this practice, e.g., HIPAA requiring security awareness training for healthcare organizations [74].
-GDPR and PCI DSS encourage or mandate similar training [76, 21].
-Industry standards and frameworks further reinforce these requirements [75], e.g., ISO 27001 calling for human resource security controls including awareness programs [35].
-2.2 Empirical Studies of Phishing Training
-Table 1: Summary of Prior Experiments on Phishing and Training Interventions.
+2. Background and Related Work
+Table 1.
+Summary of Prior Experiments on Phishing and Training Interventions.
 Studies are organized by research context (Lab vs.
@@
 Hypotheses: How these works informed the hypotheses in the present study (cf.
-Table˜4).
+Table 3).
 Work Context Population Sample size Complexity Measure Compared Training Approaches? Hypotheses
-[82] (2024) Lab University students 96 H3
-[16] (2024) Lab University students 117 (Phish Scale) H4
-[18] (2014) Real-world Employees 1,359 H2, H4
-[19] (2017) Real-world University students 150 H2, H3
-[90] (2022) Real-world Employees 119 H2
-[53] (2022) Real-world Employees 19,000+ H2
-[47] (2023) Real-world Finance employees 5,000 H2
-[54] (2024) Real-world Employees 4,554 H2, H3
-[29] (2024) Real-world Banking employees “Thousands” H2
-[48] (2025) Real-world Healthcare employees 19,500+ H2
+(schoniYouKnowWhat2024, ) (2024) Lab University students 96 H3
+(canhamNotAllVictims2024, ) (2024) Lab University students 117 (Phish Scale) H4
+(caputoGoingSpearPhishing2014, ) (2014) Real-world Employees 1,359 H2, H4
+(carellaImpactSecurityAwareness2017, ) (2017) Real-world University students 150 H2, H3
+(sumnerExaminingFactorsImpacting2022, ) (2022) Real-world Employees 119 H2
+(lain2022phishing, ) (2022) Real-world Employees 19,000+ H2
+(hillmanEvaluatingOrganizationalPhishing2023, ) (2023) Real-world Finance employees 5,000 H2
+(lain2024content, ) (2024) Real-world Employees 4,554 H2, H3
+(doingAnalysisPhishingReporting2024, ) (2024) Real-world Banking employees “Thousands” H2
+(hoUnderstandingEfficacyPhishing2025, ) (2025) Real-world Healthcare employees 19,500+ H2
 Ours Real-world Fintech org. 12,511 (Phish Scale)
-Our review and analysis of prior work is summarized in Table˜1.
+We review phishing attacks and defenses (§2.1), and prior empirical studies on phishing training (§2.2).
+2.1. Phishing Attacks and Defenses
+Phishing attacks pose substantial risks to organizations, who erect both technical and human defenses against them.
+2.1.1. Phishing Attacks
+In a phishing attack, attackers send messages (phishing lures) to targets whom they wish to persuade to act against their own interests, and the interests of their organizations.
+Attackers seek results such as sensitive information, account credentials, or the ability to control the victim’s machine (alkhalilPhishingAttacksRecent2021, ; ferreiraPersuasionHowPhishing2019, ).
+These attacks use compromised infrastructure, browser vulnerabilities(10.1145/3589334.3645535, ), look-alike domains, legitimate certificates (kim2021sooel, ) and social engineering to evade technical defenses (10.1145/3696410.3714678, ; 10.1145/3652963.3655042, ) and exploit human trust (alkhalilPhishingAttacksRecent2021, ; aleroudPhishingEnvironmentsTechniques2017, ).
+Traditionally, phishing attacks are divided into two classes (dasSoKComprehensiveReexamination2020, ): mass phishing (form emails sent to many targets, often with a phishing kit (10.1145/3696410.3714710, )) and spear phishing (custom emails sent to specific targets).
+2.1.2. Automated Defenses
+Many technical defenses combat attackers’ behavior, such as SPF, DKIM, and DMARC to verify that mail comes unaltered from authorized infrastructure (rfc7208spf2014, ; rfc6376dkim2011, ; rfc7489dmarc2015, ); ARC for preserving authentication results across forwarding (rfc8617arc2019, ); BIMI for verified logos (bimi-group, ; bimi-ietf-guidance-2025, ); and secure email gateways (SEG) for content and header inspection and URL rewriting for time-of-click inspection (msft-safe-links-2025, ).
+After delivery, organizations employ continuous monitoring and feedback loops to adapt controls to evolving campaigns (mengesContrastingSynergizingCISOs2024, ).
+For an example of the dynamics (alkhalilPhishingAttacksRecent2021, ; aleroudPhishingEnvironmentsTechniques2017, ), security operations center (SOC) analysts triage user-reported messages, promote high-confidence indicators back into blocklists, and tune SEG policies based on observed false positives/negatives.
+Appendix A describes elements of the studied organization’s email controls and feedback loops.
+2.1.3. Phishing Training as a Defense
+Because phishing attacks compromise these technical defenses (butaviciusBreachingHumanFirewall2015, ; burnsSpearPhishingBarrel2019, ), the human recipients play an important role in phishing defense (downsDecision2006, ; canfieldBetter2019, ; abroshanPhishingHappensTechnology2021, ; desoldaHuman2022, ).
+Training methods range from lecture-style training involving passive videos and readings (marshallExploringEvidenceEmail2024, ) to simulation feedback providing immediate feedback after interaction with phishing simulations (carellaImpactSecurityAwareness2017, ; caputoGoingSpearPhishing2014, ).
+Interactive training uses gamification or scenario-based learning to engage participants (bitrianGamificationWorkforceTraining2024, ; hartRiskioSeriousGame2020, ), while just-in-time training alerts users in real time during risky behavior (franklinOptimisingNudgesBoosts2019, ).
+Anti-phishing trainings are often included in the cybersecurity awareness trainings ubiquitous in organizations.
+Legal mandates drive this practice.
+For example: HIPAA requires security awareness training for healthcare organizations (hipaa-security-rule, ), and GDPR and PCI DSS encourage or mandate similar training (gdpr-article39, ; pci-dss, ).
+Industry standards reinforce these requirements (nist-csf, ), e.g., ISO 27001 calling for human security controls like awareness programs (iso-27001, ).
+2.2. Empirical Studies of Phishing Training
+Our review and analysis of prior work is summarized in Table 1.
 We contribute to the body of knowledge on phishing training by conducting another large-scale study situated in a real-world context (table columns 2-4), incorporating a complexity measure (column 5), and a comparison of multiple training approaches (column 6).
@@
 The remainder of this section describes these contributions.
-For general surveys of knowledge about human factors in phishing training, we refer the reader to Marshall et al.
-[56] and Franz et al.
-[37].
+For general surveys of knowledge about human factors in phishing training, we refer the reader to (marshallExploringEvidenceEmail2024, ; franzSoK, ).
 Lab vs.
 Real-world Studies: Many early phishing training studies were conducted in controlled laboratory settings, where researchers benefit from a high degree of experimental control.
-Studies of this sort have provided numerous insights, including that active learning and immediate feedback can improve short-term detection rates, and have explored how individual differences affect training outcomes [90, 56].
+Studies of this sort have provided numerous insights, including that active learning and immediate feedback can improve short-term detection rates, and have explored how individual differences affect training outcomes (sumnerExaminingFactorsImpacting2022, ; marshallExploringEvidenceEmail2024, ).
 However, as noted by Franz et al.
-[37], these controlled settings may lack ecological validity.
+(franzSoK, ), these controlled settings may lack ecological validity.
 In laboratory environments, participant samples are typically small and consist of homogeneous groups such as academics and students.
@@
 For example, Doing et al.
-[29] analyzed phishing behaviors in bank employees, and Ho et al.
-[48] studied healthcare employees.
+(doingAnalysisPhishingReporting2024, ) analyzed phishing behaviors in bank employees, and Ho et al.
+(hoUnderstandingEfficacyPhishing2025, ) studied healthcare employees.
 Lain et al.
-has also conducted large-scale real-world measurements of training effectiveness [53, 54] Our study echoes some of these prior results in a different organizational context.
+has also conducted large-scale real-world measurements of training effectiveness (lain2022phishing, ; lain2024content, ) Our study echoes some of these prior results in a different organizational context.
 Controlling for Lure Complexity: These studies consistently report little effectiveness of anti-phishing training.
-These works have not, however, incorporated a standardized method for assessing the difficulty of the phishing lures they use [48, 29].
-We contribute the first real-world measurement of the effect of lure complexity on subjects’ interaction behaviors, based in NIST’s Phish Scale.
-To assess phishing lure complexity, the NIST Phish Scale categorizes them along two dimensions (Table˜2).
-Phishing Cues represent observable errors or inconsistencies in the email that might alert users to its fraudulent nature, such as spelling mistakes, suspicious URLs, or formatting irregularities [16].
-Premise Alignment measures the relevance and familiarity of the email’s subject matter to the recipient’s typical email environment and organizational context [6].
-Higher alignment messages with fewer cues are the most deceptive [14].
-The premise of Phish Scale is to allow organizations to benchmark lure complexity and contextualize simulation outcomes [17].
-Barrientos et al.
-[8] conducted early validation work on the NIST Phish Scale, while Dawkins and Jacobs [27] provided the official implementation guidelines.
-Ours is the first study to validate the Phish Scale’s premise at scale.
+These works have not, however, incorporated a standardized method for assessing the difficulty of the phishing lures they use (hoUnderstandingEfficacyPhishing2025, ; doingAnalysisPhishingReporting2024, ).
+The US NIST has proposed a scale for this purpose (stevesCategorizingHumanPhishing2020, ; dawkinsScalingNISTPhish2023, ).
+We summarize it here and give more details in Appendix B.
+The NIST Phish Scale categorizes lure difficulty along two dimensions.
+Phishing Cues represent observable errors or inconsistencies in the email that might alert users to its fraudulent nature, such as spelling mistakes, suspicious URLs, or formatting irregularities (canhamNotAllVictims2024, ).
+Premise Alignment measures the relevance of the email to the recipient’s organizational context (antoniouSystematicMethodExecute2015, ).
+Higher alignment messages with fewer cues are the most deceptive (butaviciusBreachingHumanFirewall2015, ).
+The premise of Phish Scale is to allow organizations to benchmark lure complexity and contextualize simulation outcomes (canhamPhishDerbyShoring2022, ).
+Building on early validation work (barrientosScalingPhishAdvancing2021, ; dawkinsScalingNISTPhish2023, ), we contribute the first real-world measurement of the effect of lure complexity on subjects’ interaction behaviors.
 Training Approaches: Franz et al.
-[37] taxonomize five categories of phishing interventions: training, awareness, education, notifications, and design interventions.
-Most studies have focused on training, which can be divided into lecture-style annual training and embedded training (e.g., through ongoing phishing simulations sent by the organization’s security team).
-Like some prior studies, our study performs a controlled experiment to compare the effectiveness of two forms of phishing training.
-Table 2: NIST Phish Scale for Detection Difficulty [87, 27].
-Cues \ Alignment Weak Medium Strong
-Few Medium Hard Hard
-Some Easy-to-Medium Medium Hard
-Many Easy Medium Medium
-This study addresses those gaps by applying the NIST Phish Scale [27] to evaluate training effectiveness across varied phishing difficulty levels in an enterprise deployment with over 12,000 participants.
-In this study, our objectives are to: (1) operationalize the NIST framework at scale, (2) examine how difficulty influences detection and reporting across training modalities, (3) provide comparative evidence on lecture-based versus interactive approaches, and (4) offer guidance for large-scale awareness efforts under conditions of modest effectiveness [48].
+(franzSoK, ) taxonomize five categories of phishing interventions: training, awareness, education, notifications, and design interventions.
+Most studies have focused on training, which can be divided into lecture-style annual training and embedded training (e.g., via ongoing phishing simulations from the organization’s security team).
+Like a few prior studies (caputoGoingSpearPhishing2014, ; carellaImpactSecurityAwareness2017, ; lain2024content, ), we compare the effectiveness of two forms of phishing training.
+This study addresses those gaps by applying the NIST Phish Scale (dawkinsScalingNISTPhish2023, ) to evaluate training effectiveness across varied phishing difficulty levels in an enterprise deployment with over 12,000 participants.
+In this study, our objectives are to: (1) operationalize the NIST framework at scale, (2) examine how difficulty influences detection and reporting across training modalities, (3) provide comparative evidence on lecture-based versus interactive approaches, and (4) offer guidance for large-scale awareness efforts under conditions of modest effectiveness (hoUnderstandingEfficacyPhishing2025, ).
 This enables evidence-based decision-making on phishing defense strategies that account for both training design and message complexity.
-3 Methodology
+3. Methodology
+Figure 1.
+Illustration of our methodology.
+Our two-factor design randomly assigns different trainings to different subjects, and then randomly sends phishing lures of varying complexity to the subjects.
+We compared between-subject performance using a mix of standard and novel metrics to test our five hypotheses.
+:
 To summarize the preceding analysis: phishing training is widely practiced, though recent evidence sheds doubt on its utility.
 Few studies have rigorously assessed outcomes against standardized difficulty measures at sufficient scale for robust inference.
-We provide a reproduction of this observation, tested through hypotheses derived from prior works.
-In this section, we describe our methodology: study design, subjects, hypotheses, metrics, and analysis methods. For ethics considerations, see the associated appendix.
-3.1 Study Design
-Figure 2: Illustration of our methodology.
-Our two-factor design randomly assigns different trainings to different subjects, and then randomly sends phishing lures of varying complexity to the subjects.
-We compared between-subject performance using a mix of standard and novel metrics to test our five hypotheses.
-Our study design is illustrated in Figure˜2.
-We employed a two-factor between-subjects experimental design [47].
+We aim to reproduce, testing hypotheses derived from prior works.
+3.1. Study Design
+Our study design is illustrated in Figure 1.
+We employed a two-factor between-subjects experimental design (hillmanEvaluatingOrganizationalPhishing2023, ).
 Subjects were randomly assigned to one of three experimental conditions using our organization’s training management system.
@@
 We quantitatively compared subjects’ performance in handling these phishing emails between the treatments.
-Factor 1—Training: The research team collaborated with a cybersecurity training vendor who produces comprehensive security awareness training materials, phishing simulation exercises designed to augment those materials, and implementation assistance for organizational deployment.
+Factor 1: Training: The research team collaborated with a cybersecurity training vendor who produces comprehensive security awareness training materials, phishing simulation exercises designed to augment those materials, and implementation assistance for organizational deployment.
 This partnership enabled access to professional training content and established phishing templates.
@@
 The first approach, Treatment A, consists of lecture-based videos on the dangers of various social engineering attacks, including phishing recognition tactics, social engineering techniques, and organizational security protocols.
-Figure˜3 gives a screenshot.
 Comprehension is assessed through a quiz.
 The second approach, Treatment B, consists of those videos followed by a series of interactive phishing exercises that present users with samples from a library of approximately 20 phishing templates.
-These exercises guide users through identifying suspicious elements and security deficiencies in emails that roughly correspond to the cues assessed by the NIST Phish Scale, and reinforces appropriate response behaviors such as the “Report Phish” button in the email client [19].
+These exercises guide users through identifying suspicious elements and security deficiencies in emails that roughly correspond to the cues assessed by the NIST Phish Scale, and reinforces appropriate response behaviors such as the “Report Phish” button in the email client (carellaImpactSecurityAwareness2017, ).
 Finally, we included a Control Group whose members received no cybersecurity training prior to the phishing simulation phase.
 This group was included to establish baseline susceptibility measurements.
-[^3] Control group participants received the same training materials after completion of the study to ensure organizational compliance with all cybersecurity awareness requirements.
+[^3] Control group participants received the same training materials after completion of the study to ensure organizational compliance with cybersecurity awareness requirements.
 The study timing dictated a small control group, as ACME Corp.
-required that the majority of employees receive annual cybersecurity training for compliance purposes.
-Figure 3: Training for Treatment A: Videos and Quiz
-Factor 2—Phishing Lure Complexity: After training, all subjects received one simulated phishing email, of varying complexity.
-To establish controlled and consistent evaluation of participant responses, we systematically assessed and categorized the difficulty of each simulated phishing email using a structured expert review process informed by the NIST Phish Scale framework [27].
+required that most employees receive annual cybersecurity training for compliance purposes.
+Factor 2: Phishing Lure Complexity: After training, all subjects received one simulated phishing email, of varying complexity.
+To establish controlled and consistent evaluation of participant responses, we systematically assessed and categorized the difficulty of each simulated phishing email using a structured expert review process informed by the NIST Phish Scale framework (dawkinsScalingNISTPhish2023, ).
 Phishing templates were selected from the vendor’s library.
 We evaluated these templates using the official NIST Phish Scale User Guide methodology to ensure accurate difficulty classification.
-For the Phishing Cues dimension, raters assessed observable indicators that might alert users to fraudulent content, including spelling errors, suspicious URLs, formatting inconsistencies, and sender authenticity markers [16].
+For the Phishing Cues dimension, raters assessed observable indicators that might alert users to fraudulent content, including spelling errors, suspicious URLs, formatting inconsistencies, and sender authenticity markers (canhamNotAllVictims2024, ).
 For the Premise Alignment dimension, raters evaluated how well each email’s subject matter, sender, and content aligned with typical organizational communications that ACME Corp.
-employees receive in their regular workflow [2].
+employees receive in their regular workflow (abroshanPhishingHappensTechnology2021, ).
 The assessment team comprised three members of ACME Corp.’s IT security staff, each experienced in identifying real-world phishing attacks.
-Following Downs et al.’s advice, two team members independently evaluated each phishing email template [30].
+Following Downs et al.’s advice, two team members independently evaluated each phishing email template (downsDecision2006, ).
 Any disagreements regarding difficulty classification were discussed and resolved, incorporating a third rater as needed.
-This collaborative validation process ensured that our phishing simulation represented a controlled range of difficulty levels essential for evaluating training effectiveness across standardized complexity measures [8].
+This collaborative validation process ensured that our phishing simulation represented a controlled range of difficulty levels essential for evaluating training effectiveness across standardized complexity measures (barrientosScalingPhishAdvancing2021, ).
 Our final simulation campaign used nineteen phishing templates that we divided into Easy (high cues/low alignment — 5721 emails), Medium (some cues, some alignment — 2279 emails) and Hard (few cues/high alignment — 4511 emails).
-Examples are given in Figure˜4.
-A full summary is given in Table˜6 in the Appendix.
-(a) Easy: low premise, high cues.
-(b) Hard: high premise, low cues.
-Figure 4: Two Example Phishing Templates.
+A full summary is given in Table 6 in Appendix C.
 Simulation Deployment: Phishing simulations were distributed among participants within a three-month window after training completion.
@@
 Templates were randomly assigned to ensure balanced distribution across training conditions.
-3.2 Subjects
-Study Context: The subjects were employees at a U.S.-based international financial technology (“fintech”) firm.
-Participants represented a diverse cross-section of the organization, including employees from finance, technology, operations, customer service, and administrative functions.
-All participants were full-time employees with active email accounts who regularly use organizational communication systems.
-We excluded participants who: (1) had inactive email accounts during the simulation period, (2) were on extended leave during the study period.
-Additionally, participants who received phishing simulations but had technical delivery failures (bounced emails, system errors) were excluded.
+3.2. Subjects
+Study Context: The subjects were full-time employees at a U.S.-based international financial technology (“fintech”) firm.
+Subjects represented a diverse cross-section of the organization, including employees from finance, technology, operations, customer service, and administrative functions.
 Subject Recruitment: This study leveraged organizational security training data collected as part of routine cybersecurity awareness training.
-The training was introduced through official organizational communications emphasizing its importance for maintaining security compliance and protecting organizational assets.
-This training was mandatory due to the organization’s need for regulatory compliance (cf.
-Appendix˜2.1.3).
+The training was introduced in official organizational communications emphasizing its importance for maintaining security compliance and protecting organizational assets.
+This training was mandatory due to the organization’s regulatory environment.
 Data were collected from Dec.
@@
 Randomization: Subjects were randomly assigned to each treatment.
-To mitigate bias by organizational role or function, we used stratified sampling [91] across departments to ensure representative distributions in each treatment group.
+To mitigate bias by organizational role or function, we used stratified sampling (sutterAvoidingHookInfluential2022, ) across departments to ensure representative distributions in each treatment group.
 Completion Rates: Our data came from the 12,511 participants who completed the full experimental protocol.
-Table˜3 shows the completion rates by treatment type.
-The comparable completion rates across conditions (ranging from 87.3% to 89.1%) suggest that our random sample was preserved between recruitment and completion.
-Table 3: Recruitment pool and completion rates by condition. Note the comparable drop-out rates by condition.
+Table 2 shows the completion rates by treatment type.
+Comparable completion rates across conditions (ranging from 87.3% to 89.1%) suggest our random sample was preserved between recruitment and completion.
+Table 2. Recruitment pool and completion rates by condition.
 Condition Recruited Completed
@@
 Total 14,051 12,511 (89.0%)
-Table 4: Summary of study hypotheses, their theoretical grounding, and relation to prior work.
+Table 3.
+Summary of study hypotheses, their theoretical grounding, and relation to prior work.
 The hypotheses address the effects of phishing email difficulty, training, and training modality on detection and reporting behaviors.
 Hypothesis Theoretical and Empirical Basis Relation to Prior Work
+H1–Phish Scale Effect: Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will have higher click-through rates than less difficult ones.
+Grounded in the foundational NIST Phish Scale framework (dawkinsScalingNISTPhish2023, ) and validation studies (barrientosScalingPhishAdvancing2021, ).
+Supported by phishing cue research (canhamNotAllVictims2024, ) and systematic email analysis (wrightPhishingSusceptibilityContext2023, ).
+The NIST Phish Scale has been validated in lab settings (barrientosScalingPhishAdvancing2021, ).
+We provide the first large-scale enterprise validation.
+H2–Training Impact (General): Training will increase the detection and reporting of phishing emails.
+Evidence of training efficacy, reduced click rates (carellaImpactSecurityAwareness2017, ), improved detection accuracy (sumnerExaminingFactorsImpacting2022, ), and short-term benefits (marshallExploringEvidenceEmail2024, ).
+Disputed (hoUnderstandingEfficacyPhishing2025, ).
+Tests whether training effects from small-scale studies (carellaImpactSecurityAwareness2017, ; sumnerExaminingFactorsImpacting2022, ) replicate at scale, addressing Franz et al.’s call (franzSoK, ).
+H3–Interactive Training Effect (Reporting): Interactive components yield higher reporting rates.
+Per Franz et al.’s analysis (franzSoK, ), and evidence of effectiveness of active and feedback-rich training (carellaImpactSecurityAwareness2017, ; marshallExploringEvidenceEmail2024, ; schoniYouKnowWhat2024, ).
+Lab findings show interactive training benefits (carellaImpactSecurityAwareness2017, ; marshallExploringEvidenceEmail2024, ).
+We provide the first test at organizational scale.
+H4–Interaction Effect: There will be an interaction between training modality and phishing difficulty, with interactive training more beneficial for easier lures.
+Emerges from cognitive load and decision strategy research (downsDecision2006, ), difficulty-aware detection patterns (wrightPhishingSusceptibilityContext2023, ), and findings on human limitations (abroshanPhishingHappensTechnology2021, ).
+Addresses the gap identified by Franz et al.
+(franzSoK, ) regarding interaction effects between training and phishing complexity, largely unexplored empirically.
+H5–Reporting Timeliness: Trained users will report suspicious emails more quickly than untrained users.
+Grounded in studies showing that training enhances threat recognition speed (carellaImpactSecurityAwareness2017, ).
+Supported by evidence that security education improves decision-making efficiency (sumnerExaminingFactorsImpacting2022, ).
+While individual reporting speed has been measured (lain2022phishing, ), no studies have examined training effects on organizational-level reporting timeliness (stevesCategorizingHumanPhishing2020, ).
+3.3. Hypotheses
+Based on prior work on phishing training efficacy (carellaImpactSecurityAwareness2017, ; caputoGoingSpearPhishing2014, ; haneyComplianceImpactTracing2024, ; sumnerExaminingFactorsImpacting2022, ; franzSoK, ; hoUnderstandingEfficacyPhishing2025, ) and documented differences in risk perception (abroshanPhishingHappensTechnology2021, ; doingAnalysisPhishingReporting2024, ), we formulated and tested five hypotheses.
+Table 3 states each hypothesis, its basis, and its relation to prior work.
+We hypothesize that:
+  1. (1)
     H1 (Phish Scale Effect): Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will have higher click-through rates than less difficult ones.
-Grounded in the foundational NIST Phish Scale framework [27] and validation studies [8].
-Supported by phishing cue research [16] and systematic email analysis [99].
-The NIST Phish Scale has been validated in lab settings [8].
-We provide the first large-scale enterprise validation.
+  2. (2)
     H2 (Training Impact–General): Training will increase the detection and reporting of phishing emails.
-Evidence of training efficacy, reduced click rates [19], improved detection accuracy [90], and short-term benefits [56].
-Disputed [48].
-Tests whether training effects from small-scale studies [19, 90] replicate at scale, addressing Franz et al.’s call [37].
-H3 (Interactive Training Effect–Reporting): Interactive components yield higher reporting rates than lecture-only.
-From Franz et al.’s modality taxonomy [37], and evidence of the effectiveness of active and feedback-rich training [19, 56, 82].
-Builds on academic findings of interactive training benefits [19, 56].
-We provide the first test at organizational scale.
+  3. (3)
+    H3 (Interactive Training Effect–Reporting): Interactive components yield higher reporting rates than lecture-only training.
+  4. (4)
     H4 (Interaction Effect): There will be an interaction between training modality and phishing difficulty, with interactive training more beneficial for easier lures.
-Emerges from cognitive load and decision strategy research [30], difficulty-aware detection patterns [99], and findings on human limitations [2].
-Addresses the gap identified by Franz et al.
-[37] regarding interaction effects between training and phishing complexity, largely unexplored empirically.
+  5. (5)
     H5 (Reporting Timeliness): Training will improve the timeliness of phishing reporting, with trained users reporting suspicious emails more quickly than untrained users.
-Grounded in studies showing that training enhances threat recognition speed [19].
-Supported by evidence that security education improves decision-making efficiency [90].
-While individual reporting speed has been measured [53], no studies have examined training effects on organizational-level reporting timeliness [88].
-3.3 Hypotheses
-Based on prior work on phishing training efficacy [19, 18, 44, 90, 37, 48] and documented differences in risk perception [2, 29], we formulated and tested five hypotheses.
-Table˜4 states each hypothesis, its basis, and its relation to prior work.
-Our specific hypotheses are:
-  1. 1.
-    H1 (Phish Scale Effect): Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will have higher click-through rates than less difficult ones.
-  2. 2.
-    H2 (Training Impact–General): Training will increase the detection and reporting of phishing emails.
-  3. 3.
-    H3 (Interactive Training Effect–Reporting): Interactive components yield higher reporting rates than lecture-only training.
-  4. 4.
-    H4 (Interaction Effect): There will be an interaction between training modality and phishing difficulty, with interactive training more beneficial for easier lures.
-  5. 5.
-    H5 (Reporting Timeliness): Training will improve the timeliness of phishing reporting, with trained users reporting suspicious emails more quickly than untrained users.
-With respect to hypothesis novelty and prior testing: H1 has been tested in small academic settings [8] but never at enterprise scale.
-H2 and H3 build on established training effectiveness research [19, 90, 56] but applies these findings in a large-scale operational context using a standardized difficulty measure; H4 is novel, addressing the research gap identified by Franz et al.
-[37] regarding interactions between training modality and phishing complexity, which has received minimal empirical attention in prior work.
-H5 is also novel, directly addressing Steves et al.’s research question about organizational temporal patterns [88].
-3.4 Metrics
-Our outcome measures were a mix of standard metrics for comparison to prior results [19, 18], and novel metrics related to response timeliness per Steves et al.
-[88].
+With respect to hypothesis novelty and prior testing: H1 has been tested in small academic settings (barrientosScalingPhishAdvancing2021, ) but never at enterprise scale.
+H2 and H3 build on established training effectiveness research (carellaImpactSecurityAwareness2017, ; sumnerExaminingFactorsImpacting2022, ; marshallExploringEvidenceEmail2024, ) but applies these findings in a large-scale operational context using a standardized difficulty measure; H4 is novel, addressing the research gap identified by Franz et al.
+(franzSoK, ) regarding interactions between training modality and phishing complexity, which has received minimal empirical attention in prior work.
+H5 is also novel, directly addressing Steves et al.’s research question about organizational temporal patterns (stevesCategorizingHumanPhishing2020, ).
+3.4. Metrics
+Our outcome measures were a mix of standard metrics for comparison to prior results (carellaImpactSecurityAwareness2017, ; caputoGoingSpearPhishing2014, ; lopezaguilarPhishingVulnerabilityPersonality2025, ) and novel metrics related to response timeliness per Steves et al.
+(stevesCategorizingHumanPhishing2020, ).
 We performed measurements using the vendor’s simulation platform, which reported the standard metrics directly and timing information that we used to derive the novel metrics.
-Standard Metrics: We employed three standard behavioral metrics to assess training effectiveness [19, 18]:
+Standard Metrics: We employed three standard behavioral metrics to assess training effectiveness (carellaImpactSecurityAwareness2017, ; caputoGoingSpearPhishing2014, ):
   • Open Rate is the percent of recipients who open the simulated phishing email, measured via embedded tracking pixels.
-[^4] This is the standard definition, though we acknowledge it can be influenced by email client settings and security configurations [6].
-  • Click-Through Rate (CTR) is the percent of recipients who click on malicious links within the phishing emails. CTR is our primary indicator of phishing susceptibility [90].
-  • Reporting Rate is the percent of recipients who report the suspicious email to the organization’s security team using the integrated “Report Phish” button, representing proactive security behavior and successful threat recognition [29].
+[^4] This is the standard definition, though we acknowledge it can be influenced by email client settings and security configurations (antoniouSystematicMethodExecute2015, ).
+  • Click-Through Rate (CTR) is the percent of recipients who click on malicious links within the phishing emails.
+CTR is our primary indicator of phishing susceptibility (sumnerExaminingFactorsImpacting2022, ).
+  • Reporting Rate is the percent of recipients who report the suspicious email to the organization’s security team using the integrated “Report Phish” button, representing proactive security behavior and successful threat recognition (doingAnalysisPhishingReporting2024, ).
 Novel Temporal Metrics: Abroshan et al.
-suggested examining temporal patterns in user responses [2], and Steves et al.
-asked: “Is time to first report sooner than time to first click?” [88].
-Rapid reporting of phish attempts enables the feedback loop built into modern email defenses (Figure˜1).
+suggested examining temporal patterns in user responses (abroshanPhishingHappensTechnology2021, ), and Steves et al.
+asked: “Is time to first report sooner than time to first click?” (stevesCategorizingHumanPhishing2020, ).
+Rapid reporting of phish attempts enables the feedback loop built into modern email defenses (Figure 3).
 We introduce the following metrics to see organizational-level temporal security dynamics:
@@
     where $t_{\text{report},i}$ is the timestamp when user $i$ reported the email and $t_{\text{deployment}}$ is when the phishing campaign was launched.
-  • Campaign Inoculation Status determines whether an organization can mitigate a phishing campaign before exploitation occurs:
+  • Campaign Inoculation Status determines if an organization can mitigate a phishing campaign before exploitation occurs:
     $\text{Inoculation}=\begin{cases}1&\text{if }t_{\text{first report}}<t_{\text{first click}}\\ 0&\text{otherwise}\end{cases}$
     where $t_{\text{first report}}$ and $t_{\text{first click}}$ are the timestamps of the first report and first click for a given campaign, respectively.
-  • Organizational Inoculation Index quantifies the temporal advantage (or disadvantage) of reporting over clicking:
+  • Organizational Inoculation Index (OII) quantifies the temporal advantage (or disadvantage) of reporting over clicking:
     $\text{OII}=t_{\text{first click}}-t_{\text{first report}}$
-    Positive values indicate “inoculation” where reports preceded clicks, enabling organizational threat mitigation before exploitation.
-Negative values indicate that clicks occurred before reports, suggesting vulnerability windows.
-This novel organizational-level metric directly addresses Steves et al.’s research question while providing actionable insights into collective security behaviors that complement individual-focused assessments.
-The Organizational Inoculation Index enables organizations to evaluate whether their workforce’s collective reporting behavior provides sufficient early warning to enable proactive threat response.
-3.5 Analysis
+    Positive values indicate “inoculation” where reports preceded clicks, enabling organizational threat mitigation before exploitation. Negative values instead suggest vulnerability windows.
+This novel organizational-level metric directly addresses Steves et al.’s (stevesCategorizingHumanPhishing2020, ) research question while providing actionable insights into collective security behaviors that complement individual-focused assessments.
+The Organizational Inoculation Index lets organizations evaluate whether their workforce’s collective reporting behavior provides early warning to enable proactive threat response.
+3.5. Analysis
 Data Validation and Quality Assurance: Prior to analysis, we conducted data validation to ensure analytical integrity.
-This included verification of random assignment effectiveness across groups, assessment of technical delivery success rates, and examination of potential outliers or data anomalies.
-Participants with incomplete data due to technical failures, extended leave, or incomplete training were excluded from analysis to maintain dataset quality [8].
+This included verifying random assignment effectiveness across groups, assessing technical delivery success rates, and examining potential outliers or data anomalies (barrientosScalingPhishAdvancing2021, ).
 Statistical Approach: We employed two complementary analytical approaches to examine training effectiveness.
 Our primary analysis used two-way analysis of variance (ANOVA) to test main effects and interactions between training modality (Treatments A, B, and Control) and phishing difficulty level (Easy, Medium, Hard) on our outcome measures.
-Note that although ANOVA with binary outcomes violates standard normality assumptions, this approach aligns with established practices in phishing research [19, 18] and enables direct comparison with prior studies.
-To address the limitations of ANOVA with binary data, we conducted confirmatory analyses using logistic regression models appropriate for binary outcomes (clicked/not, reported/not) [83].
+Note that although ANOVA with binary outcomes violates standard normality assumptions, this approach aligns with established practices in phishing research (carellaImpactSecurityAwareness2017, ; caputoGoingSpearPhishing2014, ) and enables direct comparison with prior studies.
+To address the limitations of ANOVA with binary data, we conducted confirmatory analyses using logistic regression models appropriate for binary outcomes (clicked/not, reported/not) (seabold2010statsmodels, ).
 The logistic regression models employed maximum likelihood estimation and provided odds ratios for effect size interpretation.
 Both analytical approaches yielded consistent conclusions regarding training effectiveness, with ANOVA providing F-statistics for comparison with prior literature and logistic regression offering more statistically appropriate modeling of our binary outcomes.
-We report both approaches to provide comprehensive statistical evidence while acknowledging the methodological trade-offs inherent in each approach.
-All analyses were conducted using Python [77] with pandas [57] for data manipulation, NumPy [45] for numerical computations, SciPy [95] for statistical tests, and statsmodels [83] for regression modeling and maximum likelihood estimation.
-Visualizations were created using matplotlib [50] and seaborn [97].
-3.6 Threats to Validity
-We encourage the reader to consider several limitations when interpreting the results of our study.
-We distinguish these limitations into three kinds of threats to validity.
-Following the guidance of Verdecchia et al., we focus on what we perceive to be the most crucial ones [94].
-Construct Validity: Construct validity is concerned with whether we operationalized the desired constructs of our study appropriately [84, 98].
-In this study, we worked with two primary constructs: (1) the particular interpretations of the two different types of phishing training we deployed, which depend on our cybersecurity training vendor; and (2) our assessment of the difficulty of phishing lures.
-It is possible that our cybersecurity vendor’s training was simply lackluster, hence its failure to produce any significant improvement in subject performance.
-However, note that our results align with prior work that used different trainings.
-Meanwhile, to control for subjectivity in rating phishing lures, we applied the NIST Phish Scale and used multiple independent analysts to judge difficulty [85].
-Internal Validity: Internal validity threats affect the reliability of conclusions we have about the relationship between the cause and the effect we observe [84, 98].
+We report both approaches, while acknowledging the methodological trade-offs inherent in each approach.
+3.6. Threats to Validity
+We encourage the reader to consider several limitations when interpreting the results of our study (verdecchia2023threats, ). We distinguish these limitations into three kinds of threats to validity.
+Construct Validity: Construct validity is concerned with whether we operationalized the desired constructs of our study appropriately (shadishCookCampbell2002, ; wohlinEtAl2012, ).
+In this study, we worked with two primary constructs: (1) the phishing training we deployed, which depends on our cybersecurity training vendor; and (2) our assessment of the difficulty of phishing lures.
+First, though vendors may vary in training design, our cybersecurity vendor’s training is regulation-compliant, and our results align with prior work that used different trainings.
+Second, to control for subjectivity in rating phishing lures, we used the NIST Phish Scale with multiple analysts to judge difficulty (sjobergBergersen2023, ).
+Internal Validity: Internal validity threats affect the reliability of conclusions we have about the relationship between the cause and the effect we observe (shadishCookCampbell2002, ; wohlinEtAl2012, ).
 For our study, we used stratified random sampling to mitigate potential biases related to job role and other organizational issues.
@@
 Organizational constraints required us to limit the sample size for our control group.
-This group was still large enough to provide a baseline.
+This group was still large enough to be a baseline.
 Like all real-world phishing measurement studies, our work faces several measurement concerns.
-False positive clicks may occur when automated security systems (security gateways, antivirus, etc.), shadow IT applications, or personally installed tools / browser extensions automatically scan or prefetch URLs from emails, generating click events without actual user interaction.
-False negative reporting may also occur, wherein users report phishing emails through alternative channels, such as contacting the IT helpdesk directly, rather than the simulation platform’s designated “Report Phish” button.
+False positive clicks may occur when automated security systems (security gateways, antivirus, etc.), “shadow IT” applications, or personally installed tools / browser extensions automatically scan or prefetch URLs from emails, generating click events without actual user interaction.
+False negative reporting may also occur, in which users report phishing emails through alternative channels, such as contacting the IT help-desk directly, rather than the simulation platform’s designated “Report Phish” button.
 These behaviors would, respectively, incorrectly increase and decrease our measurements of subject interactions.
 We cannot fully eliminate such threats, but our large sample size mitigates their impact on our overall conclusions.
-External Validity: External threats impact the generalizability of our research [81, 100].
+Although such measurement noise is inherent in operational studies, the large sample size and balanced distribution across conditions mitigate its systematic influence on our results.
+Every effort was made to standardize user behavior, including explicit instruction during both training modalities on how to identify and report suspicious messages through the organization’s approved “Report Phish” mechanism.
+Future work might explicitly quantify these instrumentation effects by correlating simulation telemetry with downstream SOC logs or ticketing data to better understand their prevalence and direction of bias.
+External Validity: External threats impact the generalizability of our research (runesonHost2009, ; yin2018, ).
 Our study was situated in the USA at a fintech company.
-The fintech industry context may limit generalizability to other sectors, as financial services employees may have different baseline levels of security awareness when compared to other industries.
-However, given the minimal effect sizes measured in our study, it is unclear how large of a threat this is.
-The study’s focus on immediate post-training effects also limits our understanding of long-term training impact.
-Future research should examine whether the modest effects we observed persist over time or diminish as training recency decreases, or if they improve with volume or other dimensions.
-4 Results
-Table 5: Summary of Training Effectiveness Results.
-Average rates show that trained groups performed similarly to control group, with modest differences across conditions.
-Effect sizes are shown in Figure˜5.
+The context of the fintech industry may limit generalizability to other sectors, as financial services employees may have a higher baseline level of security awareness.
+However, given the minimal effect sizes measured in our study, this threat is not substantial.
+Our focus on immediate post-training effects also limits our understanding of long-term training impact.
+Future work might assess the effect of training recency and volume.
+3.7. Ethics
+We acknowledge the importance of informed consent as part of an empirical study.
+Employees are informed at onboarding that periodic phishing simulations are a mandatory condition of employment, consistent with industry practice in financial services organizations (doingAnalysisPhishingReporting2024, ).
+Our study then analyzed data collected as part of routine, required security awareness training, using the existing corporate platform and controls.
+Our protocol was reviewed by both the company and the partner academic institution’s IRB.
+Employees participated within the existing employment/training framework, which has been approved by the company.
+In consultation with management, we excluded high-distress scenarios (e.g., family emergency) to reduce psychological risk but preserve ecological validity.
+Because the research used de-identified operational data generated by routine commercial operations, our institutional IRB concluded that from the academic partner’s view our study did not constitute human subjects research.
+4. Results
+Table 4. Summary of Training Effectiveness Results. Average rates show that trained groups performed similarly to control group, with modest differences across conditions.
 Aspect Results
-Effect Size Generally modest, consistent with recent large-scale findings [48]
+Effect Size Consistent with large-scale findings (hoUnderstandingEfficacyPhishing2025, )
+  Training Effect $\eta^{2}=0.000$ (negligible)
+  Difficulty Effect $\eta^{2}=0.007$ (meaningful)
+  Training $\times$ Difficulty $\eta^{2}=0.001$ (negligible)
 Click Rates
@@
 Our analysis examined the effectiveness of phishing training across 12,511 participants randomly assigned to three experimental conditions.
-Table˜5 gives a summary of our key results.
-Figure˜5 complements this with a depiction of effect sizes for the controlled factors.
-Figure 5: Comparison of effect sizes ($\eta^{2}$) for key factors in phishing susceptibility.
-The NIST Phish Scale (difficulty) shows the strongest and only practically meaningful effect ($\eta^{2}=0.007$, $p<0.001$), while training and interaction effects are negligible ($\eta^{2}\leq 0.001$).
-See also Table˜5.
-4.1 H1: Phish Scale Effect
-Figure 6: Comparison of effect sizes of phishing susceptibility. Bars show the variance explained by phishing difficulty (per the NIST Phish Scale) for each training intervention.
+Table 4 gives a summary of our key results.
+Appendix 4 complements this with a depiction of effect sizes for the controlled factors.
+4.1. H1: Phish Scale Effect
+Figure 2. Comparison of effect sizes of phishing susceptibility. Bars show the variance explained by phishing difficulty (per the NIST Phish Scale) for each training intervention. :
 H1 was supported.
-Figure˜6 illustrates the comparison of effect sizes across different factors in phishing susceptibility.
+Figure 2 illustrates the comparison of effect sizes across different factors in phishing susceptibility.
 The NIST Phish Scale successfully predicted user behavior.
 Phishing lure difficulty had a highly significant effect on click-through rates (F(2, 12086)=41.415, p << 0.001, $\eta^{2}$=0.007), validating the scale’s utility for measuring phishing complexity.
-Higher difficulty emails consistently produced higher click rates across all experimental conditions:
-  • Easy difficulty: 7.0% average click rate
-  • Medium difficulty: 8.7% average click rate
-  • Hard difficulty: 15.0% average click rate
-This statistically significant difficulty gradient confirms that the NIST Phish Scale effectively captures factors that influence user susceptibility to phishing attempts, providing empirical validation for its use in organizational contexts.
-4.2 H2: Overall Training Effects
+Higher difficulty emails consistently produced higher click rates across all experimental conditions; Easy difficulty: 7.0% average click rate, Medium difficulty: 8.7% average click rate, Hard difficulty: 15.0% average click rate.
+4.2. H2: Overall Training Effects
 H2 was not supported.
@@
 Main effect analysis showed no significant difference between trained and untrained groups for click rates (F(2, 12086) = 0.800, pp=0.450) or report rates (F(2, 12086) = 0.874, pp=0.417).
-This finding aligns with recent large-scale evidence suggesting minimal effectiveness of current training approaches [48].
-4.3 H3: Interactive Training Effects
+This finding aligns with recent large-scale evidence suggesting minimal effectiveness of current training approaches (hoUnderstandingEfficacyPhishing2025, ).
+4.3. H3: Interactive Training Effects
 H3 was not supported.
@@
 The coefficient estimate for Training + Exercise relative to Training Only was modest and non-significant across all difficulty levels.
-This finding contrasts with some prior work suggesting benefits of interactive training components [19], but aligns with recent large-scale studies showing limited training effectiveness regardless of modality [48].
-4.4 H4: Training-Difficulty Interactions
+This finding contrasts with some prior work suggesting benefits of interactive training components (carellaImpactSecurityAwareness2017, ), but aligns with recent large-scale studies showing limited overall training effectiveness (hoUnderstandingEfficacyPhishing2025, ).
+4.4. H4: Training-Difficulty Interactions
 H4 was not supported.
@@
 While there was a marginal interaction effect for click rates, this primarily reflected the pattern that training appeared to increase clicks on easier emails while potentially reducing clicks on harder emails.
-Coefficient analysis revealed some interesting interaction patterns:
-  • Training Only × Hard: -9.67 percentage points (pp=0.002)
-  • Training + Exercise × Hard: -9.09 percentage points (pp=0.004)
-These significant negative interactions suggest that training may provide some protection against the most difficult phishing emails, even though overall training effects were not significant.
-It is worth considering that the training vendor created the phishing templates and training materials, which may introduce alignment between the cues emphasized in training and those present in the simulated emails.
-This alignment could artificially enhance the apparent effectiveness of training for certain email types, particularly those closely modeled after the vendor’s examples.
-As a result, the observed reductions in click rates for hard emails may partially reflect this overlap rather than a generalizable improvement in user detection skills.
-Future studies should consider using independently developed phishing templates or varying template sources to better assess the robustness of training effects across diverse phishing scenarios.
-4.5 H5: Reporting Timeliness
+Coefficient analysis revealed some interesting interaction patterns: Training Only × Hard: -9.67 percentage points (pp=0.002); Training + Exercise × Hard: -9.09 percentage points (pp=0.004) These significant negative interactions suggest that training may provide some protection against the most difficult phishing emails, even though overall training effects were not significant.
+Higher alignment messages with fewer cues are the most deceptive (butaviciusBreachingHumanFirewall2015, ).
+4.5. H5: Reporting Timeliness
 H5 was partially supported, though with important caveats about sample size limitations.
-Our analysis of reporting timeliness revealed mixed patterns in organizational-level temporal protection across treatment groups.
-Individual reporting behavior showed a median time-to-report of 21 minutes (0.35 hours), with 90% of reports occurring within 18.8 hours of phishing deployment, indicating generally rapid user response to suspicious emails.
-At the organizational level, we observed notable differences in inoculation patterns—campaigns where reports preceded clicks, enabling organizational threat mitigation before exploitation.
-Training Only achieved the highest inoculation rate at 54.5% of campaigns, with a median organizational response time of -0.33 hours (reports preceding clicks).
-In contrast, the Control group achieved inoculation in 42.9% of campaigns with a median response time of 8.09 hours, while Training + Exercise showed 36.4% inoculation with a median response time of 1.21 hours.
-However, these differences were not statistically significant due to necessarily small sample sizes at the campaign level (n=33 campaigns total).
-The Organizational Inoculation Index revealed substantial organizational resilience independent of individual training effectiveness, with over one-third of all template-group combinations demonstrating proactive reporting patterns.
-While Training Only showed numerically superior temporal performance, the limited campaign-level sample size prevents definitive conclusions about training effects on organizational-level reporting timeliness.
-These findings provide the first empirical measurement of organizational temporal security dynamics, directly addressing Steves et al.’s research question about whether "time to first report is sooner than time to first click" [88].
-The results suggest that organizational-level protective behaviors may operate independently of individual training effectiveness, though larger-scale studies are needed to establish statistical significance of these temporal patterns.
-4.6 Other Measurements
-The final question we considered was whether there was any relationship between training time and detection ability, e.g., decay of training effectiveness.
-In our study, subjects received the phishing simulation email between 0–4 months after the training.
-Figure˜7 shows the click rate by treatment group, discriminating by lure difficulty.
-We observe no discernible trend in the results as a function of time.
-Given the result of the visual inspection, we did not perform statistical tests of this aspect.
-Figure 7: Temporal Impact on Detection. We observed no discernible relationship between time-since-training and performance on phishing lures.
-4.7 Effect Sizes and Practical Significance
-While some individual coefficients reached statistical significance, effect sizes were generally small ($\eta^{2}$ << 0.01 for all main effects), raising questions about practical significance [90].
-The largest meaningful effect was the difficulty gradient ($\eta^{2}$ = 0.007 for click rates), confirming the NIST Phish Scale’s predictive validity.
-Training effects, where present, were modest in magnitude.
-Even statistically significant interaction effects represented changes of less than 10 percentage points, suggesting that while training may provide some benefits, these are limited in scope and may not justify the substantial organizational investment required for comprehensive training programs [48].
-These findings underscore the need for realistic expectations about training effectiveness and suggest that organizations should complement training with other defense mechanisms rather than relying on awareness programs as primary protection against phishing threats [37].
-5 Discussion
-Our large-scale field study provides empirical evidence for both the utility of the NIST Phish Scale and the limited effectiveness of current phishing training approaches.
-While we observed some statistically significant effects, their practical significance raises important questions about the return on investment for organizational security awareness programs.
-5.1 The Reality of Training (In)Effectiveness and Organizational Security
-Confronting Training Effectiveness Realities: Despite deploying well-structured training programs to over 12,000 employees, we found no statistically significant main effects of training on either click rates (pp=0.450) or report rates (pp=0.417).
-The small effect sizes we observed ($\eta^{2}$ < 0.01 for all main effects) indicate that even statistically significant improvements translate to minimal operational impact.
-This aligns with recent large-scale evidence from Ho et al.
-[48] and challenges the optimistic assessments often promoted by cybersecurity training vendors.
-Our findings suggest that many organizations may be operating under a dangerous illusion of training effectiveness.
-The cybersecurity community must confront this reality by demanding more rigorous evaluation standards and refusing to accept vendor claims that cannot be substantiated under realistic threat conditions.
-Only by acknowledging the true limitations of current training approaches can organizations make informed decisions about resource allocation and security strategy [37].
-The Possibility of Training Metrics Manipulation: Our findings suggest the possibility that organizations may inadvertently—or deliberately—manipulate phishing training metrics to achieve desired outcomes.
-The strong predictive power of the NIST Phish Scale (F(2, 12086) = 41.415, p < 0.001) demonstrates that practitioners can significantly influence training results simply by adjusting the difficulty of phishing lures used in simulations.
-Our observation informs broader concerns about “compliance theater” in cybersecurity [44], where organizations optimize for metrics that satisfy regulatory requirements rather than actual security improvement.
-Recent work by Ho et al.
-[48] and Lain et al.
-[53, 54] provides additional evidence that training effectiveness is often overstated.
-Our findings demonstrate a specific mechanism through which such overstatement may occur: the systematic use of phishing lures that fail to represent realistic attack sophistication.
-Such selection may not be the fault of the organizations alone.
-We note that many vendors of phishing training allow organizations to select from libraries of phishing templates without standardized difficulty assessment.
-This enables the selection—accidental or deliberate—of easier lures, resulting in inflated success metrics [18].
-5.2 Recommendations for Phish Resilience
-Layered Defense Strategy: Our findings add to the evidence that organizational phishing defense strategy must treat training as one of many layers of defense.
-The strong effect of phishing difficulty on user susceptibility (F(2,12086) = 41.415, p < 0.001) indicates that attacks are likely to succeed regardless of training efforts.
-Evidence-Based Implementation Framework: Organizations seeking meaningful security assessment must resist the temptation to achieve favorable metrics through lure selection bias.
-Our results support the following three-part framework:
-(1) Technical controls first.
-As humans cannot be trained to reliably detect phishing, automated defenses are in order.
-Prior work recommends investing primarily in filtering, attachment analysis, URL protection, and multi-factor or phishing-resistant authentication [4].
-(2) Targeted training.
-Our results suggest that generic company-wide training is ineffective [48].
-Prefer focused, role- or scenario-specific training where prior evidence suggests impact [82, 90].
-Use standardized difficulty frameworks (e.g., the NIST Phish Scale) for evaluation.
-(3) Continuous assessment. Regularly compare training costs and outcomes against alternative investments using standardized metrics that reflect behavior and resource efficiency.
-In phishing training assessment, guided by our methods, we recommend that organizations:
-  • Use standardized difficulty frameworks like the NIST Phish Scale to ensure representative threat simulation
-  • Report results separately by difficulty level rather than aggregating across lure types
-  • Benchmark against external standards rather than setting arbitrary internal targets
-  • Focus on improvement trends over time rather than absolute performance thresholds
-We do not recommend eliminating phishing training, but rather to right-size its role within a defense-in-depth strategy guided by evidence rather than marketing claims.
-The Evolving Threat Landscape: Our findings must be interpreted within the context of rapidly evolving phishing threats.
-The emergence of large language models and generative AI technologies is enabling attackers to create increasingly sophisticated and personalized phishing attempts that may render current training approaches obsolete [4, 80, 9].
-When attackers can automate the creation of personalized phishing lures and generate near-perfect replicas of legitimate communications, it may become unreasonable to expect users to reliably identify deceptive content.
-[^5] One of the authors recently failed his organization’s annual phishing training, which was experimenting with GenAI-based lures.
-The fundamental challenge lies in the inherent complexity of human factors in resistance to phishing attacks.
-Since these exploits target individual vulnerabilities, personal and contextual factors play an enormous role in individual user susceptibility [30].
-Even a single individual will vary dramatically throughout a day—users typically focus on content without paying attention to security, and are significantly impacted by cognitive constraints such as memory limitations and current cognitive load [33, 32].
-Training represents only a single layer of defense against these multifaceted human vulnerabilities.
-This technological evolution, combined with the complexity of human decision-making under cognitive load, suggests that the cybersecurity operations community must reconsider where human-centered defenses lie in the multilayered approach organizations must take to cybersecurity.
-Future approaches may need to emphasize cryptographic verification systems, zero-trust architectures, and other technical mechanisms to reduce reliance on the “human firewall”.
-Implications for Theory and Practice: Our findings have significant implications for both cybersecurity theory and organizational practice.
-The limited training effectiveness we observed suggests that current human-centered defense paradigms may be fundamentally inadequate for addressing sophisticated phishing threats.
-The modest effect sizes ($\eta^{2}$ < 0.01 for all main effects) indicate that even statistically significant improvements may not translate to meaningful operational risk reduction.
-The interaction patterns we identified—where training appeared to increase vulnerability to easier emails while potentially providing some protection against harder emails—highlight the complex and sometimes counterintuitive effects of security awareness interventions.
-These findings suggest that training programs may redistribute rather than uniformly reduce user vulnerability, creating new attack vectors while addressing others.
-For practitioners, our results support a realistic reassessment of training investments within broader security strategies.
-Organizations should view awareness programs as one component of layered defense rather than primary protection mechanisms.
-The substantial resources typically devoted to comprehensive training initiatives might achieve better risk reduction when reallocated toward technical controls, incident response capabilities, or more targeted interventions [37].
-5.3 Future Work
-Our findings open several important avenues for future research that could advance understanding of phishing training effectiveness and organizational cybersecurity resilience.
-AI-Generated Phishing and the Limits of the NIST Phish Scale: The emergence of generative AI enables highly convincing phishing attacks that may lack the flaws traditionally used to assess difficulty, such as poor grammar or implausible scenarios [80].
-This calls into question the continued relevance of the NIST Phish Scale, which assumes such imperfections are present.
-Although some phishing campaigns will continue to use low-budget approaches with human-induced errors, future research should investigate whether training methods can prepare users to recognize AI-crafted phishing attempts and whether new paradigms are needed to simulate and defend against these advanced threats.
-Automated Difficulty Assessment: The manual process of rating phishing lures using the NIST Phish Scale is subject to rater subjectivity and does not scale well to large deployments.
-Future work should explore automated or semi-automated systems—potentially powered by machine learning—to assess lure difficulty by evaluating features such as cue visibility and premise plausibility.
-Additionally, comparing lure success rates in randomized control groups with sufficiently large populations could offer a data-driven alternative for estimating difficulty.
-These approaches could improve consistency, scalability, and real-world applicability of phishing difficulty assessments.
-Alternative Training Modalities and Interventions: While our study focused on lecture-based and interactive training approaches commonly offered by vendors, other intervention types remain understudied.
-Future research should investigate the effectiveness of just-in-time training, personalized adaptive tools that adjust to users’ individual risk profiles [82], and hybrid models that combine human training with technical controls.
-Emerging interventions—such as AI-driven, real-time reviews of links and emails—also warrant exploration.
-Even widely deployed tools like spelling and grammar checkers may offer untapped potential as passive defenses against phishing.
-Organizational Inoculation Index Refinement: Our introduction of the Organizational Inoculation Index offers a starting point for measuring collective security behaviors across a workforce, but several refinements could improve its practical value.
-We suggest that the Organizational Inoculation Index or a variant might be a useful addition to large-scale phishing simulation tools, as it provides a valuable metric for assessing how quickly and effectively an organization responds to phishing threats at scale, a valuable metric worth improving upon.
-These research directions collectively aim to provide organizations with evidence-based guidance for cybersecurity defense strategies while advancing the scientific understanding of human factors in cybersecurity.
-The ultimate goal is to develop more effective approaches to organizational phishing resilience that acknowledge both the capabilities and limitations of human-centered defenses.
-6 Conclusion
-This large-scale field study of 12,511 employees provides critical empirical evidence about the effectiveness of phishing awareness training in real-world organizational contexts.
-Our systematic application of the NIST Phish Scale across three experimental conditions represents the largest validation of this standardized difficulty framework to date.
-The NIST Phish Scale demonstrated strong predictive power, with click rates ranging from 7.0% for easy lures to 15.0% for hard lures (F(2, 12086) = 41.415, p < 0.001).
-However, neither lecture-based nor interactive training produced statistically significant improvements in click rates (p = 0.450) or reporting behavior (p = 0.417), with effect sizes below 0.01 for all main training effects.
-These findings challenge conventional assumptions about training efficacy and necessitate a fundamental shift toward layered technical defenses.
-Rather than viewing training as a primary protection mechanism, organizations should integrate awareness programs as supplementary components within comprehensive security architectures that emphasize passwordless authentication, email filtering, and incident response procedures that assume successful compromise.
-The evidence suggests current training approaches provide only modest protection against evolving phishing threats.
-Ethical Considerations
-Our analysis uses operational security-training data from a U.S.
-fintech enterprise with randomized assignment to two training modalities and a (smaller, compliance-constrained) control group.
-Each participant received one simulated phish after training; difficulty was rated via the NIST Phish Scale.
-Stakeholders: Following the guidance of Davis et al.
-[26], we identify the following direct stakeholders: employees who received training and one simulated phishing email; their managers; the organization’s security team and leadership; and the training/simulation vendor.
-Indirect stakeholders are the US NIST whose Phish Scale we used, customers and the public who benefit from stronger defenses, and the research community.
-Potential Harms: Our subjects and their managers face risks such as distress from deceptive emails, reputational or employment concerns tied to phishing performance, and privacy risks from telemetry (opens, clicks, reports, timestamps).
-Given our negative results, the organization’s security and leadership teams (and comparable teams at other organizations) may be exposed to risk related to ineffective use of company funds.
-These results also expose the training/simulation vendor to the risk of lost revenue.
-Among our indirect stakeholders, we perceive no harm to business customers and the public, nor to the research community. If the NIST Phish Scale is found ineffective, NIST may face reputational harm.
-Potential organizational harms include disclosure of sensitive internal processes or phishing content that could aid adversaries.
-Publication risks center on enabling misuse if raw artifacts/templates were released without redaction.
-Mitigations Implemented: Minimal-risk, in-policy setting: The work analyzed data collected as part of routine, required security awareness training, using the existing corporate platform and controls; no extra emails beyond the single post-training simulation were sent for research purposes.
-Content safeguards: In consultation with management, we excluded high-distress scenarios (e.g., family emergency) to reduce psychological risk but preserve ecological validity.
-Fairness & access: A smaller-than-desirable control group was necessary to satisfy compliance obligations that most employees receive training; control participants received the same training after the study to avoid unequal access to potential benefits.
-Privacy & data minimization: Analyses used de-identified, minimum-necessary fields (delivery, open, click, report events; template and arm; coarse timestamps needed for the Organizational Inoculation Index).
-We report aggregate results; no identifiable outcomes are published.
-Any user-level logs remain within the organization.
-Operational safety: Simulations were limited to standard vendor templates vetted by the organization; delivery failures (bounces/system errors) were excluded; security operations continued their normal protections and triage process to avoid increased organizational risk.
-Artifact release controls: To support Open Science while preventing harm, we provide anonymized analysis code and aggregated/synthetic data sufficient to reproduce statistics and figures; raw email bodies/headers and user-level logs are not released due to privacy and security risk.
-Links are anonymized for review per the CFP and replaced with non-anonymous DOIs post-acceptance.
-Deception and Employee Welfare: The study necessarily involved deception (simulated phishing) to preserve realism.
-We mitigated risk by (i) restricting distressing content, (ii) limiting contact to a single simulation per participant, and (iii) focusing analyses and publication on aggregate, non-identifying outcomes (e.g., click, report, and campaign-level temporal metrics), not on individual performance.
-Oversight and Consent Context: The company and the academic institution’s IRB reviewed the protocol.
-Employees participated within the existing employment/training framework, which has been approved by the company.
-Because the research used de-identified operational data generated by that framework, our institutional IRB concluded that our study did not constitute human subjects research under applicable federal guidelines.
-Decision to Proceed and to Publish: We judged that expected benefits—independent, large-scale evidence about real-world training effectiveness; the first enterprise-scale validation of a standardized phishing difficulty measure; and a practical organizational temporal metric—outweigh residual risks under the above mitigations.
-We further reduce publication risk by withholding sensitive raw artifacts and releasing only what is necessary for independent verification.
-Open Science
-Our anonymized artifact is available at: https://anonymous.4open.science/r/phish-2F58/README.md.
-Artifact Availability and Reproducibility: This research fully embraces USENIX Security 2026’s Open Science requirements through comprehensive artifact sharing and complete reproducibility.
-All materials necessary for independent verification and extension are provided following conference guidelines for responsible disclosure in cybersecurity research.
-Our complete research package includes: (1) anonymized dataset (anonymized_phishing_data.csv, 12,511 participants, 2.3MB compressed) with comprehensive metadata and data dictionary; (2) complete analysis pipeline (phishing_analysis_toolkit/) enabling one-click reproduction of all paper results; and (3) statistical validation suite with automated testing and sensitivity analyses.
-All artifacts have been independently tested across Windows, macOS, and Linux environments with runtime of 10-20 minutes on standard hardware (4GB RAM minimum).
-Privacy and Ethical Compliance: Following USENIX Security’s emphasis on responsible disclosure, our data sharing employs privacy-first anonymization with k-anonymity $\geq$ 5, external privacy audit by certified data protection officer, and re-identification risk < 0.1%.
-Security-aware disclosure excludes organizational identifiers that could enable targeted attacks while presenting vulnerability information constructively for defense improvement.
-All procedures maintain GDPR Article 89 compliance and IRB approval (Protocol #ANON-2024-PHI-001).
-Complete Computational Environment: The reproducibility framework provides: platform-independent Python implementation with version-pinned dependencies (23 packages), fixed random seeds for deterministic execution, and automated validation procedures.
-Statistical analyses employ both ANOVA (for literature comparison) and logistic regression (methodologically appropriate for binary outcomes) using Python [77] with pandas [57], NumPy [45], SciPy [95], and statsmodels [83].
-Cross-verification using alternative software confirms all results.
-Documentation and Community Support: Multi-level documentation includes: README.md with complete project overview, data/data_dictionary.md defining all 47 variables, and REPLICATION_GUIDE.md for independent verification.
-The modular code architecture supports longitudinal extensions and provides standardized variable definitions for cross-study comparison.
-Post-publication materials will include institutional repository deposit with permanent DOI and ongoing community support for replication questions.
-All artifacts comply with USENIX Security 2026 requirements while maintaining the highest standards of privacy protection and responsible disclosure in cybersecurity research.
-Complete materials are provided in the submission package for reviewer evaluation, with permanent public access following publication acceptance.
+Individual reporting behavior showed a median time-to-report of 21 minutes (0.35 hours), with 90% of reports occurring within 18.8 hours of phishing deployment.
+We analyzed 17 templates that received both clicks and reports.
+Overall, 52.9% achieved inoculation (reports preceded clicks), with median Organizational Inoculation Index (OII) of +0.4 hours.
+By treatment group, Control achieved inoculation on 66.7% of templates (median OII: +0.3 hours), Training Only on 75% (median OII: +1.37 hours), and Training + Exercise on only 28.6% (median OII: -0.2 hours).
+These differences were not statistically significant due to limited sample size.
+For successfully inoculated templates, the median OII was approximately 30 minutes, suggesting automated response mechanisms would be necessary to mitigate threats before exploitation.
+These findings provide the first empirical measurement of organizational temporal security dynamics, responding to Steves et al.’s research question about whether “time to first report is sooner than time to first click” (stevesCategorizingHumanPhishing2020, ).
+The results suggest that organizational-level protective behaviors operate independently of individual training effectiveness.
+Studying the opportunities and limitations of this feedback loop is a promising direction for further study.
+4.6. Effect Sizes and Practical Significance
+While some coefficients reached statistical significance, effect sizes were generally small ($\eta^{2}$ << 0.01), raising questions about practical significance (sumnerExaminingFactorsImpacting2022, ).
+The only meaningful effect was the difficulty gradient ($\eta^{2}$ = 0.007 for click rates), confirming the NIST Phish Scale’s predictive validity.
+Even statistically significant interaction effects represented changes of less than 10 percentage points.
+These findings underscore the need for realistic expectations about training effectiveness and suggest organizations should complement training with other defense mechanisms rather than relying on awareness programs as primary protection (franzSoK, ).
+5. Discussion
+5.1. Implications for Practice
+After deploying regulation-compliant training programs to over 12,000 employees, we found no statistically significant main effects of training on either click rates (pp=0.450) or report rates (pp=0.417).
+The small effect sizes we observed ($\eta^{2}$ ¡ 0.01 for all main effects) indicate that even statistically significant improvements translate to minimal operational impact.
+This result aligns with recent large-scale evidence from Ho et al.
+(hoUnderstandingEfficacyPhishing2025, ) and challenges the optimistic assessments promoted by cybersecurity training vendors.
+The predictive power of the NIST Phish Scale (F(2, 12086)=41.415, p ¡ 0.001) demonstrates that practitioners can significantly influence training results by adjusting the difficulty of phishing lures used in simulations.
+Our data shows click rates vary dramatically based on lure difficulty: from 7.0% for easy emails to 15.0% for hard emails.
+Organizations setting specific performance targets (such as click-through rates around 2%) may achieve these by using low-difficulty lures that fail to represent sophisticated attacks organizations actually face (alkhalilPhishingAttacksRecent2021, ).
+This creates a misalignment between perceived and actual security posture.
+These findings suggest several actionable opportunities for both training vendors and organizational security teams.
+Vendors can enhance their value by incorporating standardized measures of lure difficulty and adaptive testing frameworks aligned with models such as the NIST Phish Scale.
+This would produce interpretable results that organizations could use to calibrate difficulty levels to track progress over time, personalize training to employee risk profiles, and benchmark outcomes across departments or industries.
+However, the relatively low impact of phishing training suggests that awareness programs must be part of a layered defense strategy, complemented by technical controls and incident response preparedness (franzSoK, ), to gain risk reduction rather than nominal compliance (haneyComplianceImpactTracing2024, ).
+In this way, security awareness initiatives can evolve from static, one-size-fits-all modules into data-driven, context-aware components of an organization’s continuous improvement process.
+5.2. Future Work
+Refinement of Organizational Inoculation Index: Our introduction of the Organizational Inoculation Index offers a starting point for measuring collective security behaviors across a workforce.
+We suggest that this metric might be a useful addition to large-scale phishing simulation tools, as it provides a valuable assessment of how quickly and effectively an organization responds to bulk phishing threats at scale, although its utility against spear phishing (especially in the age of LLMs) is unclear.
+AI-Generated Phishing and the Limits of the NIST Phish Scale: The emergence of generative AI enables highly convincing phishing attacks at scale that lack traditional flaws used to assess difficulty, such as unsecured websites (kim2021sooel, ) or poor grammar (royFromChatbotsPhishbots2024, ).
+This calls into question the continued relevance of the NIST Phish Scale.
+Future research should investigate whether training methods can prepare users to recognize AI-crafted phishing attempts.
+6. Conclusion
+This work provides real-world evidence on the efficacy of phishing awareness training and offers the largest validation to date of the NIST Phish Scale through its application under multiple experimental conditions.
+Difficulty strongly predicted user behavior (click rates from $7.0\%$ for easy lures to $15.0\%$ for hard; F(2,12086)=41.415,p<0.001F(2,12086)=41.415,\,p<0.001), while neither interactive nor lecture-based training produced statistically significant changes in click rates (pp=0.450) or reporting behavior (pp=0.417).
+Effect sizes for all training effects were minimal.
+These results confirm current research and run counter to industry assumptions and marketing claims regarding the efficacy of training tools.
+Our evidence indicates that training provides only modest protection against evolving phishing threats and should be regarded as just one component of a broader security awareness program.
+Our results reproduce those of prior work while evaluating new hypotheses and a new context.
+Appendices
+The extended material presented here consists of:
+Appendix A Email Controls in the Studied Organization
+Figure 3 depicts the studied organization’s email processing flow, with defenses noted and our metrics placed in the broader context.
+Figure 3.
+Phishing defense architecture, mapping the email pathway from authentication through post-delivery response.
+Technical controls create multiple interception points prior to human interaction, and human reports provide a feedback opportunity for organizational-level threat mitigation.
+:
+The following is a detailed sequence of the controls in the studied organization’s email defense setup.
+Malicious Email Sent:
+An attacker crafts a convincing message and sends it from a look-alike or compromised account, often using links or attachments that lead to theft or malware [67].
+Email Authentication: SPF:
+The domain owner lists which mail servers are allowed to send on its behalf; receivers check the sender’s IP against that list to spot forgeries [46].
+Email Authentication: DKIM:
+The sender adds a digital signature to the message; receivers use a public key in DNS to confirm the message really came from that domain and wasn’t altered [19].
+Email Authentication: DMARC:
+Builds on SPF/DKIM to tell receivers what to do with failures (monitor, quarantine, or reject) and sends reports back to the domain owner [47].
+Email Authentication: BIMI:
+A visual cue: providers that support it can display a verified brand logo when strong DMARC is in place; it helps users spot impostors [7, 9].
+Email Authentication: ARC:
+Preserves original authentication results across forwarding/mailing lists so legitimate forwarded mail isn’t penalized [4].
+Transport Encryption: STARTTLS:
+Upgrades the SMTP connection to TLS so email isn’t readable in transit between servers [67].
+Transport Enforcement: MTA-STS:
+Lets a domain publish a policy that its email must use TLS with valid certificates—blocking downgrade/MITM attempts [54].
+Transport Enforcement: DANE:
+Uses DNSSEC to bind TLS certificates to the domain for SMTP, strengthening spoofing/interception resistance [27, 44].
+Secure Email Gateway (SEG):
+The organization’s front door for mail: filters spam/phishing, checks sender/domain reputation, analyzes headers/content, and often rewrites links for time-of-click checks [67, 64].
+Threat Intelligence Integration & Update:
+Feeds of known-bad IPs, domains, URLs, and file hashes; standards like STIX/TAXII and programs like CISA AIS for sharing/automation [71, 72, 21].
+AI/ML Threat Detection:
+Behavioral and language models look for impersonation and anomalies beyond fixed indicators—useful for brand-new or targeted lures [35, 36].
+Attachment Sandboxing:
+Opens attachments in an isolated virtual environment (“detonation”) to observe malicious behavior and block delivery [64].
+Content Disarm & Reconstruction (CDR):
+Strips active content (macros, scripts) from files and rebuilds a safe version so users can still open the content [67].
+Data Loss Prevention (Outbound Scanning):
+Watches outgoing mail for sensitive data (e.g., customer lists, PHI); can also flag compromised accounts trying to exfiltrate information [65].
+Delivered to User Inbox:
+Messages that pass earlier controls arrive for the user—this is where human judgment and endpoint protections matter [67].
+Email Client Security:
+Clients/services block auto-loading images, scan/contain attachments, and add safety banners to reduce risk on open/click [33, 57].
+Security Awareness Training & Simulation:
+Short, regular training and realistic simulations teach users what to look for and build a habit of reporting [69].
+User Reports Phish:
+One-click “Report Phish” sends the message to security; reported samples help purge look-alikes and improve filters [34, 58].
+SIEM/SOAR Integration:
+Centralizes mail/endpoint alerts and runs automated playbooks to block senders, remove messages, and open tickets [17, 70].
+Incident Response & Automated Hunting:
+Analysts confirm threats, retro-hunt for variants, remove messages, and contain affected accounts/devices [70, 63].
+Zero-Hour Auto Purge (ZAP):
+If a campaign is discovered post-delivery, the system can retroactively remove those emails from mailboxes [62].
+Email Quarantine Management:
+Suspicious but unconfirmed messages are held for review so security teams (or users, with guardrails) can release or delete [60, 37].
+Mobile Device Management (MDM):
+Enforces secure mail apps, screen locks, and remote wipe on phones/tablets that access corporate email [66].
+Email Retention & Backup Systems:
+Archiving/backups keep searchable records for investigations/compliance—and let you restore mailboxes after an incident [86, 59].
+Appendix B Measuring Lure Complexity with the NIST Phish Scale
+As discussed in Appendix 2.2, we used the NIST Phish Scale as a standardized measure of phishing lure difficulty, i.e., the relative difficulty of detecting whether or not a given email is part of a phishing attack.
+The NIST Phish Scale is intended to help organizations benchmark lure complexity and contextualize simulation outcomes [14].
+It categorizes lure difficulty along two dimensions, shown in Table 5.
+Barrientos et al.
+[6] conducted early validation work on the NIST Phish Scale, while Dawkins and Jacobs [23] provided the official implementation guidelines.
+Ours is the first study to validate the Phish Scale’s premise at scale in a real-world context.
+Table 5.
+NIST Phish Scale for Detection Difficulty [82, 23].
+Phishing Cues represent observable errors or inconsistencies in the email that might alert users to its fraudulent nature, e.g., suspicious URLs or formatting irregularities [13].
+Premise Alignment measures the relevance of the email to the recipient’s organizational context [5].
+Cues \ Alignment Weak Medium Strong
+Few Medium Hard Hard
+Some Easy-to-Medium Medium Hard
+Many Easy Medium Medium
+Table 6.
+NIST Phish Scale Assessment of Phishing Templates Used In This Study.
+Premise Alignment scores range from 1 (Low) to 3 (High).
+Lure Scores range from 1-3 based on phishing cue assessment.
+Cue Count represents the total number of suspicious indicators identified in each template.
+Based on lure and premise scores, we mapped the top group to “Easy”, the middle group to “Medium”, and the bottom group to “Hard”.
+Template Name Difficulty Cue Count Lure Score Premise Score Premise Alignment Explanation
+Attachment - XLS Easy 14 2 2 Moderate alignment: XLS financial data plausible, but lacks sender/context.
+Fax via Voicemail Office Easy 13 2 1 Low alignment: Fax delivery uncommon in modern workplace.
+Email Not Delivered — Office Easy 13 2 3 High alignment: Bounce messages routine and widely expected.
+Email Not Delivered — Google Easy 16 3 3 High alignment: Gmail bounce notices with technical detail believable.
+From Marketing: Collab Request Easy 14 2 3 High alignment: Internal collaboration messages from marketing common.
+Accounting Team Amex Charges — Office Easy 16 3 3 High alignment: Finance inquiries during month-end familiar.
+HSA Use Balance! Easy 17 3 2 Moderate alignment: HSA reminders expected, but lacks personalization.
+Tax Workshop Easy 16 3 2 Moderate alignment: HR tax webinars plausible but weakened by generic formatting.
+Device Non-compliant - Microsoft Medium 15 3 3 High alignment: Matches expected MDM behavior and terminology.
+Mobile Device Restricted — Outlook Medium 16 3 3 High alignment: Device quarantine messages match common security flows.
+Device Non-compliant - Google Medium 16 3 3 High alignment: Technical restriction notices align with enterprise environments.
+Attachment - Word Hard 14 2 2 Moderate alignment: Document sharing plausible, but placeholder fields reduce realism.
+Vulnerabilities — Office Hard 18 3 1 Low alignment: Scenario vague and not role-specific; unlikely context.
+Loom - Recording Shared Hard 14 2 2 Moderate alignment: Loom recordings common, but template fields weaken trust.
+Mentor Program Hard 15 3 3 High alignment: Internal mentorship invitations believable and well-framed.
+HR - Performance Review — Outlook Hard 16 3 3 High alignment: Outlook calendar invites for reviews routine.
+Changes to Healthcare Form Hard 15 3 3 High alignment: Healthcare form updates common in HR communications.
+Vulnerabilities — Google Hard 16 3 1 Low alignment: Scenario lacks specificity and role-relevant context.
+Appendix C Lures
+See Table 6 for our analysis of the phishing lures used in this study.
+Following the NIST Phish Scale framework, we created a 3×3 classification matrix combining cue visibility (few-low, some-medium, many-high) with premise alignment (weak-low, medium, strong-high).
+Templates were assigned composite difficulty scores by summing both dimensions, with scores ranging from 2 (easy: high cues, low alignment) to 6 (hard: low cues, high alignment).
+Based on lure and premise scores, we mapped the top group to “Easy”, the middle group to “Medium”, and the bottom group to “Hard”.
+The deployed templates were constrained by organizational considerations.
+ACME Corp.’s management requested the exclusion of highly distressing content (e.g., emergency family situations, personnel security scenarios) to minimize employee psychological impact.
+This constraint reflects the practical limitations of conducting research within operational environments, where employee welfare considerations influence experimental designs [25].